### v1

# Sentences over Random Groups I: Existential Sentences

Abstract
Random groups of density $d<\frac{1}{2}$ are infinite hyperbolic, and of density $d>\frac{1}{2}$ are finite. We prove that for any given system of equations $\Sigma$, all the solutions of $\Sigma$ over a random group of density $d<\frac{1}{2}$ are projected from solutions of $\Sigma$ over the free group $F_{k}$, with overwhelming probability, where $k$ is the rank of the group. We conclude that any given sentence in the Boolean algebra of universal sentences, is a truth sentence over $F_{k}$ if and only if it is a truth sentence over random groups of density $d<\frac{1}{2}$, with overwhelming probability.
1 Introduction
Around 1945, a well-known question was presented by Alfred Tarski on the first order theory of free groups. He asked if every two non-abelian finitely generated free groups are elementary equivalent. His question was answered affirmatively by Z. Sela through his seminal work, which was published in a series consisting of seven papers titled “Diophantine Geometry”.
Theorem.
(Z. Sela, [Se1, Se2, Se3, Se4, Se5, Se6, Se7]) Every two non-abelian free groups share the same first order theory.
For a different approach to Tarski’s problem, see also [Kh-My].
Actually, Sela obtained a classification of all the f.g. (finitely generated) groups that are elementary equivalent to a non-abelian f.g. free group.
Hence, all the non-abelian f.g. free groups share the same collection of truth sentences. This collection is called the theory of free groups. In this paper, and in the coming papers of this series, we will be interested in the groups that can or cannot be distinguished from the free groups by a single sentence. More precisely, we aim to prove that “almost all the groups” cannot be distinguished from the free groups by a single sentence.
In order to describe the mathematical meaning of the term “almost all the groups”, we need to fix a model of groups through which one can pick a group at random. A model for picking a finitely presented group at random was suggested by M. Gromov in [Gr1]. This model has two fixed parameters $0\leq d\leq 1$ and an integer $k\geq 2$. The value of $d$ is called the density of the model, and $k$ is the rank of a fixed free group $F_{k}$ with a fixed basis $a=(a_{1},...,a_{k})$. All the groups in this model are given by presentations in terms of the generating set $a$. The number of relations in a group of level $l$ in the density model, grows exponentially with $l$. This is in contrast to the few relators model, also introduced by Gromov in [Gr2], where the number of relations is fixed in advance.
Picking a group at random in the density model, enables us to consider the probability of the satisfaction of a given property in that random group. For example, in [Gr1] Gromov proved that in all the interesting densities $d<\frac{1}{2}$, a random group is infinite hyperbolic in overwhelming probability (tends to $1$). This fact gives a possible answer for the question “What does a generic group look like?”. The answer is “It is hyperbolic almost always”. It was proved in [Gr1] that at densities $d>\frac{1}{2}$, the random groups are either trivial or isomorphic to $\mathbb{Z}/2\mathbb{Z}$, with overwhelming probability.
In 2013, J. Knight raised the following natural question: for any given first order sentence $\psi$, the sentence $\psi$ is a truth sentence over free groups if and only if $\psi$ is a truth sentence over random groups (in overwhelming probability) in the few relators model. In [Kh-Sk], O. Kharlampovich and R. Sklinos stated a similar conjecture, but for random groups in the density model with densities $d<\frac{1}{16}$. We extend these conjectures to the following more general conjecture.
Conjecture 1.1.
Let $0\leq d<\frac{1}{2}$ and let $\psi$ be a first order sentence in the language of groups. The random group of density $d$ satisfies $\psi$ (in overwhelming probability) if and only if the free group $F_{k}$ satisfies $\psi$.
In this paper, we prove 1.1 for sentences that belong to the Boolean algebra of universal sentences. In the next papers we consider general sentences.
Theorem 1.2.
Let $0\leq d<\frac{1}{2}$. Let $\psi$ be a first order sentence in the language of groups, and assume that $\psi$ belongs to the Boolean algebra of universal sentences. Then the random group of density $d$ satisfies $\psi$ (in overwhelming probability) if and only if the free group $F_{k}$ satisfies $\psi$.
The statement of 1.2 was proved for densities $d<\frac{1}{16}$ in [Kh-Sk], using a different argument than the one we present in this paper.
In [Se1, Se2, Se3, Se4, Se5, Se6, Se7], Sela introduced new mathematical tools, that can be applied on a given sentence $\psi$ over a non-abelian free group. Roughly speaking, when a sentence $\psi$ is a truth sentence over some non-abelian f.g. free group, these tools give a syntactic “proof” for the correctness of that sentence. These tools, together with the “proof” that is “written” by them, are shared uniformly by the various non-abelian f.g. free groups. In particular, the “proof” of $\psi$ obtained over a single non-abelian f.g. free group can be used as a “proof” for the correctness of $\psi$ over all the other f.g. free groups.
Our strategy is proving that all of these tools, developed over free groups, can be applied over a random group without changing them or their functionality.
The most basic among these tools is the Makanin-Razborov diagram, which was introduced in [Se1]. The Makanin-Razborov diagram of a given system of equations over a free group is a finite diagram that describes the structure of the set of solutions of that system. The development of the more advanced tools introduced by Sela, rely heavily on the Makanin-Razborov diagram.
Hence, it is natural that our initial step towards the proof of the conjecture is showing that the Makanin-Razborov diagram of a given system of equations can be used uniformly for describing the sets of solutions of that system over random groups.
In order to do so, we will prove the following statement.
Theorem 1.3.
(Main Theorem) Let $k\geq 2$ be an integer, let $d<\frac{1}{2}$ be a real number, and consider the density model with $k$ generators and density $d$. Let $\Sigma_{0}$ be a system of equations. Then the random presentation $\pi_{\Gamma}:F_{k}\rightarrow\Gamma$ of density $d$ satisfies the following property with overwhelming probability.
The set of solutions of the system $\Sigma_{0}$ in $\Gamma$ equals the projection to $\Gamma$ of the set of solutions of $\Sigma_{0}$ in $F_{k}$.
The majority of the work in this paper is devoted to the proof of 1.3. For doing that, we rely partially on some of the techniques introduced originally in [Gr1] and stated formally in [Ol1]. These techniques lead us to a better understanding for the hyperbolic nature of random groups, an understanding that leads to an application of a shortening procedure. The shortening procedure was introduced originally in [Ri-Se], and used in [Se1] mainly for constructing the Makanin-Razborov diagram of a given system of equations over free groups.
The paper is organized as follows.
In 2, we recall the definition of the Gromov density model.
In 3, we recall the definition of a Van-Kampen diagram and a decorated Van-Kampen diagram. Decorated Van-Kampen diagrams appeared originally in [Ol1], and were used in [Kh-Sk]. We need to handle problems regarding lengths in the Cayley graphs of Random groups, rather than triviality problems. Therefore, we found it necessary to extend this object. Then, in 3, we continue by introducing the procedure of “generally reducing” a decorated Van-Kampen diagram. This procedure will be necessary for the purposes of bounding probabilities of the satisfaction of equations over random groups.
In 3, we also introduce “mining modification”, which will be applied on decorated diagrams. The purpose of this modification is to enable us to handle variables which are supposed to get values of small lengths, and that occur in equations over random groups. Also in 3 we further estimate bounds for the probabilities of a random group to fulfill a decorated diagram.
In 4, we recall some facts from [Ol1], regarding the hyperbolicity of random groups, and we bound the number of decorated Van-Kampen diagrams in terms of appropriate parameters.
In 5, we prove some facts that hold in free groups, and that are related to commutativity in free groups. We will count on these facts in order to be able to handle diagrams with no cells over the random group (i.e., trees).
In 6, we prove that there are no new solutions of bounded lengths of a given system over random groups, i.e., all such solutions are projected from solutions in free groups.
In 7, we start to handle lengths problems in random groups. In particular, we bound the “time” for which the axes of two non-commutative elements could travel together (in close distance). This bound is interpreted in terms of the level of the random group under consideration (and the density parameter $d$). This bound will play an essential role during the application of the shortening procedure in the coming sections. LABEL:Sec:Axes-of-Elements-in-Random-Groups is the most technically involved section in this paper.
In 8, we use the facts proved in the previous sections in order to understand the hyperbolic nature of a random group.
In 9, we apply a shortening procedure on sequences of solutions of systems of equations over random groups. As stated above, the shortening procedure was introduced originally in [Se1] in order to construct the Makanin-Razborov diagram of a given system over free groups. Here, in 9, our goal in applying the shortening procedure is proving the non-existence of “non-lift sequences” over random groups, i.e., the non-existence of sequences of solutions that are not projected from free groups.
Finally, in 10, we deduce 1.2, and that the Makanin-Razborov diagram of a given system of equations over free groups, encodes all the solutions of that system over a random group.
Acknowledgment. I am indebted to my advisor Zlil Sela who introduced me to this problem. He shared his knowledge and ideas with me, and without his input, I couldn’t have completed this work.
2 The Gromov Density Model
We fix an integer $k\geq 2$, a free group $F_{k}$, and a basis $a=(a_{1},...,a_{k})$ of $F_{k}$. Given a property $P$, and a tuple $T$ of words in $F_{k}$, we may consider the following claim: the group presentation $\langle a:T\rangle$ satisfies the property $P$. For example, we may choose the property $P$ to be “the group is Hopfian”. Then, given a tuple $T$ of words in $F_{k}$, our claim becomes: the group $\langle a:T\rangle$ is Hopfian. The property $P$ may also be a property of a presentation, such as “the group presentation is small cancellation $C^{\prime}(1/6)$”. Then, given a tuple $T$ of words in $F_{k}$, the claim will be: “the presentation $\langle a:T\rangle$ is $C^{\prime}(1/6)$”. Since every presentation with generators $a$ has a natural interpretation of the elements of the free group $F_{k}$, the claim $P$ can contain constants too. For example, $P$ can be the claim “the word $a_{1}a_{2}a_{1}^{-1}a_{3}^{2}$ is trivial”.
Suppose further that the tuples that we consider are ordered in levels, i.e., for each integer $l$, we are given a collection $\mathcal{M}_{l}$ of tuples consisting of words in the free group $F_{k}$. Then, the description of the claim $P$ could also include the parameter $l$, such as the claim “the group is $2l$-hyperbolic”.
The group presentations defined by the tuples in the collection $\mathcal{M}_{l}$ are called the groups of level $l$. Once we define the groups of level $l$, for all integers $l$, we obtain a probability model of groups.
The way that a probability model of groups is used in general, is as follows. We fix a claim $P$ and a collection of tuples ordered in levels $\mathcal{M}_{l}$. And then, for every integer $l$, we consider the following number:
$p_{l}:=\frac{|\{T\in\mathcal{M}_{l}:\text{The group presentation }\langle a:T\rangle\text{ satisfies the property }P\}|}{|\mathcal{M}_{l}|}\,.$
This number is called the probability that the random group of level $l$ (in the relevant model) satisfies the property $P$. Then, our interest will be in the asymptotic behavior of $p_{l}$. I.e., we check if the limit $\underset{l\rightarrow\infty}{\lim}p_{l}$ exists. If the limit
$p_{0}=\underset{l\rightarrow\infty}{\lim}p_{l}$
do exist, we say that the probability that the random group (in the relevant model) satisfies the property $P$ is the real number $p_{0}$. If in addition we have that $p_{0}=1$, we say then that the random group (in the relevant model) satisfies the property $P$ in overwhelming probability.
In this paper, we are interested in a specific model, which was introduced in [Gr1] by M. Gromov.
Definition 2.1.
(Gromov density model). Fix a natural number $k\geq 2$, fix a free group $F_{k}$ with a fixed basis $a=(a_{1},...,a_{k})$, and fix a real number $0\leq d\leq 1$. The number $d$ is called the density of the model. Given an integer $l$, we denote by $B_{l}$ the set of cyclically reduced words of length $l$ in $F_{k}$
$B_{l}:=\{r\in F_{k}:|r|=l,\ r\text{ is cyclically reduced}\}\,.$
We also define the set $\mathcal{M}_{l}$, called the set of group presentations of level $l$, to be the collection of all the tuples $T=\left(r_{1},...,r_{s(l)}\right)$ consisting of $s(l)=|B_{l}|^{d}$ elements $r_{i}\in B_{l}$, $i=1,...,\left[s(l)\right]$.
Given a property $P$, we call the number
$p_{l}:=\frac{|\{T\in\mathcal{M}_{l}:\text{The group presentation }\langle a:T\rangle\text{ satisfies the property }P\}|}{|\mathcal{M}_{l}|}\,,$
the probability that the random group of level $l$ and density $d$ satisfies the property $P$. If the limit $p_{0}=\underset{l\rightarrow\infty}{\lim}p_{l}$ exists, then we call it the probability that the random group of density $d$ satisfies the property $P$.
If $p_{0}=1$, we say that the random group of density $d$ satisfies the property $P$ with overwhelming probability. In this case we say also that almost all the groups of density $d$ satisfy the property $P$. If $p_{0}=0$ in contrast, we say then that the property $P$ is negligible in density $d$.
[…]
Note that if the probability that the random group of density $d$ satisfies the property $P$ is $p_{0}$, then the probability that the random group of density $d$ satisfies the property $\neg P$ is $1-p_{0}$. In particular, the random group of density $d$ satisfies the property $P$ if and only if the property $\neg P$ is negligible in density $d$.
As the reader may note, 2.1 defines a family of models rather than a single one. For each value $0\leq d\leq 1$ of $d$, there is a model defined. Actually, random groups in those various models may behave differently, depending on the fixed density, even for properties that are naturally brought up in the context of groups. For example, we have the following facts.
Theorem 2.2.
([Ol2])
  1. 1.
    For $d<\frac{1}{5}$, almost all the presentations of density $d$ are Dehn presentations.
  2. 2.
    For $d>\frac{1}{5}$, almost all the presentations of density $d$ are not Dehn presentations.
In fact, for densities $d>\frac{1}{2}$, the density models are trivial.
Theorem 2.3.
([Gr1]) For $d>\frac{1}{2}$, almost all the groups of density $d$ are either trivial or isomorphic to $\mathbb{Z}/2\mathbb{Z}$.
3 Van-Kampen Diagrams
3.1 Van-Kampen Diagrams over Group Presentations
Let $l\geq 1$ be an integer. We want to define Van-Kampen diagrams over groups (presentations) with relations each of length $l$.
First, we start by defining topological Van-Kampen diagrams.
Definition 3.1.
A topological Van-Kampen diagram (of level $l$), usually denoted by $TD$, is a planar graph whose bounded faces are tiled by cells (as disks). The boundary of each cell in the diagram $TD$ consists of exactly $l$ edges. The boundary of each cell admits a starting vertex and a direction (between two: clockwise or counterclockwise), and the boundary of the whole diagram $TD$ admits a starting vertex and a direction.
An equivalence relation is defined between the cells, whose classes are called numberings, and are linearly ordered. In particular, different cells may admit the same numbering. The direction of an edge $e$ with respect to the cell $c$, where $e$ belongs to the boundary of $c$, is defined to be the direction of the cell $c$, and the order of $e$ in $c$ is the natural one induced by the starting point and the direction of $c$. An edge $e$ in the diagram $TD$ is called a filament, if $e$ does not belong to (the boundary of) a cell.
A topological Van-Kampen diagram without numbering is to forget the numberings of the cells in the topological Van-Kampen diagram (i.e., to forget the equivalence relation defined on the cells).
For every topological Van-Kampen diagram $TD$ we associate an auxiliary graph $K$, defined as follows. Assume that the cells of $TD$ admit exactly $n$ different numberings, which we denote $1,...,n$. For each $i=1,...,n$, we introduce $l$ different vertices, that correspond to the (directed) edges of a (single copy of a) cell numbered $i$ in $TD$. The matchings between the edges of the various cells inside the diagram $TD$, naturally suggests the following way for connecting the vertices in the graph $K$. Let $1\leq i,j\leq n$ be two numberings, and let $v_{i},v_{j}$ be two vertices associated to the numberings $i,j$ in $K$ respectively. If an edge $e_{i}$ represented by the vertex $v_{i}$ in some cell $c_{i}$ numbered $i$ inside the diagram $TD$ equals the edge $e_{j}$ represented by the vertex $v_{j}$ in some cell $c_{j}$ numbered $j$ in $TD$, then, we connect $v_{i}$ to $v_{j}$. If further the direction of $e_{i}$ w.r.t. $c_{i}$ equals the direction of $e_{j}$ w.r.t. $c_{j}$, we label the edge between $v_{i}$ and $v_{j}$ by $-1$, and by $1$ otherwise.
The topological Van-Kampen diagram $TD$ is called reduced, if there exist no loops (edges between a vertex and itself) in the associated graph $K$. Hence, by iteratively eliminating pairs of cells $c,c^{\prime}$ with the same numbering that share a common edge $e$ of the same order in $c$ and in $c^{\prime}$ (and identifying pairs of edges in the boundary of the constructed hole accordingly), every non-reduced topological Van-Kampen diagram $TD$ can be converted to a reduced one while reducing the number of cells, and without changing the boundary (length) of the diagram. We call the procedure of eliminating all such pairs of cells from the diagram standard reduction of the diagram.
Once we assign for each (directed) edge in a topological Van-Kampen diagram $TD$ a letter in the alphabet $a^{\pm}$, the boundary of each cell in $TD$ will be given some word in $F_{k}$, called the contour of that cell. Similarly, the boundary of the whole diagram will be given a word in $F_{k}$, called the contour of the diagram.
Definition 3.2.
Let $\Gamma=\langle a:\mathcal{R}\rangle$ be a group (presentation) whose relations are all (reduced words) of length $l$. A Van-Kampen diagram over $\Gamma$, denoted usually by $VKD$, is a topological Van-Kampen diagram $TD$ without numbering, together with assigning to each edge in $TD$ a letter in the alphabet $a^{\pm}$, so that the contour of every cell is a relator of $\Gamma$. If there exists no relator $r$ of $\Gamma$ so that $r$ is the contour of two distinct cells $c,c^{\prime}$ in $TD$ that share an edge $e$ so that $e$ has the same orders in $c,c^{\prime}$ (but the direction of $e$ in $c$ differs from the direction of $e$ in $c^{\prime}$), then, we say that the Van-Kampen diagram $VKD$ is reduced.
If $TD$ is a topological Van-Kampen diagram without numbering, then, we say that the group (presentation) $\Gamma$ fulfills the topological diagram without numbering $TD$ if one can assign to each edge in $TD$ a letter in the alphabet $a^{\pm}$, so that the obtained diagram is a Van-Kampen diagram over $\Gamma$.
Let $TD$ be a topological Van-Kampen diagram (with numbering), that admits $n$ distinct numberings. Let $t=(t_{i})$ be an $n$-tuple of reduced words in $F_{k}$. We say that the $n$-tuple $t$ fulfills the topological diagram $TD$ if one can assign to each edge in $TD$ a letter in the alphabet $a^{\pm}$, so that for every numbering $i=1,...,n$, the contours of the cells numbered $i$ in $TD$ are all equal to the word $t_{i}$. We say that the group (presentation) $\Gamma$ fulfills the topological diagram $TD$ if there exists an $n$-tuple $t$ consisting of pairwise distinct relators of $\Gamma$, so that $t$ satisfies $TD$.
Remark 3.3.
Let $TD$ be a topological Van-Kampen diagram, and consider its associated graph $K$.
  1. 1.
    If $n$ is the amount of distinct numberings in $TD$, then, the group $\Gamma$ fulfills $TD$, if and only if $\Gamma$ admits an $n$-tuple of pairwise distinct relators that satisfy the restrictions defined by $K$.
  2. 2.
    If the group $\Gamma$ fulfills $TD$, then the obtained Van-Kampen diagram $VKD$ is reduced if and only if the topological diagram $TD$ is reduced.
A more general notion than topological Van-Kampen diagram, is a decorated topological Van-Kampen diagram.
Definition 3.4.
A decorated topological Van-Kampen diagram (of level $l$), usually denoted by $DTD$, is a topological Van-Kampen diagram (of level $l$), who is equipped with the following extra structure. The directed path (cycle) that represents the boundary of the diagram $DTD$ is partitioned into consecutive parts (i.e., they can pairwise intersect in at most two points and they cover all the boundary cycle), that are called the parts of the boundary of $DTD$. An equivalence relation is defined on these parts, and each part is given a direction (between two directions). Each class in this equivalence relation is called a variable, and the equivalence relation together with the parts of the boundary and their directions are called the decoration of $DTD$. If two parts represent the same variable, then they are required to be of the same length (number of edges). If $e$ is an edge that belongs to some part $p$ of the boundary, then, the direction of $e$ w.r.t. $p$ is defined to be the direction of $p$, and the order of $e$ in $p$ is the natural one induced by the starting point and the direction of $p$.
Informally, a variable $x$ can be written along a part of the boundary, and $x^{-1}$ (the same variable but the other direction) could be written on another part of the boundary of the same length as the previous one. A decoration of the boundary could be explained briefly by a word in the variables, e.g. $xyx^{-1}y^{-1}$.
We declare a subset (maybe empty) of the variables to be rigid, and another subset (maybe empty) to be constant. Every (directed) edge that lie in a part of the boundary that represents a constant variable, is given a labeling in advance, i.e., a letter from the alphabet $a^{\pm}$. Moreover, as a part of the definition, we require that each rigid variable contains only one element (only one part of the boundary represents that rigid variable). If $e$ is an edge in the diagram $DTD$, then $e$ is said to be a rigid edge if $e$ belongs to a rigid part of the boundary, and $e$ is called a constant edge if it belongs to a constant part of the boundary. A cell $c$ in the diagram $DTD$ is called an isolated cell, if it admits a rigid edge. A numbering $i$ of a cell in $DTD$ is called an isolated numbering, if all the cells with numbering $i$ in the diagram are isolated cells.
A decorated topological Van-Kampen diagram without numbering is to forget the numberings of the cells in the decorated topological Van-Kampen diagram.
A decorated topological Van-Kampen diagram is called reduced, if the underlying topological Van-Kampen diagram is reduced.
For every decorated topological Van-Kampen diagram $DTD$ we associate an auxiliary graph $K$. This graph is initialized to be the graph associated to the underlying topological Van-Kampen diagram. However, the structure of the decorated diagram $DTD$ (the existence of the decoration and possible cancellations between the parts along the boundary - filaments) implies a natural generalization of the standard notion of “contiguity” (matching) between edges in a topological Van-Kampen diagram.
This contiguity is defined by declaring two edges $e_{1},e_{2}$ in the diagram to be contiguous if $e_{1},e_{2}$ belong to some parts $P_{1}x,P_{2}x$ (respectively) of the boundary of $DTD$, so that the orders of $e_{1},e_{2}$ in $P_{1}x,P_{2}x$ (respectively) are equal, and $P_{1}x,P_{2}x$ represent the same variable $x$ in the decoration. The obtained minimal equivalence relation on the set of edges in $DTD$ is called the contiguity between edges in the decorated diagram $DTD$. Edges in the same contiguity class are called contiguous edges.
Motivated by this notion of contiguity, we extend the graph $K$ as follows. Let $1\leq i,j\leq n$ be two numberings, and let $v_{i},v_{j}$ be two vertices associated to the numberings $i,j$ in $K$ respectively. If an edge $e_{i}$ represented by the vertex $v_{i}$ in some cell $c_{i}$ numbered $i$ inside the diagram $DTD$ is contiguous to an edge $e_{j}$ represented by the vertex $v_{j}$ in some cell $c_{j}$ numbered $j$ in $DTD$, then, we connect $v_{i}$ to $v_{j}$. If further the direction of $e_{i}$ w.r.t. $c_{i}$ equals the direction of $e_{j}$ w.r.t. $c_{j}$, we label the edge between $v_{i}$ and $v_{j}$ by $-1$, and by $1$ otherwise.
A connected component of the graph $K$ that contains a vertex that represents some constant edge, is called a constant component. A vertex $v$ of $K$ is called a constant vertex, if it belongs to a constant component of $K$, and $v$ is called an isolated vertex, if $\{v\}$ is a non-constant connected component of $K$. An edge $e$ in a decorated topological Van-Kampen diagram $DTD$ is called a cusp edge, if $e$ is a non-rigid filament edge, so that $\{e\}$ is a contiguity class.
Remark 3.5.
  1. 1.
    Note that the only edge in the diagram $DTD$ that is contiguous to a non-filament rigid edge $e$, is $e$ itself.
  2. 2.
    A cusp edge could occur only when a variable is represented by exactly two parts on the boundary, and these parts meet in “dual” edges.
Definition 3.6.
Let $\Gamma=\langle a:\mathcal{R}\rangle$ be a group (presentation) whose relations are all (reduced words) of length $l$.
Let $DTD$ be a decorated topological Van-Kampen diagram without numbering. We say that the group (presentation) $\Gamma$ fulfills the decorated topological diagram without numbering $DTD$, if one can assign to each edge in $DTD$ a letter in the alphabet $a^{\pm}$, so that the obtained diagram is a Van-Kampen diagram over $\Gamma$, and so that the decoration is satisfied, i.e., for every part $p$ of the boundary, the word that is given to the part $p$ (according to the assignment), equals (graphical equality - not only in $\Gamma$) to the word given to any other equivalent part of the boundary.
Let $DTD$ be a decorated topological Van-Kampen diagram (with numbering), that admits $n$ distinct numberings. Let $t=(t_{i})$ be an $n$-tuple of reduced words in $F_{k}$. We say that the $n$-tuple $t$ fulfills the decorated topological diagram $DTD$, if one can assign to each edge in $DTD$ a letter in the alphabet $a^{\pm}$, so that for every numbering $i=1,...,n$, the contours of the cells numbered $i$ are all equal to the word $t_{i}$, and so that the decoration is satisfied. We say that the group (presentation) $\Gamma$ fulfills the decorated topological diagram $DTD$, if there exists an $n$-tuple $t$ consisting of pairwise distinct relators of $\Gamma$, so that $t$ fulfills $DTD$.
Remark 3.7.
Let $DTD$ a decorated topological Van-Kampen diagram, and consider its associated graph $K$. Denote by $n$ the amount of distinct numberings in $DTD$. Then, the group $\Gamma$ fulfills $DTD$, if and only if, $\Gamma$ admits an $n$-tuple of pairwise distinct relators that satisfy the restrictions defined by $K$.
3.2 Topological Van-Kampen Diagrams over Random Groups
Lemma 3.8.
Let $DTD$ be a decorated topological Van-Kampen diagram, that admits $n$ distinct numberings. Let $K$ be the graph associated to the decorated diagram $DTD$, and let $u$ be the amount of distinct non-constant connected components in $K$. Denote by $B_{l}$ the set of cyclically reduced words of length $l$ in $F_{k}$. Then, the probability that a random $n$-tuple of words in $B_{l}$ fulfills the diagram $DTD$ is at most
$\frac{(2k)^{n}\cdot(2k-1)^{u}}{B_{l}^{n}}\,.$
Proof.
To construct an $n$-tuple of cyclically reduced words that fulfills $DTD$, we start by specifying a label from $a^{\pm}$ for the first edge in every cell, and then we choose one label from $a^{\pm}$ for each non-constant connected component in the auxiliary graph $K$ associated to $DTD$.
Hence, we deduce that there exist at most $(2k)^{n}\cdot(2k-1)^{u}$ distinct $n$-tuples of cyclically reduced words of length $l$ that fulfill $DTD$.
Then, the probability that a random $n$-tuple of cyclically reduced words fulfills $DTD$ is at most
$\frac{(2k)^{n}\cdot(2k-1)^{u}}{B_{l}^{n}}\,.$
Lemma 3.9.
Let $DTD$ be a decorated topological Van-Kampen diagram, that admits $n$ distinct numberings. Denote by $B_{l}$ the set of cyclically reduced words of length $l$ in $F_{k}$. Then, the probability that the random group $\Gamma$ of level $l$ and density $d$, fulfills the decorated diagram $DTD$, is at most
$B_{l}^{nd}\cdot p\,,$
where $p$ is the probability that a random $n$-tuple of words in $B_{l}$, fulfills the decorated diagram $DTD$.
Hence, the probability that the random group $\Gamma$ of level $l$ and density $d$, fulfills the decorated diagram $DTD$ is at most
$\frac{(2k)^{n}\cdot(2k-1)^{u}}{B_{l}^{n\cdot(1-d)}}\,,$
where $u$ is given in 3.8.
Proof.
Recall that the random group $\Gamma$ of level $l$ and density $d$, is defined by choosing in random a $B_{l}^{d}$-tuple of cyclically reduced words of length $l$, and note that, by definition, the group $\Gamma$ fulfills $DTD$ if and only if $\Gamma$admits an $n$-tuple of pairwise distinct relators that fulfills the diagram $DTD$.
Hence, the probability that the random group of level $l$ and density $d$ fulfills $DTD$, is at most the probability that when choosing in random a $B_{l}^{d}$ words $w=(w_{i})$ from $B_{l}$, picked one after the other, one of the $n$-tuples $(w_{i_{1}},...,w_{i_{n}})$, where $1\leq i_{1},...,i_{n}\leq B_{l}^{d}$, fulfills $DTD$.
Hence, the probability that the random group of level $l$ and density $d$ fulfills $DTD$, is bounded by
$T_{n}\cdot p\,,$
where $T_{n}$ is the amount of distinct $n$-tuples consisting of elements in the sequence $1,2,...,B_{l}^{d}$. That is, the probability that the random group of level $l$ and density $d$ fulfills $DTD$, is bounded by
$B_{l}^{nd}\cdot p\,.$
Lemma 3.10.
Let $DTDWN$ be a decorated topological Van-Kampen diagram without numbering, and let $\mathcal{A}$ be the collection of all the reduced decorated topological Van-Kampen diagrams $DTD$ (with numberings) for which forgetting the numbering from $DTD$ yields the diagram $DTDWN$.
Then, the probability that the random group $\Gamma$ of level $l$ fulfills the diagram $DTDWN$ so that the obtained Van-Kampen diagram over $\Gamma$ is reduced, is at most the probability that the random group $\Gamma$ fulfills one of the diagrams in $\mathcal{A}$.
Proof.
Obvious. ∎
Example 3.11.
Consider the diagram composed of two cells, each of length $l$ (i.e., $l$ edges), sharing more than $\frac{l}{6}$ consecutive edges. Assume that the cells share the same starting point, and admit opposite directions. This information gives a topological Van-Kampen diagram without numberings $D$. The probability that the random group of level $l$ fulfills $D$ is of course $1$, since any (non-free) group (presentation) fulfills this diagram.
On the other side, in this example there is only two ways to write numberings in the cells, which are $1,1$, and $1,2$ (the two cells are equivalent, or not otherwise). Let $D^{\prime}$ be the one with the first numbering, and $D^{\prime\prime}$ be the second.
The probability that the random group $\Gamma$ of level $l$ fulfills $D$ so that the obtained Van-Kampen diagram over $\Gamma$ is reduced, is bounded by the probability that $\Gamma$ satisfies $D^{\prime\prime}$, since $D^{\prime}$ is not reduced as a topological Van-Kampen diagram (and hence it will not provide any reduced Van-Kampen diagram over $\Gamma$). Let $K$ be the graph associated to $D^{\prime\prime}$. Then $K$ admits at most $\frac{11l}{6}$ connected components. Hence, the probability that the random group of level $l$ and density $d$ satisfies $D^{\prime\prime}$ is bounded by
$\frac{(2k)^{2}\cdot(2k-1)^{\frac{11l}{6}}}{B_{l}^{2\cdot(1-d)}}\leq\frac{(2k)^{2}}{(2k-1)^{(\frac{1}{6}-2d)\cdot l}}\,.$
In particular, this argument shows that if the density $d$ satisfies $d<\frac{1}{12}$, then the random group of density $d$ cannot admit two distinct relators that begin with the same $\frac{l}{6}$-piece. Of course, this argument can easily extended to prove that the random group of density $d<\frac{1}{12}$, is $C^{\prime}(\frac{1}{6})$-small cancellation (it was shown in [Gr1] that the random group of density $d<\frac{r}{2}$ is $C^{\prime}(r)$-small cancellation).
3.3 General Reduction of a Decorated Topological Van-Kampen Diagram
Let $DTD$ be a decorated topological Van-Kampen diagram. We describe the following modifications on $DTD$ that does not change the decoration of the diagram $DTD$.
Definition 3.12.
Let $DTD$ be a decorated topological Van-Kampen diagram, and let $K$ be the auxiliary graph associated to $DTD$.
Let $v$ be an isolated vertex of $K$ (i.e., $\{v\}$ is a non-constant connected component of $K$), that represents some edge $e$ in (the boundary of) some cell numbered $i$ in the diagram $DTD$. Let $C$ be the contiguity class of the edge $e$ in $DTD$. Recall that, since every rigid variable in the decoration corresponds only to one part of the boundary, if $C$ contains a non-filament rigid edge, then $C=\{e\}$ is a singleton.
  • If $C$ does not contain a rigid edge, we eliminate all the cells that contain an edge in $C$ (all of these cells share the numbering $i$, since $v$ is isolated), and stretch the filament edges that belong to $C$ accordingly (while keeping the parts of the boundary that does not contain an edge in $C$ unchanged - so that the decoration is preserved).
  • If $C$ contains a rigid edge, then, if not all of the cells containing an edge in $C$ are isolated, we eliminate all the cells that contain an edge in $C$, we stretch the filament non-rigid edges that belong to $C$ accordingly, and on each of the (filament) rigid edges in $C$, we paste a copy of the cell numbered $i$ accordingly (since we don’t want to change the rigid parts, we do not stretch the filament edges in $C$ that belong to rigid parts. Instead, we paste a copy of the corresponding cell on that edges correspondingly, without changing the rigid parts).
For clarifying the geometric picture in the last case, if $C$ contains a non-filament rigid edge, then we do nothing, and otherwise all the rigid edges in $C$ must be filaments, and in particular we paste on each of them a copy of the cell numbered $i$ accordingly, in condition that not all of the cells containing edges in $C$ are isolated (if all of them are isolated, in the last case, we do nothing).
We call this modification, a $(v,C)$-isolation applied on the decorated diagram $DTD$.
Lemma 3.13.
Let the notation be as in 3.12, and let $DTD^{\prime}$ be the decorated diagram obtained by applying a $(v,C)$-isolation on $DTD$. Let $K^{\prime}$ be the graph associated to $DTD^{\prime}$. Assume that the $(v,C)$-isolation did not reduce the amount of numberings. Then, the vertex (corresponding to) $v$ is (again) an isolated vertex in $K^{\prime}$.
Proof.
We keep the notation of 3.12. Let $f$ be an edge represented by $v$ in the diagram $DTD$, and let $F$ be the contiguity class of $f$ in $DTD$. Assume that $f$ is not contiguous to $e$ in $DTD$. Since $v$ is an isolated numbering of $K$, all of the cells in $DTD$ that admit an edge that belong to $F$ share the numbering $i$, and, each of these edges have the same order $r$ in that cells, which equals also to the order of $e$ in the corresponding cell.
Hence, by the definition of $(v,C)$-isolation, neither the filament edges in $F$, nor the non-filament edges in $F$, are affected by this modification. Hence, the collection of edges in $DTD^{\prime}$ corresponding to the edges in $F$, forms again a contiguity class in $DTD^{\prime}$.
Recalling that a $(v,C)$-isolation can add cells to the diagram only by pasting cells numbered $i$ on the rigid parts along the edges of order $r$, we conclude that every contiguity class of the edge of order $r$ in a cell numbered $i$ in the obtained diagram $DTD^{\prime}$, contains only filaments and edges that are of order $r$ in cells numbered $i$. That is, if the $(v,C)$-isolation did not eliminate the numbering $i$ totally from the diagram, then $v$ will keep to be an isolated vertex of $K^{\prime}$. ∎
Definition 3.14.
Let the notation be as in 3.12. In accordance with 3.13, the procedure of applying iterative $(v,C_{j})$-isolations on the decorated diagram $DTD$, is called isolating the numbering $i$, or briefly, a numbering isolation.
Lemma 3.15.
The numbering isolation procedure terminates after finitely many steps.
Proof.
Let $DTD$ be a decorated diagram, and let $v$ be an isolated vertex in the graph associated to $DTD$, corresponding to an edge in a cell numbered $i$ in $DTD$. By definition, every $(v,C_{j})$-isolation either reduces the amount of non-isolated cells with numbering $i$, or otherwise, it reduces the amount of cells with numbering $i$ in the diagram. Moreover, a $(v,C_{j})$-isolation does not increase the amount of non-isolated cells with numbering $i$.
Hence, the procedure of isolating the numbering $i$ must stop to reduce the amount of non-isolated cells numbered $i$ eventually. Hence, it must reduce the amount of cells numbered $i$ eventually. Hence, it must terminate eventually. ∎
Lemma 3.16.
Let $DTD$ be a decorated topological Van-Kampen diagram. An (effective) numbering isolation applied on $DTD$, keeps isolated numberings isolated, and does not increase the amounts of their (isolated) cells in the diagram. And, it either
  1. 1.
    Isolates a new numbering (that wasn’t isolated before), or
  2. 2.
    Reduces the amount of (isolated) cells of an isolated numbering, or
  3. 3.
    Reduces the amount of numberings in $DTD$.
Proof.
Let $K$ be the graph associated to $DTD$, and let $v$ be an isolated vertex of $K$ corresponding to an edge in a cell numbered $i$ in $DTD$. We consider the numbering isolation procedure corresponding to $v$.
First, note that if $j\neq i$ is another numbering, then the iteration will not change the amount of cells numbered $j$ in the diagram. In particular, if $j$ was an isolated numbering then the procedure above keeps it isolated, and preserves the amount of (isolated) cells numbered $j$.
Now assume that $i$ was isolated. Then, the procedure will not paste any new cell to the diagram (it only removes cells in this case). Hence, the numbering $i$ will stay isolated after the iteration, and, assuming the procedure was effective, the procedure reduces the amount of cells numbered $i$ in this case.
Now assume that $i$ was not isolated. Then, the procedure will eliminate every non-isolated cell numbered $i$ in the diagram, and (maybe) will paste cells numbered $i$ on rigid edges. Hence, $i$ will be an isolated numbering after the procedure. ∎
Definition 3.17.
Let $DTD$ be a decorated topological Van-Kampen diagram. Let $x$ be a non-constant variable in the decoration, and let $P_{1}x,...,P_{q}x$ be all the parts of the boundary of $DTD$ representing the variable $x$. Assume that $P_{1}x$ contains a self-intersection $S$.
For all $i=1,...,q$,
  • if the part $P_{i}x$ contains a self-intersection compatible with $S$, we eliminate it from the part $P_{i}x$, and
  • otherwise, we paste $S$ on the corresponding edges in $P_{i}x$ (filling the cut sub-diagram $S$ into $P_{i}x$ from the direction of the unbounded face of the graph $DTD$).
We call this modification an elimination of self-intersection. Note that self-intersection elimination modification, reduces the boundary length of the diagram.
Lemma 3.18.
Let $DTD$ be a decorated topological Van-Kampen diagram. An (effective) self-intersection elimination modification, reduces the boundary length of the diagram. Moreover, it keeps isolated numberings isolated, and does not increase the amounts of their (isolated) cells in the diagram.
Proof.
Let $x$, $P_{1}x,...,P_{q}x$, and $S$ be as in 3.17. If $x$ is a rigid variable, then, $q=1$, and $x$ is represented only by the part $P_{1}x$, so the claim is obvious. Assume now that $x$ is a non-rigid variable, and let $i$ be an isolated numbering in the diagram. Since $i$ is an isolated numbering, and $x$ is not a rigid variable, the cut sub-diagram $S$ does not contain any cell numbered $i$. ∎
Finally, we note that we can extend our “reduction process” to include also standard reduction:
Lemma 3.19.
Let $DTD$ be a decorated topological Van-Kampen diagram. Applying standard reduction on the diagram $DTD$ (with forgetting the decoration), reduces the amount of cells in the diagram. Moreover, it keeps isolated numberings isolated, and does not increase the amounts of their (isolated) cells in the diagram.
Proof.
Obvious. ∎
Lemma 3.20.
Let $DTD$ be a decorated topological Van-Kampen diagram. The procedure of consecutively applying numbering isolation, self-intersection elimination, and standard reduction, applied on the diagram $DTD$, terminates after finitely many steps.
Proof.
According to 3.16, 3.18, and 3.19, every iteration of the procedure, keeps isolated numberings isolated, and does not increase the amounts of their (isolated) cells.
However, since the sequence of amounts of numberings in the diagrams obtained along the procedure is descending, it must be stabilized eventually. Hence, the procedure must stop to isolate new numberings eventually. Hence, according to 3.16, every (effective) numbering isolation, will reduce the amount of (isolated) cells of the isolated numberings eventually. Hence, the procedure must stop to apply (effective) numbering isolation eventually.
But self-intersection elimination modification reduces the boundary length of the diagram, and standard reduction does not change the boundary length. Hence, the procedure must stop to apply self-intersection elimination modifications eventually.
However, standard reduction reduces the amount of cells in the diagram. Hence, the procedure then must stop to apply standard reduction eventually. ∎
Definition 3.21.
Let $DTD$ be a decorated topological Van-Kampen diagram. In light of 3.20, we summarize our terminology so that the procedure of consecutively applying numbering isolation, self-intersection elimination, and standard reduction, on the diagram $DTD$, is called generally reducing the decorated diagram $DTD$.
Hence, a generally reduced decorated diagram $DTD$ will mean a reduced decorated diagram with no self-intersections, and for which for every isolated vertex $v$ in the graph associated to $DTD$, the contiguity class $C$ of any edge $e$ in $DTD$ represented by $v$, contains a rigid edge, but all the cells that contain an edge in $C$ are isolated.
Remark 3.22.
Note that running general reduction on a decorated diagram, could affect the rigid parts of the boundary only by eliminating self-intersections from them. It is not hard to see that running general reduction on a decorated diagram, does not affect the constant parts of the boundary.
Theorem 3.23.
Let $DTD$ be a generally reduced decorated topological Van-Kampen diagram, with associated graph $K$. Let $A$ be the amount of isolated vertices in $K$, and let $R$ be the total amount of rigid edges in $DTD$. Then
$A\leq R\,.$
Moreover, if $i$ is a numbering of a cell in the diagram, so that there exists an isolated vertex $v$ in $K$ corresponding to the numbering $i$, then $i$ is an isolated numbering. In particular, if there exists no isolated numberings in the generally reduced diagram $DTD$, then $A=0$.
Proof.
Let $v$ be an isolated vertex of $K$, and let $C$ be a contiguity class of an edge in $DTD$ that is represented by the vertex $v$. If $C$ does not contain a rigid edge, then, by the definition of general reduction, the decorated diagram is not generally reduced. If $v^{\prime}\neq v$ is another isolated vertex of $K$, and $C^{\prime}$ is a contiguity class for an edge in the diagram represented by the vertex $v^{\prime}$, then $C$ and $C^{\prime}$ are disjoint sets (otherwise, $v,v^{\prime}$ will be connected by an edge).
For the last statement, assume that $v$ is an isolated vertex of $K$, and let $e$ be an edge represented by $v$ in $DTD$. Denote by $c$ a cell containing the edge $e$ in the diagram. If $c$ is not isolated, then, by the definition of general reduction, the diagram is not generally reduced. ∎
3.4 Mining in Decorated Van-Kampen Diagrams
We now describe a modification applied on a decorated diagram, a modification that could change the rigid parts.
Definition 3.24.
Let $DTD$ be a decorated topological Van-Kampen diagram. A mining modification in $DTD$ is eliminating all the (isolated) cells of an isolated numbering, from the diagram, without changing the non-rigid variables.
Lemma 3.25.
Let $DTD$ be a decorated topological Van-Kampen diagram. Let $DTD^{\prime}$ be the diagram obtained after an (effective) mining modification on $DTD$. Then,
  1. 1.
    The diagram $DTD^{\prime}$ contains less numberings than $DTD$.
  2. 2.
    The probability that the random group $\Gamma$ of level $l$ and density $d$, fulfills the original diagram $DTD$, is at most the probability that $\Gamma$ fulfills the obtained diagram $DTD^{\prime}$.
Proof.
Point 1 is obvious. For Point 2, we note that every group that fulfills $DTD$, must fulfill the obtained diagram $DTD^{\prime}$. ∎
After applying a mining modification on a decorated diagram, we generally reduce the obtained diagram. Thus, we extend the meaning of a mining modification, to include further generally reducing the obtained diagram.
3.5 Decorated Topological Van-Kampen Diagrams over Random Groups
Lemma 3.26.
Let $DTD$ be a decorated topological Van-Kampen diagram. Let $DTD^{\prime}$ be the diagram obtained by generally reducing $DTD$. Then, the probability that the random group $\Gamma$ of level $l$ and density $d$, fulfills the original diagram $DTD$, is at most the probability that $\Gamma$ fulfills the obtained generally reduced diagram $DTD^{\prime}$.
Proof.
We note that every group that fulfills $DTD$, must fulfill the diagram obtained by applying numbering isolation, or self-intersection elimination, or standard reduction, on $DTD$. ∎
Theorem 3.27.
Let $DTD$ be a generally reduced decorated topological Van-Kampen diagram, that admits $n$ distinct numberings. Let $R$ be the total amount of rigid edges in $DTD$. Denote by $B_{l}$ the set of cyclically reduced words of length $l$ in $F_{k}$.
Then, the probability that a random $n$-tuple of words in $B_{l}$ fulfills the diagram $DTD$ is at most
$\frac{(2k)^{n}\cdot(2k-1)^{\frac{nl+R}{2}}}{B_{l}^{n}}\,.$
Moreover, if there exist no isolated numberings in $DTD$, then the probability that a random $n$-tuple of words in $B_{l}$ fulfills the diagram is at most
$\frac{(2k)^{n}\cdot(2k-1)^{\frac{nl}{2}}}{B_{l}^{n}}\,.$
Proof.
Let $K$ be the auxiliary graph associated to the decorated diagram $DTD$. Since $DTD$ is generally reduced, we have, according to 3.23, that the amount $A$ of isolated vertices in $K$ is at most $A\leq R$, and in case $DTD$ contains no isolated numberings, we have that $A=0$.
Denote by $c$ the amount of constant vertices in $K$, and denote by $V$ the amount of non-isolated non-constant vertices of $K$. Of course, the graph $K$ consists of $nl$ vertices. Hence,
$V=nl-A-c\,,$
and by definition, every non-isolated non-constant vertex belongs to a connected component of $K$ of size at least $2$.
Hence, the amount of distinct non-constant connected components in $K$ is at most
$u\leq\frac{V}{2}+A\leq\frac{nl+A}{2}\,.$
Now 3.8 completes the proof. ∎
Theorem 3.28.
Let $DTD$ be a generally reduced decorated topological Van-Kampen diagram, that admits $n$ distinct numberings. Let $R$ be the total amount of rigid edges in $DTD$. Denote by $B_{l}$ the set of cyclically reduced words of length $l$ in $F_{k}$.
Then, the probability that the random group $\Gamma$ of level $l$ and density $d$, fulfills the decorated diagram $DTD$, is at most
$\frac{(2k)^{n}\cdot(2k-1)^{\frac{nl+R}{2}}}{B_{l}^{n\cdot(1-d)}}\leq\left(\frac{2k}{(2k-1)^{(\frac{1}{2}-d-\frac{R}{2nl})l}}\right)^{n}\,.$
Moreover, if there exist no isolated numberings in $DTD$, then the probability that $\Gamma$ fulfills the diagram is at most
$\left(\frac{2k}{(2k-1)^{(\frac{1}{2}-d)l}}\right)^{n}\,.$
Proof.
We only used 3.27 and evaluated in 3.9. ∎
4 Isoperimetric Inequality in Random Groups
4.1 Isoperimetric Inequality and the Hyperbolicity Constant
Theorem 4.1.
([Ol1]). At density $d<\frac{1}{2}$, for every $\epsilon>0$, with overwhelming probability, every reduced Van-Kampen diagram $D$ over a random group at density $d$ satisfies
$|\partial D|\geq(1-2d-\epsilon)\cdot l\cdot|D|\,,$
where $|D|$ is the number of faces in $D$, and $|\partial D|$ is the length of its boundary.
In particular, for $\epsilon=\frac{1}{2}-d$, we have with overwhelming probability that
$|D|\leq\frac{|\partial D|}{l\cdot(\frac{1}{2}-d)}\,.$
Remark 4.2.
Let $d<\frac{1}{2}$. Assume that we have a collection $\mathcal{V}$ of reduced decorated topological Van-Kampen diagrams (whose cells are of boundary length $l$), and that we are interested in bounding the probability that the random group $\Gamma$ of level $l$ and density $d$ fulfills some diagram in the collection $\mathcal{V}$.
Then, according to 4.1, we can drop all the diagram $D$ in our collection $\mathcal{V}$, for which $|D|>\frac{|\partial D|}{l\cdot(\frac{1}{2}-d)}$, since they add nothing (a negligible fraction) to the probability that we wanted to bound. In accordance with this, we will say that a diagram $D$ is among those diagrams that the random group $\Gamma$ (of level $l$ and density $d$) could satisfy, if $|D|\leq\frac{|\partial D|}{l\cdot(\frac{1}{2}-d)}$ (and each of the cells in $D$ are of boundary length $l$).
Theorem 4.3.
Let $d<\frac{1}{2}$, and denote
$C_{0}=\frac{\frac{1}{2}-d}{4}\,.$
Let $\Gamma$ be the random group of level $l$ and density $d$.
Then, with overwhelming probability, the ball of radius $C_{0}l$ in the Cayley graph of $\Gamma$, is a tree.
Proof.
Let $D$ be a reduced Van-Kampen diagram over $\Gamma$ for some relation of length at most $2C_{0}l$ in $\Gamma$. Then, according to 4.1, the number of cells in $D$ is bounded by
$|D|\leq\frac{1}{2}\,.$
Hence, $|D|=0$. ∎
Theorem 4.4.
([Ol2, Ol1]). At density $d<\frac{1}{2}$, for every $\epsilon>0$, with overwhelming probability, the random group $\Gamma$ of level $l$ is torsion-free $\frac{12l}{(1-2d-\epsilon)^{2}}$-hyperbolic.
In particular, $\Gamma$ is $\alpha_{0}l$-hyperbolic, where
$\alpha_{0}=\frac{48}{(1-2d)^{2}}\,.$
4.2 Counting Topological Van-Kampen Diagrams
Theorem 4.5.
Let $W$ be a word in some free group. Denote by $B(m,S)$ the amount of decorated topological Van-Kampen diagrams $DTD$ with decoration $W$, with at most $m$ cells each of boundary length $l$, and with boundary length at most $S$, so that every part of the boundary of $DTD$ (representing a variable in the decoration $W$), contains no self-intersection.
Then, the number $B(m,S)$ is bounded by
$B(m,S)\leq m\cdot\left(\left(4(ml)^{3}\right)^{m}\cdot S^{3}\right)^{m}\cdot 2S^{3|W|+1}\cdot|W|\cdot\left(2(|W|+2)\cdot\left(4(|W|+2)^{2}S\right)^{2(|W|+2)}\right)^{|W|}\,.$
In particular, there exists a polynomial $Q$, so that, if $m,|W|\leq\ln l$, Then,
$B(m,S)\leq\left(Q(l)\right)^{\ln^{2}l}\,.$
Proof.
First, let us bound the amount $C(m)$ of circular (no filaments) topological Van-Kampen diagrams with at most $m$ cells, each of length $l$. To construct a diagram in $C(m)$, we have to paste a new cell on a diagram in $C(m-1$). For that we have to choose two points on the boundary of that diagram in $C(m-1)$, and a direction (where to paste). Then, we choose a starting vertex, a direction, and a numbering for the added cell. Hence,
$C(m)\leq 2(ml)^{2}\cdot l\cdot 2\cdot m\cdot C(m-1)=4(ml)^{3}\cdot C(m-1)\,.$
Since $C(0)=1$, we get by induction that
$C(m)\leq\left(4(ml)^{3}\right)^{m}\,.$
Every diagram in $B(m,S)$, consists of at most $m$ circular components that belong to $C(m)$, and that are connected by (maybe complicated) trees. For constructing a diagram in $B(m,S)$, we start by connecting the circular components by lines, and after that we attach the remaining filament parts.
For doing that we start by choosing the amount $n\leq m$ of circular components. Then, we choose iteratively $n$ diagrams from $C(m)$, and in each iteration, we choose a point on the boundary of the diagram obtained in the previous iteration, we choose a point on the boundary of the chosen diagram in the present iteration, and then we connect these two points by a line that we have to choose its length. Since we want to get at the end a diagram in $B(m,S)$, whose boundary length is at most $S$, we can assume that the boundary length of each of the diagrams obtained or chosen in any iteration is at most $S$. Hence, for choosing $n\leq m$ circular components from $C(m)$ and connecting them so that the obtained diagram is a sub-diagram of a diagram in $B(m,S)$, we have at most
$m\cdot\left(C(m)\cdot S^{3}\right)^{m}$
different choices. Denote by $C^{\prime}(m)$ the collection of diagrams that are constructed in the last paragraph.
In order to construct a diagram in $B(m,S)$, we choose a diagram in $C^{\prime}(m)$ and attach the possible remaining filament components. These remaining filament components can be described by cancellation trees for solutions (in a free group) of length at most $S$ of equations of the form $UwV$, where $w$ is a sub-word of some cyclic permutation of $W$, and $U,V$ denote new variables. According to 4.6, the amount of such trees is bounded by
$T(|W|+2,S)\leq 2(|W|+2)\cdot\left(4(|W|+2)^{2}S\right)^{2(|W|+2)}\,.$
According to the assumption, every part of the boundary of a decorated diagram in $B(m,S)$ representing some variable in the decoration $W$, contains no self-intersection. Thus, the amount of trees of cancellation that we should attach to a diagram in $C^{\prime}(m)$ for obtaining a diagram in $B(m,S)$, is at most $t\leq|W|$. Hence, in order to construct a diagram in $B(m,S)$, we choose a diagram $D$ in $C^{\prime}(m)$, we choose the amount $t\leq|W|$ of trees from $T(|W|+2,S)$ to be attached to the boundary of $D$, we choose $t$ points $p_{i}$ on the boundary of $D$, we choose $t$ trees $T_{i}$ in $T(|W|+2,S)$, we choose a point $q_{i}$ on $T_{i}$ for each $i$, and we identify the points $p_{i}$ and $q_{i}$ for all $i$. For doing this, we have at most
$m\cdot\left(C(m)\cdot S^{3}\right)^{m}\cdot|W|\cdot S^{|W|}\cdot\left(T(|W|+2,S)\right)^{|W|}\cdot S^{|W|}\,.$
Let $C^{\prime\prime}(m)$ be the diagrams obtained in the last paragraph.
Finally, in order to construct a diagram in $B(m,S)$, we choose a diagram $D$ in $C^{\prime\prime}(m)$, we choose a starting vertex and a direction for the boundary of that diagram $D$. And then, for “writing” the decoration $W$ along the boundary of $D$, it remains only to specify the length of each variable in $W$. We conclude that
$B(m,S)\leq m\cdot\left(C(m)\cdot S^{3}\right)^{m}\cdot|W|\cdot S^{|W|}\cdot\left(T(|W|+2,S)\right)^{|W|}\cdot S^{|W|}\cdot 2S\cdot S^{|W|}\,.$
Lemma 4.6.
Let $L$ be an integer. Then, the amount $T(L,S)$ of combinatorial (counting the number of edges but not the labels) cancellation trees for all the solutions of circumference $S$ in a free group of the equation $W=1$, for some word $W$ of length at most $L$, is bounded by
$T(L,S)\leq 2L\cdot\left(4L^{2}S\right)^{2L}\,.$
Proof.
We note that any tree of cancellation in $T(L,S)$, can be topologically described by a tree with at most $2L$ vertices, and thus at most $2L$ edges. Let $b(L)$ be the total amount of such trees. Now for each tree in $b(L)$ we have to choose lengths for its edges. For that each edge has at most $S$ options, and hence in total we have at most $S^{2L}$ options. Finally, for creating a tree in $b(W)$, we have to choose the number $v\leq 2L$ of vertices, and then to choose $v-1$ pairs from a set of size $v$. Hence
$b(W)\leq 2L\cdot\left((2L)^{2}\right)^{2L}\,.$
5 Axes of Elements in Free Groups
Definition 5.1.
Let $G\in F_{k}$ (reduced word), and suppose that we have the graphical equality (with no cancellations)
$G=YgY^{-1}\,,$
for (reduced) words $Y,g\in F_{k}$, so that $g$ is cyclically reduced.
Then, the path in the Cayley graph of $F_{k}$ induced by the elements $Zg^{n}$, $n\in\mathbb{Z}$, is called the axis of the element $G$ in $F_{k}$, and is usually denoted by $A_{G,F_{k}}$. The translation length of $G$, denoted by $[G]_{F_{k}}$, is defined to be
$[G]_{F_{k}}=|g|\,.$
Lemma 5.2.
Let $G,H\in F_{k}$ be two elements. If $G,H$ does not commute, then
$\text{diam}(A_{G,F_{k}}\cap A_{H,F_{k}})<2\max\{[G]_{F_{k}},[H]_{F_{k}}\}\,.$
Proof.
Assume $[G]_{F_{k}}\geq[H]_{F_{k}}$, and without loss of generality, assume that $G=g$, and $H=h$, are cyclically reduced words. Assume that
$\text{diam}(A_{G,F_{k}}\cap A_{H,F_{k}})\geq 2\max\{[G]_{F_{k}},[H]_{F_{k}}\}\,.$
By replacing $g$ with $g^{-1}$ if necessary, we may assume that $g,h$ translates in the same directions on their axes.
Now the segment $[1,g^{2}]$, which is a sub-segment of $A_{G,F_{k}}$, lies entirely in $A_{H,F_{k}}$. Denote
$x=gh\,,$
$y=hg\,.$
Since $g\in A_{H,F_{k}}$, we have that $y\in A_{H,F_{k}}$ and
$d(1,y)=|g|+|h|\,.$
Since $|h|\leq|g|$, and $h$ and $g$ translates in the same direction, and $[1,g]$ is a sub-segment of $A_{H,F_{k}}$, we have that $h\in[1,g]$. In particular, $h\in A_{G,F_{k}}$, and we get that $x\in[1,g^{2}]$ and
$d(1,x)=|g|+|h|\,.$
Since $g,h$ translates in the same direction, then $x$ and $y$ lies on the same side of $1$ on the line $A_{H,F_{k}}$. Hence,
$x=y\,.$
Lemma 5.3.
Let $p,q,G,H$ be reduced words in the free group $F_{k}$, and suppose that we have the graphical equalities (with no cancellations)
$G=YgY^{-1}\,,$
$H=ZhZ^{-1}\,,$
for some $g,h,Y,Z\in F_{k}$ (reduced words), so that $g,h$ are cyclically reduced. Assume that there exists an integer $n_{0}>10$ so that neither $p$ nor $q$ contains a sub-word of the form $g^{\pm n_{0}},h^{\pm n_{0}}$.
Then, for all integers $r,s>110n_{0}$, if
$pG^{r}q^{-1}H^{-s}=1\,,$
then the elements $pGp^{-1}$ and $H$ commute.
Proof.
Without loss of generality, assume that
$|g|\geq|h|\,.$
According to the assumption, we have that
$pYg^{r}Y^{-1}q^{-1}Zh^{-s}Z^{-1}=1\,.$
Let $T$ be a tree of cancellation for the last equation. We distinguish $6$ special segments in the tree $T$; those $6$ segments that correspond to the $6$ elements
$p,\,Yg^{r}Y^{-1},\,q,\,Zh^{s}Z^{-1}\,,g^{r},\,h^{s}\,.$
Since neither $p$ nor $q$ contains a sub-word of the form $g^{\pm n_{0}},h^{\pm n_{0}}$, the tree $T$, topologically, consists of two disjoint segments corresponding to the elements $p,q$, that are connected by a non-degenerate segment $I$ (so that $I$ intersects each of them in exactly one point). Denote by $A_{1},A_{2}$ and $B_{1},B_{2}$, the initial and the terminal points of the segments corresponding to the elements $p$ and $q$ correspondingly (so that $p=[A_{1},A_{2}]$ and $q=[B_{1},B_{2}]$). Denote also $I=[C_{1},C_{2}]$, where $C_{1}$ is the point in $I\cap[A_{1},A_{2}]$, and $C_{2}$ is the point in $I\cap[B_{1},B_{2}]$.
Now Denote by $E$ and $F$ the middle points of the segments $[A_{1},B_{1}]$ and $[A_{2},B_{2}]$ (corresponding to the elements $Zh^{s}Z^{-1}$ and $Yg^{r}Y^{-1}$) respectively. Note that $I=[A_{1},B_{1}]\cap[A_{2},B_{2}]$, and that $E$ and $F$ must lie on $I$ (since neither $p$ nor $q$ contains a sub-word of the form $g^{\pm n_{0}},h^{\pm n_{0}}$). We define a metric on $T$ by declaring that every edge is of length $1$. A (compatible) directions are defined on the segments $[A_{1},B_{1}]$ and $[A_{2},B_{2}]$ by declaring that $A_{1}$ and $A_{2}$ are their left endpoints respectively.
We further define “coordinates” for the points in $T$ by denoting $D(x)$ and $U(x)$ the points of (directed) distance $x$ from the point $E$ and that lie on the segments $[A_{1},B_{1}]$ and $[A_{2},B_{2}]$ respectively, for all (possible) real $x$. Note that $D(x)=U(x)$ when $D(x)$ or $U(x)$ lie on $I$. In such a case, we denote the point $D(x)=U(x)$, simply by $x$.
Let $[L^{h},M^{h}]$ and $[L^{g},M^{g}]$ be the sub-segments of $[A_{1},B_{1}]$ and $[A_{2},B_{2}]$, corresponding to the elements $h^{s}$ and $g^{r}$ respectively. Since $|g|\geq|h|$, and $g,h$ are cyclically reduced words, it suffices to show that a sub-segment of length $2|g|$ of $[L^{g},M^{g}]$ matches a sub-segment of $[L^{h},M^{h}]$ (by 5.2). Assume by contradiction that this is not the case.
Let $R_{E}$ and $R_{F}$ be the reflections of the segments $[A_{1},B_{1}]$ and $[A_{2},B_{2}]$ around the points $E$ and $F$ respectively. We note that
$Lab(e_{1})=Lab(R_{E}(e_{1}))^{-1}\,,$
$Lab(e_{2})=Lab(R_{F}(e_{2}))^{-1}\,,$
for all edges $e_{1}$ and $e_{2}$ in the sets $[A_{1},L^{h}]\cup[M^{h},B_{1}]$ and $[A_{2},L^{g}]\cup[M^{g},B_{2}]$ respectively, where $Lab(e)$ denotes the label of the edge $e$ read from left to right.
Since neither $p$ nor $q$ contains a sub-word of the form $g^{\pm n_{0}}$, there must be a sub-segment $[L_{1}^{g},M_{1}^{g}]$ of $[L^{g},M^{g}]$ that lies entirely in $I$, and whose length is at least
$|[L_{1}^{g},M_{1}^{g}]|\geq 108n_{0}|g|\,.$
Hence, there exists a sub-segment $[L_{2}^{g},M_{2}^{g}]$ of $[L_{1}^{g},M_{1}^{g}]$ that lies entirely in $I$ on either the left hand side or the right hand side of the point $E$, and that is of length at least
$|[L_{2}^{g},M_{2}^{g}]|\geq 54n_{0}|g|\,.$
Without loss of generality, assume that $[L_{2}^{g},M_{2}^{g}]$ lies on the left hand side of the point $E$.
Since we have assumed that there is no sub-segment of length $2|g|$ of $[L^{g},M^{g}]$ that matches a sub-segment of $[L^{h},M^{h}]$, we deduce that the point $L^{h}$ lies in the sub-segment $[M_{2}^{g}-2|g|,E]$.
Thus, the sub-segment $[L_{2}^{g},M_{2}^{g}-2|g|]$ of $[L^{g},M^{g}]$ lies entirely in $I\cap[A_{1},L^{h}]$, and is of length at least
$|[L_{2}^{g},M_{2}^{g}-2|g|]|\geq 53n_{0}|g|\,.$
Let $[L_{3}^{g},M_{3}^{g}]$ be the longest sub-segment of $[L^{g},M^{g}]$ with this property. In particular, if $E$ lies in $[A_{1},M^{g}]$, then, $M_{3}^{g}=L^{h}$.
Then, since $[L_{3}^{g},M_{3}^{g}]\subset[A_{1},L^{h}]$, we have
$Lab([R_{E}(M_{3}^{g}),R_{E}(L_{3}^{g})])=Lab(R_{E}([L_{3}^{g},M_{3}^{g}]))=Lab([L_{3}^{g},M_{3}^{g}])^{-1}\,.$
Hence, since the label of the segment $Lab([L^{g},M^{g}])=g^{r}$ cannot contain $g^{-1}$ as a sub-word ($g\neq 1$), the segment
$[R_{E}(M_{3}^{g}),R_{E}(L_{3}^{g})]$
cannot match $[L^{g},M^{g}]$ in a sub-segment of length more than $2|g|$.
And since $q$ does not contain a sub-word of the form $g^{\pm n_{0}}$, the segment
$[R_{E}(M_{3}^{g}),R_{E}(L_{3}^{g})]$
cannot match the segment $[B_{1},B_{2}]$ (corresponding to $q$) in a sub-segment of length more than $(n_{0}+1)|g|$.
We deduce that the segment
$[L_{3}^{g},M_{3}^{g}]$
contains a sub-segment $J$ (that lies in $J\subset I\cap[A_{1},L^{h}]$), so that its reflection around $E$, $R_{E}(J)$, is a sub-segment of $I$ that lies entirely on the right hand side of $M^{g}$, and so that the length of $J$ equals
$|J|=50n_{0}|g|\,.$
Since $J$ lies entirely on the right hand side of $M^{h}$, this implies in particular that $M^{h}$ lies in $[E,C_{2}-50n_{0}|g|]\subset I$.
Now if $E$ lies in $[A_{1},M^{g}]$, then, as stated before, we have $M_{3}^{g}=L^{h}$. In particular, the segment $[L^{h}-10|g|,L^{h}]$, is a sub-segment of $[L^{g},M^{g}]$ and lies entirely in $I$. Hence, applying $R_{E}$ on $[L^{h}-10|g|,L^{h}]$, gives a sub-segment of $I$ that cannot match $[L^{g},M^{g}]$ in a sub-segment of length more than $2|g|$. But
$R_{E}([L^{h}-10|g|,L^{h}])=[M^{h},M^{h}+10|g|]\,.$
Thus, the point $M^{g}$ must lie in $[M^{h},M^{h}+2|g|]$.
In particular, we deduce that the point $F$ lies on the left hand side of the point $E$, and the distance between them is at least
$E-F\geq 50n_{0}|g|\,.$
Since $E-F\geq 50n_{0}|g|$, we get that for all $x\in I$, if $R_{E}(x)$ belongs to $I$, then $R_{F}\circ R_{E}(x)$ is a translation of $x$ by at least $100n_{0}|g|$ to the left on the segment $[A_{2},B_{2}]$. Similarly, for all $x\in I$, if $R_{F}(x)$ belongs to $I$, then $R_{E}\circ R_{F}(x)$ is a translation of $x$ by at least $100n_{0}|g|$ to the right on the segment $[A_{1},B_{1}]$.
Hence, there must exists an (unique) integer $n$, so that the segment $\left(R_{F}\circ R_{E}\right)^{n}(J)$ does not lie entirely in $I$, or the segment $R_{E}\circ\left(R_{F}\circ R_{E}\right)^{n}(J)$ does not lie entirely in $I$. Assume without loss of generality that $\left(R_{F}\circ R_{E}\right)^{n}(J)$ does not lie entirely in $I$. Since $p$ does not contain a sub-word of the form $g^{\pm n_{0}}$, the segment $\left(R_{F}\circ R_{E}\right)^{n}(J)$ cannot match the segment $[A_{1},A_{2}]$ (corresponding to $p$) in a sub-segment of length more than $(n_{0}+1)|g|$, and hence, the segment $\left(R_{F}\circ R_{E}\right)^{n}(J)$ admits a sub-segment $J^{\prime}$ that lies entirely in $I$ on the left hand side of $L^{h}$, and whose length is
$|J^{\prime}|=45n_{0}|g|\,.$
Since $q$ does not contain a sub-word of the form $g^{\pm n_{0}}$, the segment $R_{E}(J^{\prime})$ cannot match the segment $[B_{1},B_{2}]$ (corresponding to $q$) in more than $(n_{0}+1)|g|$. Hence, $R_{E}(J^{\prime})$ contains a sub-segment $J^{\prime\prime}$ that lies entirely in $I$ on the right hand side of $M^{g}$, and whose length is
$|J^{\prime\prime}|=40n_{0}|g|\,.$
But $R_{F}(J^{\prime\prime})$ lies entirely in the segment $[A_{2},C_{1}]$ which is a sub-segment of the segment $[A_{1},A_{2}]$ corresponding to $p$. Thus, $p$ contains a sub-word of the form $g^{-20n_{0}}$, a contradiction. ∎
Finally, we state the following simple fact, which will help us in bounding probabilities in what follows.
Lemma 5.4.
For every $a>1$ and every integers $n,k$, the function
$\frac{x^{n\ln^{k}x}}{a^{x}}$
approaches 00 as $x$ goes to $\infty$.
Proof.
We have that:
$\frac{x^{n\ln^{k}x}}{a^{x}}=\frac{e^{n\ln^{k+1}x}}{a^{x}}=\frac{c^{\ln^{k+1}x}}{a^{x}}\,,$
where $c=e^{n}$. Let $\epsilon>0$. Then, for $x>0$, we have
$\frac{c^{\ln^{k+1}x}}{a^{x}}<\epsilon$
$\iff$ $c^{\ln^{k+1}x}<\epsilon a^{x}$
$\iff$ $\ln^{k+1}x\cdot\ln c<\ln\epsilon+x\ln a$
$\iff$ $\frac{\ln^{k+1}x}{x}<\frac{\ln a}{\ln c}+\frac{\ln\epsilon}{x\ln c}\,.$
For all large enough $x$, we have $-1<\frac{\ln\epsilon}{x\ln c}$, and $\frac{\ln^{k+1}x}{x}<\frac{\ln a}{\ln c}-1$. ∎
6 Lifting Solutions of Bounded Lengths over Random Groups
Definition 6.1.
Let $y=(y_{1},...,y_{q})$ be a free basis, and let $\Sigma=\Sigma(y,a)$ be a system of equations with variables $y$ and with (or without) constants in the alphabet $a^{\pm}$. Let $\Gamma=\langle a:\mathcal{R}\rangle$ be a group presentation.
  • A tuple of elements $y_{0}\in\Gamma^{q}$, is called a solution for the system $\Sigma$ over the group $\Gamma$, if each of the equations defining $\Sigma$ is satisfied in $\Gamma$ when we evaluate $y=y_{0}$ in it. Briefly, in this case we write $y_{0}\in\Gamma$ and $\Sigma(y_{0},a)=1$ over $\Gamma$. We often treat $y_{0}$ as (a tuple of) element(s) of $F_{k}$, but in those cases, we intend that $y_{0}$ is a reduced word that represents a geodesic in $\Gamma$ (a geodesic word).
  • We define the circumference $C_{\Sigma}(y_{0})$ of a solution $y_{0}$ over the group $\Gamma$ of a system of equations $\Sigma(y,a)$, to be
[…]
    where $w$ runs on the equations in $\Sigma$, $y_{0}$ represent geodesic words in $\Gamma$, and $||w(y_{0})||$ is the length of the word $w(y_{0})$ in a free semi-group (without canceling).
  • The length $L_{\Sigma}(y_{0})$ of the solution $y_{0}$ over $\Gamma$, is defined to be the length of the longest specialization of the variables given by the tuple $y_{0}$ (the lengths measured in $\Gamma$).
  • The number of equations in the system $\Sigma$ is denoted by $N(\Sigma)$. And the length of the system $\Sigma(y,a)$, denoted $|\Sigma|$, is defined to be the sum of the lengths of the words defining the equations of $\Sigma$ (measured in a free group).
  • Since the group is given by presentation $\Gamma=\langle a:\mathcal{R}\rangle$, it is equipped with an epimorphism $\pi_{\Gamma}:F_{k}\rightarrow\Gamma$ (mapping $a$ to $a$ recall that $F_{k}$ denotes a free group with a fixed basis $a$). Let $y_{0}\in\Gamma$ be a solution of $\Sigma$ over $\Gamma$. We say that the solution $y_{0}=(y_{1}^{0},...,y_{q}^{0})$ can be lifted to a solution of $\Sigma$ in $F_{k}$, if for all $i=1,...,q$, there exists an element $\tilde{y}_{i}^{0}$ in the set $\pi_{\Gamma}^{-1}(y_{i}^{0})$ of preimages of $y_{i}^{0}$, so that $\tilde{y}_{0}=(\tilde{y}_{1}^{0},...,\tilde{y}_{q}^{0})$ is a solution for $\Sigma$ over $F_{k}$.
Remark 6.2.
Note that for every solution $y_{0}\in\Gamma$ of the system $\Sigma$, we have
$C_{\Sigma}(y_{0})\leq|\Sigma|\cdot L_{\Sigma}(y_{0})\,.$
Example 6.3.
Consider the equation $y^{2}=1$, and the group $\Gamma=\langle a:a_{1}^{2}\rangle$. The solution $y_{0}=a_{1}$ of the equation $y^{2}=1$ over $\Gamma$, cannot be lifted to a solution in $F_{k}$.
Theorem 6.4.
Let $d<\frac{1}{2}$. Let $\Sigma$ be a system of equations, and let $\Gamma$ be the random group of level $l$ and density $d<\frac{1}{2}$. We denote by $p_{l}$ the probability that $\Gamma$ admits a solution of length at most $l\ln^{2}l$, that cannot be lifted to a solution of $\Sigma$ in $F_{k}$.
Then
$p_{l}\overset{l\rightarrow\infty}{\longrightarrow}0\,.$
Proof.
Let $\mathcal{V}^{0}$ be the collection of all the reduced decorated families of topological Van-Kampen diagrams $DFTD$ with decoration $\Sigma$ and boundary length at most $|\Sigma|\cdot l\ln^{2}l$ (the boundary length of a family of diagrams is the sum of the lengths of the boundaries of the diagrams in that family), that $\Gamma$ could satisfy. For each such $DFTD$ in $\mathcal{V}^{0}$, we declare no rigid parts in the boundary, and the parts in the boundary corresponding to the constants appearing in $\Sigma$, are declared to be the constant parts of the boundary.
Since the boundary length of each of the diagrams in each of the decorated families in $\mathcal{V}^{0}$ is at most $|\Sigma|\cdot l\ln^{2}l$, we have according to 4.1, that the amount of cells $m$ in any diagram in $\mathcal{V}^{0}$ is bounded by $m\leq\beta_{0}\ln^{2}l$, where $\beta_{0}=\frac{|\Sigma|}{\frac{1}{2}-d}$ (see 4.2). Thus, according to 4.5, there exists a polynomial $Q=Q_{d}$, so that
$|\mathcal{V}^{0}|\leq Q(l)^{\ln^{4}l}\,.$
We generally reduce each of the decorated families in $\mathcal{V}^{0}$, and denote the collection of the obtained generally reduced decorated families by $\mathcal{V}$.
We denote by $\mathcal{V}_{t}$ the sub-collection of decorated families in $\mathcal{V}$ that contain no cells, i.e., those decorated families consisting of $N(\Sigma)$ trees (where $N(\Sigma)$ is the number of equations in $\Sigma$). And we denote by $\mathcal{V}_{1}$ the remaining decorated families in $\mathcal{V}$.
We denote by $\mathcal{V}_{t}^{0}$ and $\mathcal{V}_{1}^{0}$, the sub-collections consisting of those diagrams in $\mathcal{V}^{0}$ who admit an associated diagram in $\mathcal{V}_{t}$ and $\mathcal{V}_{1}$ respectively.
By the definition of the collection $\mathcal{V}_{1}$, every decorated family in $\mathcal{V}_{1}$ admits a positive amount $n\geq 1$ of distinct numberings, but contains no isolated numberings (since there are no rigid variables in the decoration). Hence, according to 3.28, for every decorated family $DFTD$ in $\mathcal{V}_{1}$, the probability that the random group $\Gamma$ of level $l$ and density $d$ fulfills $DFTD$ is at most
$\left(\frac{2k}{(2k-1)^{(\frac{1}{2}-d)l}}\right)^{n}\leq\frac{2k}{(2k-1)^{(\frac{1}{2}-d)l}}\,.$
Thus, we conclude that the probability that there exists some decorated family $DFTD$ in the collection $\mathcal{V}_{1}$ so that $\Gamma$ fulfills $DFTD$, is at most
$|\mathcal{V}_{1}|\cdot\frac{2k}{(2k-1)^{(\frac{1}{2}-d)l}}\leq|\mathcal{V}^{0}|\cdot\frac{2k}{(2k-1)^{(\frac{1}{2}-d)l}}\leq\frac{2k\cdot Q(l)^{\ln^{4}l}}{(2k-1)^{(\frac{1}{2}-d)l}}\,\overset{l\rightarrow\infty}{\longrightarrow}0$
which, according to 5.4, approaches 00 as $l$ goes to infinity.
Recall according to 3.26 that general reduction can only increase the probability of fulfilling the decorated family. Hence, we conclude that the probability $p_{1}(l)$ that there exists some decorated family $DF$ in the collection $\mathcal{V}_{1}^{0}$ so that the random group $\Gamma$ of level $l$ and density $d$ fulfills $DF$, approaches 00 as $l$ goes to infinity
$p_{1}(l)\overset{l\rightarrow\infty}{\longrightarrow}0\,.$
Hence, we may assume that the random group $\Gamma$ of level $l$ and density $d$, does not fulfill any of the diagrams in the collection $\mathcal{V}_{1}^{0}$.
Finally, assume that $\Gamma$ admits a solution $y_{0}$ of length (over $\Gamma$) at most $L_{\Sigma}(y_{0})\leq l\ln^{2}l$ for the system $\Sigma$. We need to prove that $y_{0}$ admits a lift $\tilde{y}_{0}\in F_{k}$, so that $\tilde{y}_{0}$ is a solution of $\Sigma$ in $F_{k}$.
Since $\Sigma(y_{0})=1$ in $\Gamma$, and $L_{\Sigma}(y_{0})\leq l\ln^{2}l$, the group $\Gamma$ must fulfill some reduced decorated family $DFTD$ with decoration $\Sigma$ in the collection $\mathcal{V}^{0}$, so that the obtained family $FVKD$ of Van-Kampen diagrams consists of reduced diagrams over $\Gamma$, and their contours are given by $\Sigma(y_{0})$ (i.e., the contour of the $i$-th diagram is $w_{i}(y_{0})$, where $w_{i}$ is the $i$-th word in $\Sigma$, for all $i=1,...,N(\Sigma)$).
However, since we have already proved that the collection of groups of level $l$ and density $d$ that satisfy a decorated family in $\mathcal{V}_{1}^{0}$ is negligible, we can assume that $DFTD$ belongs to the collection $\mathcal{V}_{ab}^{0}$.
Since the decorated family $DFTD$ of topological diagrams, by the definition of $\mathcal{V}_{ab}^{0}$, can be converted to a tree by means of applying topological modifications (general reduction) on its diagrams, the family $FVKD$ of Van-Kampen diagrams can be converted to a tree $T$ by applying the same sequence of modifications on its diagrams. However, these modifications do not change the values (in $\Gamma$) of the variables along the boundary.
Hence, the procedure of applying those modifications on $FVKD$, provides lifts $\tilde{y}_{0}\in F_{k}$ for the elements $y_{0}\in\Gamma$, so that
$\Sigma(\tilde{y}_{0})=1$
in $F_{k}$. As claimed. ∎
Definition 6.5.
Let $\Sigma(y,a)$ be a system of equations, and let $l$ be an integer. Let $\pi:F_{k}\rightarrow\Gamma$ be a group presentation. We say that $\Gamma$ satisfies the $\Sigma(y,a)$-$lln^{2}$ lifting property, or briefly $\Sigma$-$lln^{2}$ l.p., if every solution of $\Sigma$ of length at most $l\ln^{2}l$ in $\Gamma$, can be lifted to a solution of $\Sigma$ in $F_{k}$.
Theorem 6.6.
Let $\mathcal{H}$ be the collection of all finite systems of equations. Let $d<\frac{1}{2}$. Then, we can order the systems in $\mathcal{H}$ in an ascending sub-collections $H_{1}\subset H_{2}\subset H_{3}\subset...$, so that the following property is satisfied.
For all $l$, let $\Gamma_{l}$ be the random group of level $l$ and density $d$. Then, If $p_{l}$ is the probability that the random group $\Gamma_{l}$ satisfies the $\Sigma$-$lln^{2}$ l.p. for all $\Sigma\in H_{l}$, then $p_{l}$ converges to $1$ as $l$ approaches $\infty$.
Proof.
Let $q_{l}=1-p_{l}$. Of course, assuming that $H_{l}$ was defined, $q_{l}$ equals the probability that there exists some system $\Sigma$ in $H_{l}$, so that the random group of level $l$, $\Gamma_{l}$, does not satisfy the $\Sigma$-$lln^{2}$ l.p..
Given a system $\Sigma\in\mathcal{H}$, we denote by $q_{l}^{\Sigma}$, the probability that $\Gamma_{l}$ does not satisfy the $\Sigma$-$lln^{2}$ l.p.. Fixing a system $\Sigma$, we already know, according to 6.4, that $q_{l}^{\Sigma}\rightarrow 0$ as $l\rightarrow\infty$.
Since there exists only a countable amount of finite systems, we can write them in a sequence $\Sigma_{1},\Sigma_{2},\Sigma_{3},...$. Now let $1=l_{1}\leq l_{2}\leq l_{3},...$ be an ascending sequence of integers, so that for all $n\in\mathbb{N}$, we have that $q_{l}^{\Sigma_{1}}+q_{l}^{\Sigma_{2}}...+q_{l}^{\Sigma_{n}}<\frac{1}{n}$ for all $l\geq l_{n}$.
For every integer $l$, let $n$ be the maximal integer for which $l_{n}\leq l$, and we define $H_{l}$ to be $H_{l}=\{\Sigma_{1},...,\Sigma_{n}\}$. Of course, the probability $q_{l}$ that there exists some system $\Sigma$ in $H_{l}$, so that the random group of level $l$, $\Gamma_{l}$, does not satisfy the $\Sigma$-$lln^{2}$ l.p., satisfies $q_{l}\leq q_{l}^{\Sigma_{1}}+q_{l}^{\Sigma_{2}}...+q_{l}^{\Sigma_{n}}<\frac{1}{n}\leq\frac{1}{l}$.
Hence $q_{l}\rightarrow 0$ as $l\rightarrow\infty$. As required. ∎
7 Axes of Elements in Random Groups
We recall according to 4.4 that for a fixed $d<\frac{1}{2}$, the random group of level $l$ and density $d$ can be assumed to be torsion-free $\alpha_{0}l$-hyperbolic, where
$\alpha_{0}=\frac{48}{(1-2d)^{2}}\,.$
In what follows, we will use this fact extensively, and hence without explicit reference. We use the abbreviation nbhd for neighborhood.
7.1 Conjugates of Small Elements in Random Groups
Lemma 7.1.
Let $d<\frac{1}{2}$. Let $\Gamma$ be the random group of level $l$ and density $d$.
Then, with overwhelming probability, if $u,v\in\Gamma$ are two conjugate elements, then there exists an element $z\in\Gamma$ in the ball of radius $|u|+|v|+(8\alpha_{0}+\zeta_{0})l$, so that
$u=zvz^{-1}\,,$
where $\zeta_{0}=\frac{100\alpha_{0}}{(\frac{1}{2}-d)}$.
Proof.
Suppose $U,V,Z\in F_{k}$ are geodesic words, so that $U=ZVZ^{-1}$ in $\Gamma$, and $Z$ is of minimal length with this property. Assume by contradiction that
$|Z|>|U|+|V|+8\alpha_{0}l+\zeta_{0}l\,.$
Consider a geodesic rectangle $\Pi$ in the Cayley graph of $\Gamma$ that corresponds to the equation $U=ZVZ^{-1}$. Denote by $p_{1},p_{2}$ and $q_{1},q_{2}$ the initial and terminal points of the sides of $\Pi$ corresponding to $U$ and $V$ respectively.
Since we have assumed that $|Z|>|U|+|V|+8\alpha_{0}l+\zeta_{0}l$, and since the sides of $\Pi$ are geodesics, there must exist two points $x,y$ on the side $p_{1}q_{1}$ of $\Pi$, so that $x$ lies between $p_{1}$ and $y$ say, neither $x$ nor $y$ belongs to the $2\alpha_{0}l$-nbhd of $p_{1}p_{2}\cup q_{1}q_{2}$, and
$d(x,y)=\zeta_{0}l+1\,.$
Let $x^{\prime},y^{\prime}$ be the dual points of $x,y$ respectively that lie on the side $p_{2}q_{2}$ of $\Pi$. That is, the points $x^{\prime},y^{\prime}$ are those points on $p_{2}q_{2}$ for which
$d(p_{2},x^{\prime})=d(p_{1},x)\,,$
$d(q_{2},y^{\prime})=d(q_{1},y)\,.$
We want to show now that $d(y,y^{\prime}),\>d(x,x^{\prime})\leq 10\alpha_{0}l$. By the $\alpha_{0}l$-hyperbolicity of $\Gamma$, there exist two points $z_{x},z_{y}$ on the side $p_{2}q_{2}$ of $\Pi$, so that
$d(x,z_{x}),\,d(y,z_{y})\leq 2\alpha_{0}l\,.$
If, by contradiction, $d(x^{\prime},z_{x})>2\alpha_{0}l$, then, in particular, $x^{\prime}$ lies between $p_{2}$ and $z_{x}$, or $z_{x}$ lies between $p_{2}$ and $x^{\prime}$. Assume without loss of generality the first case, and consider the element $Z^{\prime}$ of $\Gamma$ that labels the path $p_{1}xz_{x}q_{2}$. Then, $Z^{\prime}$ is of length at most
$|Z^{\prime}|\leq d(p_{1},x)+d(x,z_{x})+d(z_{x},q_{2})\leq|Z|-(d(x^{\prime},z_{x})-2\alpha_{0}l)<|Z|\,.$
However, we have that $Z^{\prime}=ZV$ in $\Gamma$, and hence
$Z^{\prime}VZ^{\prime-1}=ZVZ^{-1}=U\,,$
in $\Gamma$, which contradicts the minimality of (the length of) $Z$ with this property. Thus, we conclude that
$d(x,x^{\prime}),\>d(y,y^{\prime})\leq 10\alpha_{0}l\,.$
Now consider the three geodesic words $u$, $v$, and $z$ that label geodesic segments $xx^{\prime}$, $yy^{\prime}$, and $xy$ respectively in the Cayley graph of $\Gamma$. By the construction of the dual points $x^{\prime}$ and $y^{\prime}$ of the points $x$ and $y$, we know that
$u=zvz^{-1}$
in $\Gamma$. Moreover, since
$d(x,y)=|z|=\zeta_{0}l+1\,,$
if we show that the elements $u,v$ can be conjugated by an element $z^{\prime}\in\Gamma$ of length at most
$|z^{\prime}|\leq\zeta_{0}l\,,$
then we conclude that the original elements $U,V$ can be conjugated by an element of length less that $|Z|$, which is a contradiction that implies the claim.
Indeed, we consider the decoration $\hat{u}\hat{z}\hat{v}^{-1}\hat{z}^{-1}$, and declare the the rigid variables to be $\hat{u},\hat{v}$, and no constant variables. Let $\mathcal{V}^{0}$ be the collection of all the reduced decorated topological Van-Kampen diagrams $DTD$ with decoration $\hat{u}\hat{z}\hat{v}^{-1}\hat{z}^{-1}$ and boundary length at most $l\ln l$, that $\Gamma$ could satisfy, and so that the lengths of the parts of the boundary of $DTD$ corresponding to the variables $\hat{u},\hat{v}$, sum to at most $2\nu_{0}l$, where $\nu_{0}=10\alpha_{0}$.
Since the boundary length of each of the diagrams in $\mathcal{V}^{0}$ is at most $l\ln l$, we have according to 4.1, that the amount of cells $m$ in any diagram in $\mathcal{V}^{0}$ is bounded by $m\leq\beta_{0}\ln l$, where $\beta_{0}=\frac{1}{\frac{1}{2}-d}$ (see 4.2). Thus, according to 4.5, there exists a polynomial $Q=Q_{d}$, so that
$|\mathcal{V}^{0}|\leq Q(l)^{\ln^{2}l}\,.$
We generally reduce each of the diagrams in $\mathcal{V}^{0}$, and denote the collection of the obtained generally reduced diagrams by $\mathcal{V}$. Recall according to 3.26 that general reduction can only increase the probability of fulfilling the diagram.
For a given generally reduced decorated topological Van-Kampen diagram $DTD$ in $\mathcal{V}$, with $m$ cells numbered by $1,...,n$, for some $1\leq n\leq m$, we describe the following iterative procedure. If $DTD$ contains an isolated numbering, we apply a mining modification on $DTD$. If the obtained diagram, denoted $DTD$ again, contains an isolated numbering, we apply another mining modification. We continue iteratively, until we arrive to a generally reduced decorated diagram $DTD$ with no isolated numberings, or we apply at most
$k_{0}=\left\lceil\frac{\nu_{0}}{\frac{1}{2}-d}\right\rceil+1\,.$
iterative mining modifications. Note that $k_{0}=k_{0,d}$ depends on $d$ only.
We denote by $\mathcal{V}_{1}$ the sub-collection of $\mathcal{V}$ consisting of those diagrams $DTD$ in $\mathcal{V}$ for which the procedure performs $k_{0}$ iterations.
We denote by $\mathcal{V}_{F}$ the sub-collection of $\mathcal{V}\backslash\mathcal{V}_{1}$ consisting of those diagrams $DTD$ in $\mathcal{V}$ for which the procedure stops (before applying $k_{0}$ iterations) in a diagram with no cells (a tree).
And we denote by $\mathcal{V}_{2}$ the remaining diagrams of $\mathcal{V}$; i.e., $\mathcal{V}_{2}=\mathcal{V}\backslash\mathcal{V}_{1}\cup\mathcal{V}_{F}$ consists of those diagrams in $\mathcal{V}$ for which the procedure stops (before applying $k_{0}$ iterations) in a diagram with a positive number of cells. We replace the diagrams in $\mathcal{V}_{2}$ with the diagrams obtained by applying the procedure on them.
We denote by $\mathcal{V}_{1}^{0},\mathcal{V}_{F}^{0},\mathcal{V}_{2}^{0}$ the sub-collections consisting of those diagrams in $\mathcal{V}^{0}$ who admit an associated diagram in $\mathcal{V}_{1},\mathcal{V}_{F},\mathcal{V}_{2}$ respectively.
By the definition of the procedure above, and according to 3.25, the amount of distinct numberings $n$ in any diagram in the collection $\mathcal{V}_{1}$, is at least
$n\geq k_{0}=\left\lceil\frac{\nu_{0}}{\frac{1}{2}-d}\right\rceil+1\,.$
Hence, according to 3.28, for every diagram $DTD$ in $\mathcal{V}_{1}$ with $n$ distinct numberings, the probability that the random group $\Gamma$ of level $l$ and density $d$ fulfills $DTD$ is at most
$\left(\frac{2k}{(2k-1)^{(\frac{1}{2}-d-\frac{\nu_{0}}{n})l}}\right)^{n}\leq\frac{2k}{(2k-1)^{\gamma_{0}l}}\,,$
where $\gamma_{0}=\frac{1}{2}-d-\frac{\nu_{0}}{k_{0}}>0$ (recall that the rigid parts in $DTD$ sum to at most $2\nu_{0}l$).
On the other hand, by the definition of the collection $\mathcal{V}_{2}$, every diagram in $\mathcal{V}_{2}$ admits a positive amount $n\geq 1$ of distinct numberings, but contains no isolated numberings. Hence, according to 3.28, for every diagram $DTD$ in $\mathcal{V}_{2}$, the probability that $\Gamma$ fulfills $DTD$ is at most
$\left(\frac{2k}{(2k-1)^{(\frac{1}{2}-d)l}}\right)^{n}\leq\frac{2k}{(2k-1)^{\gamma_{0}l}}\,.$
Thus, we conclude that the probability that there exists some diagram $DTD$ in the collection $\mathcal{V}_{1}\cup\mathcal{V}_{2}$ so that the random group $\Gamma$ of level $l$ and density $d$ fulfills $DTD$, is at most
$|\mathcal{V}_{1}\cup\mathcal{V}_{2}|\cdot\frac{2k}{(2k-1)^{\gamma_{0}l}}\leq|\mathcal{V}^{0}|\cdot\frac{2k}{(2k-1)^{\gamma_{0}l}}\leq\frac{2k\cdot Q(l)^{\ln^{2}l}}{(2k-1)^{\gamma_{0}l}}\,\overset{l\rightarrow\infty}{\longrightarrow}0$
which, according to 5.4, approaches 00 as $l$ goes to infinity.
Recall according to 3.25 that mining modifications can only increase the probability of fulfilling the diagram. Hence, we conclude that the probability $p_{1,2}(l)$ that there exists some diagram $D$ in the collection $\mathcal{V}_{1}^{0}\cup\mathcal{V}_{2}^{0}$ so that the random group $\Gamma$ of level $l$ and density $d$ fulfills $D$, approaches 00 as $l$ goes to infinity
$p_{1,2}(l)\overset{l\rightarrow\infty}{\longrightarrow}0\,.$
Hence we may assume that the random group $\Gamma$ of level $l$ and density $d$, does not fulfill any of the diagrams in the collection $\mathcal{V}_{1}^{0}\cup\mathcal{V}_{2}^{0}$.
Now recall our elements $u$, $v$, and $z$ who satisfy
$uzv^{-1}z^{-1}=1$
in $\Gamma$, explained in the top of the proof. Since $|u|,|v|\leq\nu_{0}l$, and since $|z|\leq l\ln l$, we get that the group $\Gamma$ must fulfill some reduced decorated Van-Kampen diagram $DTD$ with decoration $\hat{u}\hat{z}\hat{v}^{-1}\hat{z}^{-1}$ in the collection $\mathcal{V}^{0}$, so that the contour of the obtained (reduced) Van-Kampen diagram $VKD$ is $uzv^{-1}z^{-1}$.
However, since we have already proved that the collection of groups of level $l$ and density $d$ that satisfy a diagram in $\mathcal{V}_{1}^{0}\cup\mathcal{V}_{2}^{0}$ is negligible, we can assume that $DTD$ belongs to the collection $\mathcal{V}_{F}^{0}$. Since the topological diagram $DTD$, by the definition of $\mathcal{V}_{F}^{0}$, can be converted to a tree by means of applying topological modifications (mining and general reduction) on it, the Van-Kampen diagram $VKD$ can be converted to a tree $T$ by applying the same sequence of modifications on it. However, these modifications do not change the values (in $\Gamma$) of the variables along the boundary.
Hence, the procedure of applying those modifications on $VKD$, provides lifts $\tilde{u},\tilde{v},\tilde{z}\in F_{k}$ (reduced words - the obtained words on the boundary of $T$ could be non-reduced, but we simply reduce them) for the elements $u,v,z\in\Gamma$ respectively, so that
$\tilde{u}\tilde{z}\tilde{v}^{-1}\tilde{z}^{-1}=1$
in the free group $F_{k}$. However, according to the definition of $\mathcal{V}_{F}^{0}$, the tree $T$ is obtained by applying a sequence of alternate general reduction and mining modifications on $VKD$, applied for at most $k_{0}=k_{0,d}$ iterations. Now, by definition, general reduction can change the rigid parts $u,v$ of the boundary, only by removing self-intersections from them, but a single mining modification can change a rigid part by eliminating only cells that share some edge with that part. Thus, the lifts $\tilde{u},\tilde{v}$ that are obtained as a consequence of applying at most $k_{0}=k_{0,d}$ mining modifications, lie, as paths in the Cayley graphof $\Gamma$, in the $k_{0}l$-nbhd of the geodesic segments $u,v$ in the Cayley graph of $\Gamma$, respectively.
In particular, each of the lifts $\tilde{u}$ and $\tilde{v}$, as paths in the Cayley graph of $\Gamma$, is contained in a ball of radius
$\frac{|u|}{2}+k_{0}l,\,\frac{|v|}{2}+k_{0}l\leq 2k_{0}l\,.$
Finally, we consider the obtained lifts $\tilde{u},\tilde{v}\in F_{k}$ for the elements $g,h\in\Gamma$. Write $\tilde{u}=A^{\prime}eA^{\prime-1}$, and $\tilde{v}=B^{\prime}fB^{\prime-1}$ (graphical equality), for reduced words $A^{\prime},B^{\prime}\in F_{k}$ and $e,f\in F_{k}$ cyclically reduced. Then, since $\tilde{u}$ and $\tilde{v}$ are conjugates in the free group $F_{k}$, the words $e$ and $f$ must be cyclic permutations to each other. Write $f=f_{1}f_{2}$ (graphical equality), so that $e=f_{2}f_{1}$. Then, we have
$\tilde{u}=A^{\prime}(B^{\prime}f_{1})^{-1}\tilde{v}(A^{\prime}(B^{\prime}f_{1})^{-1})^{-1}\,,$
which translates in $\Gamma$ as the equality
$u=AB^{-1}v(AB^{-1})^{-1}\,,$
where $A$ and $B$ denote geodesic words with $A=A^{\prime}$ and $B=B^{\prime}f_{1}$ in $\Gamma$. However, as explained above, a path labeled $\tilde{u}=A^{\prime}eA^{\prime-1}$ or $\tilde{v}=B^{\prime}f_{1}f_{2}B^{\prime-1}$ in the Cayley graph of $\Gamma$, stays inside a ball of radius $2k_{0}l$. In particular, we have that the length of each of the elements $A,B$ is at most
$|A|,|B|\leq 4k_{0}l\,.$
Hence, the length of the element $z^{\prime}=AB^{-1}$ is at most
$|z^{\prime}|\leq 8k_{0}l\leq\zeta_{0}l\,.$
As required. ∎
7.2 Isometries of a Hyperbolic Space
The readers who are not familiar with hyperbolic geometry in the context of group theory, are recommended to see the first section in [Co].
Definition 7.2.
Let $\Gamma=\langle a:\mathcal{R}\rangle$ be a torsion-free non-elementary hyperbolic group, and let $X$ be its Cayley graph (w.r.t. the generating set $a$). Let $G$ be an element of $\Gamma$.
  1. 1.
    The translation length of $G$, denoted by $[G]_{X}$, or simply $[G]$, is defined to be
    $[G]=\underset{x\in X}{\min}d(Gx,x)\,.$
  2. 2.
    Let $x\in X$. The asymptotic translation length of $g$, denoted by $[G]_{X}^{\infty}$, or simply $[G]^{\infty}$, is defined to be
    $[G]^{\infty}=\underset{n\rightarrow\infty}{lim}\frac{d(G^{n}x,x)}{n}\,.$
Note that the definition of the asymptotic translation length does not depend on the choice of the point $x$ (see [Br-Ha], Chapter II.6, exercises 6.6).
We denote by $\partial X$ the Gromov boundary of $X$ (see [Br-Ha], Chapter III.H). Let $G$ be a non-trivial element of $\Gamma$. Each one of the orbits $\{G^{n}x:n\in\mathbb{Z},\ n\geq 0\}$ and $\{G^{n}x:n\in\mathbb{Z},\ n\leq 0\}$ has exactly one accumulation point in $\partial X$. These two points are distinct and we denote them by $G_{+}$ and $G_{-}$ respectively. $G_{+}$ and $G_{-}$ do not depend on the choice of $x$. They are the only points of $\partial X$ fixed by the action of $G$.
Lemma 7.3.
Let $\Gamma=\langle a:\mathcal{R}\rangle$ be a torsion-free non-elementary $\delta$-hyperbolic group, and let $X$ be its Cayley graph (w.r.t. the generating set $a$). Let $G$ be an element of $\Gamma$. Let $k\in\mathbb{Z}$.
Then
  1. 1.
    $[G]^{\infty}\leq[G]\leq[G]^{\infty}+32\delta$;
  2. 2.
    $[G^{k}]\leq|k|\cdot[G]$;
  3. 3.
    $[G^{k}]^{\infty}=|k|\cdot[G]^{\infty}$.
Proof.
For Point 1, see [[Co], p.14]. Points 2&3 are a consequence of triangle inequality, and that the sequence $(\frac{d(G^{kn}x,x)}{kn})_{n}$ is a subsequence of $(\frac{d(G^{n}x,x)}{n})$ (for $k>0$). ∎
7.3 Axes of Elements in Random Groups
Definition 7.4.
Let $\Gamma=\langle a:\mathcal{R}\rangle$ be a hyperbolic group, and let $G\in\Gamma$ be a hyperbolic element. Let $Z,g\in F_{k}$ be two geodesic words in $\Gamma$ so that $G=ZgZ^{-1}$, and the length of $g$ equals the translation length of $G$ in $\Gamma$; $|g|=[G]$. Let $X$ be the Cayley graph of $\Gamma$ (w.r.t. the generating set $a$).
  • The path in the Cayley graph of $\Gamma$ induced by the elements $Zg^{n}$, $n\in\mathbb{Z}$, is called a nerve for $G$. The graphical word $g$, as an element of $F_{k}$, is called the word of the nerve. A nerve of $G$ is denoted usually by $\gamma_{G}$.
  • An axis of $G$ is any geodesic connecting the points at infinity $G_{-}$ and $G_{+}$. An axis of $G$ is usually denoted by $A_{G}$.
  • Given a real number $C\geq 0$, and a subset $U\subset X$, we denote the closed $C$-neighborhood of $U$ by
    $U^{+C}=\{x\in X:d(x,U)\leq C\}\,.$
  • If $X$ is $\delta$-hyperbolic, we define the parameter
    $D(\Gamma,X)=\{A_{G}^{+60\delta}\cap A_{H}^{+60\delta}:\,[G,H]\neq 1,\,[G],[H]\leq 1000\delta\}\,.$
    This parameter will play an important role when considering limit actions over an ascending sequences of random groups.
Remark 7.5.
  1. 1.
    A nerve $\gamma_{G}$ of a hyperbolic element $G\in\Gamma$, is a $[G]$-local geodesic. To see this, Let $Z,g$ be as in 7.4. If, by contradiction, a cyclic shift $g^{\prime}$ of $g$ is not a geodesic word in $\Gamma$, then $[g]=[g^{\prime}]\leq|g^{\prime}|<|g|$, but $[g]=[G]=|g|$, a contradiction.
  2. 2.
    If $A_{G}$ is an axis for $G$, then $GA_{G}$ is an axis for $G$ as well. In general, if $H,G$ are hyperbolic elements of $\Gamma$, and $A_{G}$ and $\gamma_{G}$ are an axis and a nerve with word $w$ for $G$, then $HA_{G}$ is an axis for the element $HGH^{-1}$, and $H\gamma_{G}$ is a nerve for $HGH^{-1}$ with word $w$.
Lemma 7.6.
([Br-Ha], Theorem 1.13) Let $\Gamma=\langle a:\mathcal{R}\rangle$ be a $\delta$-hyperbolic group, and let $X$ be its Cayley graph.
Then, the Hausdorff distance between any two $200\delta$-local geodesics connecting the same two endpoints in the boundary of $X$, $\partial X$, is at most $5\delta$. In particular, if $G\in\Gamma$ is a hyperbolic element with $[G]\geq 200\delta$, then the $5\delta$-nbhd of any nerve or axis of $G$, contains all the nerves and all the axes of $G$.
Lemma 7.7.
([Xi]) Let $\Gamma=\langle a:\mathcal{R}\rangle$ be a $\delta$-hyperbolic group, and let $X$ be its Cayley graph. Let $G\in\Gamma$ be a hyperbolic element, and let $c$ be an axis of $G$. For all $y\in X$, let $P_{c}(y)$ be a projection of $y$ on $c$ (i.e., $P_{c}(y)$ is a closest point on $c$ to $y$).
If $x\in c$ and $d(x,Gx)\geq 20\delta$, then, for all integers $0<i<j$, the point $P_{c}(G^{j}x)$ lies between $P_{c}(G^{i}x)$ and $G_{+}$.
Corollary 7.8.
Let $d<\frac{1}{2}$, and let $\Gamma$ be the random group of level $l$ and density $d$.
With overwhelming probability, every two nerves of any two commuting elements $G_{1},G_{2}\in\Gamma$ are of Hausdorff distance at most $2\cdot(3350\alpha_{0}+\zeta_{0})l$. And hence, every nerve and every axis of a hyperbolic element, are of Hausdorff distance at most $\rho_{0}l$, where $\rho_{0}=2\cdot(3355\alpha_{0}+\zeta_{0})$ (and $\zeta_{0}$ is the one given in 7.1).
Proof.
According to 7.6, it suffices to assume that $[G_{1}]=[G]<200\alpha_{0}l$, and to show then that for every nerve $\gamma_{G}$ of $G$, there exists an element $H\in\Gamma$ that commutes with $G$ and with translation length $[H]\geq 200\alpha_{0}l$, and there exists a nerve $\gamma_{H}$ for $H$, so that the Hausdorff distance between $\gamma_{G}$ and $\gamma_{H}$ is at most $(3200\alpha_{0}+\zeta_{0})l$.
Indeed, Without loss of generality, assume that $G=g$ is a geodesic word with $|g|=[G]<200\alpha_{0}l$, and let $\gamma_{G}$ be a nerve of $G$. Let $k$ be a large enough integer so that the translation length of the element $H=g^{k}$ satisfies $200\alpha_{0}l\leq[H]\leq 400\alpha_{0}l$. Let $Z,w$ be geodesic words so that $H=ZwZ^{-1}$, and $|w|=[H]$. Denote by $\gamma_{H}^{1}$ the nerve of $H$ induced by the elements $Zw^{n}$, $n\in\mathbb{Z}$. Then, since $[H]\leq 400\alpha_{0}l$, and $Z$ lies on $\gamma_{H}$, we deduce using 7.6 and 7.7 that $d(Z,gZ)\leq 450\alpha_{0}l$. In other words, if we denote $w^{\prime}=Z^{-1}gZ$, then $|w^{\prime}|\leq 450\alpha_{0}l$. Note that $w^{\prime k}=w$.
Hence, according to 7.1, there exists some geodesic word $z$ of length $|z|\leq(2000\alpha_{0}+\zeta_{0})l$ so that $g=zw^{\prime}z^{-1}$. Consider now the nerve $\gamma_{H}^{2}$ induced by the elements $zw^{n}$, $n\in\mathbb{Z}$, and the path $\gamma_{G}^{2}$ induced by the elements $zw^{\prime n}$, $n\in\mathbb{Z}$. Note that the Hausdorff distance between $\gamma_{G}^{1}=\gamma_{G}$ and $\gamma_{G}^{2}$ is at most $[g]+[w^{\prime}]+|z|\leq(2800\alpha_{0}+\zeta_{0})l$. Hence, it remains to show that $\gamma_{G}^{2}$ and $\gamma_{H}^{2}$ are of Hausdorff distance at most $500\alpha_{0}l$.
Indeed, since $[H]\geq 200\alpha_{0}l$, we have that $\gamma_{H}^{2}$ is of Hausdorff distance at most $5\alpha_{0}l$ from an axis $c$ of $g$. Hence, for all $n$, the path $g^{n}\gamma_{H}^{2}$ is of Hausdorff distance at most $5\alpha_{0}l$ from the axis $g^{n}c$ of $g$. Hence, the path $g^{n}\gamma_{H}^{2}$ is of Hausdorff distance at most $10\alpha_{0}l$ from $\gamma_{H}^{2}$. In particular, $d(zw^{\prime n},\gamma_{H}^{2})=d(g^{n}z,\gamma_{H}^{2})\leq 10\alpha_{0}l$. Thus, we deduce that $\gamma_{G}^{2}$ is contained in the $(10\alpha_{0}l+[w^{\prime}])$-nbhd of $\gamma_{H}^{2}$. On the other hand, of course, $\gamma_{H}^{2}$ is contained in the $[w]$-nbhd of $\gamma_{G}^{2}$. As required. ∎
Lemma 7.9.
Let $d<\frac{1}{2}$, and let $\Gamma$ be the random group of level $l$ and density $d$. Let $X$ be the Cayley graph of $\Gamma$ (w.r.t. the fixed generating set $a$).
Denote by $p_{l}$ the probability that $D(\Gamma,X)\geq 10l\cdot\ln l$. Then
$p_{l}\overset{l\rightarrow\infty}{\longrightarrow}0\,.$
Proof.
Assume that $G,H\in\Gamma$ so that $[G],[H]\leq 1000l$ and
$\text{diam}(A_{G}^{+60\alpha_{0}l}\cap A_{H}^{+60\alpha_{0}l})\geq 10l\cdot\ln l\,,$
where $A_{G}$ and $A_{H}$ are two axes of $G$ and $H$ respectively. According to 7.6, we can assume that
$200\alpha_{0}l\leq[G],[H]\leq 1000\alpha_{0}l\,,$
and that
$\text{diam}(\gamma_{G}^{+65\alpha_{0}l}\cap\gamma_{H}^{+65\alpha_{0}l})\geq 10l\cdot\ln l\,,$
where $\gamma_{G}$ and $\gamma_{H}$ are two nerves for $G$ and $H$ respectively. Write
$G=YgY^{-1},\qquad H=ZhZ^{-1}\,,$
so that $Y,Z,g,h\in F_{k}$ are geodesic words, and the words $g,h$ are the words of the nerves $\gamma_{G}$ and $\gamma_{H}$ respectively.
Using the $\alpha_{0}l$-hyperbolicity of $\Gamma$, we can find two pairs of points $p_{1},p_{2}$ and $q_{1},q_{2}$, in the Cayley graph of $\Gamma$, representing elements of the form $Zh^{s_{1}},Zh^{s_{2}}$ and $Yg^{r_{1}},Yg^{r_{2}}$ respectively, $r_{i},s_{i}\in\mathbb{Z}$, so that
$d(p_{1},p_{2}),\,d(q_{1},q_{2})\geq 9l\ln l\,,$
$d(p_{1},q_{1}),\,d(p_{2},q_{2})\leq\nu_{0}l\,,$
where $\nu_{0}=1130\alpha_{0}$. Without loss of generality, we assume that $r_{2}-r_{1},\,s_{2}-s_{1}\geq 0$ (if not, we replace $G$ or $H$ by $G^{-1}$ or $H^{-1}$).
We consider two geodesic words $x,y\in F_{k}$, so that
$x=p_{1}^{-1}q_{1}\,,$
$y=p_{2}^{-1}q_{2}\,.$
That is, $x,y$ labels geodesics $[p_{1},q_{1}],\,[p_{2},q_{2}]$ in the Cayley graph of $\Gamma$.
Under the assumption that
$\text{diam}(A_{G}^{+60\alpha_{0}l}\cap A_{H}^{+60\alpha_{0}l})\geq 10l\cdot\ln l\,,$
we need to prove that $G,H$ commute in $\Gamma$. Hence, it suffices to prove that the elements $xgx^{-1}$ and $h$ commute, since, for example, this implies that the axes (the whole axes) of $G$ and $H$ are of bounded Hausdorff distance from each other.
We consider the rectangle
$xg^{r}y^{-1}h^{-s}$
in the Cayley graph of $\Gamma$, where $r=r_{2}-r_{1}$, and $s=s_{2}-s_{1}$. Since
$d(p_{1},p_{2}),\,d(q_{1},q_{2})\geq 9l\ln l\,,$
we have that
$r,s\geq\frac{\ln l}{1000\alpha_{0}}\,.$
And since the nerve $\gamma_{G}$ is a $200\alpha_{0}l$-local geodesic, and hence a $(\frac{204}{196},2\alpha_{0}l)$-quasi-geodesic, we have that
$r\cdot 200\alpha_{0}l\leq r\cdot[g]\leq\frac{204}{196}\cdot|g^{r}|+2\alpha_{0}l\leq\frac{204}{196}\cdot 10l\ln l+2\alpha_{0}l$
$\implies$ $r\leq\frac{\frac{204}{196}\cdot 10\ln l+2\alpha_{0}}{200\alpha_{0}}\leq\ln l\,.$
Similarly,
$s\leq\ln l\,.$
Hence, we have that
$\ln l\geq r,s\geq\frac{\ln l}{1000\alpha_{0}}\,.$
For every $\ln l\geq i,j\geq\frac{\ln l}{1000\alpha_{0}}$, Let $\mathcal{V}^{i,j}$ be the collection of all the reduced decorated topological Van-Kampen diagrams $DTD$ with decoration $\hat{x}\hat{g}^{i}\hat{y}^{-1}\hat{h}^{-j}$, and boundary length at most $4000\alpha_{0}l\ln l$, that $\Gamma$ could satisfy, and so that the lengths of the parts of the boundary of $DTD$ corresponding to the variables $\hat{x},\hat{y}$, sum to at most $2\nu_{0}l$. Let $\mathcal{V}^{0}$ be the union of all of these collections,
$\mathcal{V}^{0}=\underset{i,j}{\cup}\mathcal{V}^{i,j}\,.$
For each diagram $DTD$ in $\mathcal{V}^{0}$, we declare $\hat{x}$ and $\hat{y}$ to be the rigid variables of the decoration, and no constant variables.
Since the boundary length of each of the diagrams in $\mathcal{V}^{0}$ is at most $4000\alpha_{0}l\ln l$, we have according to 4.1, that the amount of cells $m$ in any diagram in $\mathcal{V}^{0}$ is bounded by $m\leq\beta_{0}\ln l$, where $\beta_{0}=\frac{4000\alpha_{0}}{\frac{1}{2}-d}$ (see 4.2). Moreover, the decoration $\hat{x}\hat{g}^{i}\hat{y}^{-1}\hat{h}^{-j}$ of each diagram in $\mathcal{V}^{0}$ is of length at most $22\ln l+$ 2. Hence, according to 4.5, for all $i,j$, there exists a polynomial $Q^{i,j}=Q_{d}^{i,j}$, so that
$|\mathcal{V}^{i,j}|\leq Q^{i,j}(l)^{\ln^{2}l}\,.$
And hence, there exists a polynomial $Q=Q_{d}$, so that
$|\mathcal{V}^{0}|\leq Q(l)^{\ln^{2}l}\,.$
We generally reduce each of the diagrams in $\mathcal{V}^{0}$, and denote the collection of the obtained generally reduced diagrams by $\mathcal{V}$. Recall according to 3.26 that general reduction can only increase the probability of fulfilling the diagram.
For a given generally reduced decorated topological Van-Kampen diagram $DTD$ in $\mathcal{V}$, with $m$ cells numbered by $1,...,n$, for some $1\leq n\leq m$, we describe the following iterative procedure. If $DTD$ contains an isolated numbering, we apply a mining modification on $DTD$. If the obtained diagram, denoted $DTD$ again, contains an isolated numbering, we apply another mining modification. We continue iteratively, until we arrive to a generally reduced decorated diagram $DTD$ with no isolated numberings, or we apply at most
$k_{0}=\left\lceil\frac{\nu_{0}}{\frac{1}{2}-d}\right\rceil+1$
iterative mining modifications. Note that $k_{0}=k_{0,d}$ depends on $d$ only.
We denote by $\mathcal{V}_{1}$ the sub-collection of $\mathcal{V}$ consisting of those diagrams $DTD$ in $\mathcal{V}$ for which the procedure performs $k_{0}$ iterations.
We denote by $\mathcal{V}_{ab}$ the sub-collection of $\mathcal{V}\backslash\mathcal{V}_{1}$ consisting of those diagrams $DTD$ in $\mathcal{V}$ for which the procedure stops (before applying $k_{0}$ iterations) in a diagram with no cells (a tree).
And we denote by $\mathcal{V}_{2}$ the remaining diagrams of $\mathcal{V}$; i.e., $\mathcal{V}_{2}=\mathcal{V}\backslash\mathcal{V}_{1}\cup\mathcal{V}_{ab}$ consists of those diagrams in $\mathcal{V}$ for which the procedure stops (before applying $k_{0}$ iterations) in a diagram with a positive number of cells. We replace the diagrams in $\mathcal{V}_{2}$ by the obtained diagrams by applying the procedure on them.
We denote by $\mathcal{V}_{1}^{0},\mathcal{V}_{ab}^{0},\mathcal{V}_{2}^{0}$ the sub-collections consisting of those diagrams in $\mathcal{V}^{0}$ who admit an associated diagram in $\mathcal{V}_{1},\mathcal{V}_{ab},\mathcal{V}_{2}$ respectively.
By the definition of the procedure above, and according to 3.25, the amount of distinct numberings $n$ in any diagram in the collection $\mathcal{V}_{1}$, is at least
$n\geq k_{0}=\left\lceil\frac{\nu_{0}}{\frac{1}{2}-d}\right\rceil+1\,.$
Hence, according to 3.28, for every diagram $DTD$ in $\mathcal{V}_{1}$ with $n$ distinct numberings, the probability that the random group $\Gamma$ of level $l$ and density $d$ fulfills $DTD$ is at most
$\left(\frac{2k}{(2k-1)^{(\frac{1}{2}-d-\frac{\nu_{0}}{n})l}}\right)^{n}\leq\frac{2k}{(2k-1)^{\gamma_{0}l}}\,,$
where $\gamma_{0}=\frac{1}{2}-d-\frac{\nu_{0}}{k_{0}}>0$ (recall that the rigid parts in $DTD$ sum to at most $2\nu_{0}l$).
On the other hand, by the definition of the collection $\mathcal{V}_{2}$, every diagram in $\mathcal{V}_{2}$ admits a positive amount $n\geq 1$ of distinct numberings, but contains no isolated numberings. Hence, according to 3.28, for every diagram $DTD$ in $\mathcal{V}_{2}$, the probability that $\Gamma$ fulfills $DTD$ is at most
$\left(\frac{2k}{(2k-1)^{(\frac{1}{2}-d)l}}\right)^{n}\leq\frac{2k}{(2k-1)^{\gamma_{0}l}}\,.$
Thus, we conclude that the probability that there exists some diagram $DTD$ in the collection $\mathcal{V}_{1}\cup\mathcal{V}_{2}$ so that the random group $\Gamma$ of level $l$ and density $d$ fulfills $DTD$, is at most
$|\mathcal{V}_{1}\cup\mathcal{V}_{2}|\cdot\frac{2k}{(2k-1)^{\gamma_{0}l}}\leq|\mathcal{V}^{0}|\cdot\frac{2k}{(2k-1)^{\gamma_{0}l}}\leq\frac{2k\cdot Q(l)^{\ln^{2}l}}{(2k-1)^{\gamma_{0}l}}\,\overset{l\rightarrow\infty}{\longrightarrow}0$
which, according to 5.4, approaches 00 as $l$ goes to infinity.
Recall according to 3.25 that mining modifications can only increase the probability of fulfilling the diagram. Hence, we conclude that the probability $p_{1,2}(l)$ that there exists some diagram $D$ in the collection $\mathcal{V}_{1}^{0}\cup\mathcal{V}_{2}^{0}$ so that the random group $\Gamma$ of level $l$ and density $d$ fulfills $D$, approaches 00 as $l$ goes to infinity
$p_{1,2}(l)\overset{l\rightarrow\infty}{\longrightarrow}0\,.$
Hence we may assume that the random group $\Gamma$ of level $l$ and density $d$, does not fulfill any of the diagrams in the collection $\mathcal{V}_{1}^{0}\cup\mathcal{V}_{2}^{0}$.
Now recall our rectangle
$xg^{r}y^{-1}h^{-s}$
in the Cayley graph of $\Gamma$, explained in the top of the proof. Since $|x|,|y|\leq\nu_{0}l$, since $|g|,|h|\leq 1000\alpha_{0}l$, and since $\ln l\geq r,s\geq\frac{\ln l}{1000\alpha_{0}}$, the group $\Gamma$ must fulfill some reduced decorated Van-Kampen diagram $DTD$ with decoration $\hat{x}\hat{g}^{r}\hat{y}^{-1}\hat{h}^{-s}$ in the collection $\mathcal{V}^{0}$, so that the contour of the obtained (reduced) Van-Kampen diagram $VKD$ is $xg^{r}y^{-1}h^{-s}$.
However, since we have already proved that the collection of groups of level $l$ and density $d$ that satisfy a diagram in $\mathcal{V}_{1}^{0}\cup\mathcal{V}_{2}^{0}$ is negligible, we can assume that $DTD$ belongs to the collection $\mathcal{V}_{ab}^{0}$. Since the topological diagram $DTD$, by the definition of $\mathcal{V}_{ab}^{0}$, can be converted to a tree by means of applying topological modifications (mining and general reduction) on it, the Van-Kampen diagram $VKD$ can be converted to a tree $T$ by applying the same sequence of modifications on it. However, these modifications do not change the values (in $\Gamma$) of the variables along the boundary.
Hence, the procedure of applying those modifications on $VKD$, provides lifts $\tilde{x},\tilde{y},\tilde{g},\tilde{h}\in F_{k}$ (reduced words - the obtained words on the boundary of $T$ could be non-reduced, but we simply reduce them) for the elements $x,y,g,h\in\Gamma$ respectively, so that
$\tilde{x}\tilde{g}^{r}\tilde{y}^{-1}\tilde{h}^{-s}=1$
in the free group $F_{k}$. However, according to the definition of $\mathcal{V}_{ab}^{0}$, the tree $T$ is obtained by applying a sequence of alternate general reduction and mining modifications on $VKD$, applied for at most $k_{0}=k_{0,d}$ iterations. Now, by definition, general reduction can change the rigid parts $x,y$ of the boundary, only by removing self-intersections from them, but a single mining modification can change a rigid part by eliminating only cells that share some edge with that part. Thus, the lifts $\tilde{x},\tilde{y}$ that are obtained as a consequence of applying at most $k_{0}=k_{0,d}$ mining modifications, lie, as paths in the Cayley graphof $\Gamma$, in the $k_{0}l$-nbhd of the geodesic segments $x,y$ in the Cayley graph of $\Gamma$, respectively.
In particular, each of the lifts $\tilde{x}$ and $\tilde{y}$, as paths in the Cayley graph of $\Gamma$, is contained in a ball of radius
$\frac{|x|}{2}+k_{0}l,\,\frac{|y|}{2}+k_{0}l\leq 2k_{0}l\,.$
Finally, we consider the obtained lifts $\tilde{g},\tilde{h}\in F_{k}$ for the elements $g,h\in\Gamma$. Write $\tilde{g}=Y^{\prime}uY^{\prime-1}$, and $\tilde{h}=Z^{\prime}wZ^{\prime-1}$ (graphical equality), for reduced words $Y^{\prime},Z^{\prime}\in F_{k}$ and $u,w\in F_{k}$ cyclically reduced. Since the translation length of $g$ in $\Gamma$ satisfies $[g]\geq 200\alpha_{0}l$, we have that for all $k\in\mathbb{Z}$, the length of $u^{k}$ in $\Gamma$ satisfies:
$|u^{k}|\geq[u^{k}]=[g^{k}]\geq[g^{k}]_{\infty}=k\cdot[g]_{\infty}\geq k\cdot 168\alpha_{0}l\,.$
Similarly,
$|w^{k}|\geq k\cdot 168\alpha_{0}l\,.$
In particular, a path induced by the powers of the word $u$ in the Cayley graph of $\Gamma$, can count at most
$\left\lceil\frac{4k_{0}}{168\alpha_{0}}\right\rceil=n_{0}=n_{0,d}$
“vertices” inside a ball of radius $2k_{0}l$ (by vertices we mean the points on the path induced by the powers $w^{n}$, $n\in\mathbb{Z}$). Similarly, at most $n_{0}$ powers of $w$ can lie inside a ball of radius $2k_{0}l$ in the Cayley graph of $\Gamma.$
In particular, since each of the lifts $\tilde{x}$ and $\tilde{y}$, as paths in the Cayley graph of $\Gamma$, is contained in a ball of radius $2k_{0}l$, we deduce that neither $\tilde{x}$, nor $\tilde{y}$, admits a sub-word of the form $u^{\pm(n_{0}+1)},w^{\pm(n_{0}+1)}$. Hence, and since
$r,s\geq\frac{\ln l}{1000\alpha_{0}}\geq 110n_{0}\,,$
we get according to 5.3, that the words $\tilde{x}\tilde{g}\tilde{x}^{-1}$ and $\tilde{h}$ commute. Hence, of course, the elements $xgx^{-1}$ and $h$ of $\Gamma$, are commutative. As claimed. ∎
8 Hyperbolic Geometry over Random Groups
Lemma 8.1.
(See Lemma 2.10 in [Co]) Let $\Gamma=\langle a:\mathcal{R}\rangle$ be a $\delta$-hyperbolic group, and let $X$ be its Cayley graph. Let $c$ be a geodesic line in $X$. Let $x,y\in\Gamma$ be two points, and let $p,q\in c$ be projections on $c$ for $x,y$ respectively.
If $d(p,q)\geq 50\delta$, then
$d(x,y)+50\delta\geq d(x,p)+d(p,q)+d(q,y)\,.$
Proof.
We consider the rectangle whose (geodesic) sides are $[x,y],\,[x,p],\,[p,q],\,[y,q]$.
First note that if a point on $[x,p]$ is $2\delta$-close to a point on $[y,q]$, then $p,q$ must be $10\delta$-close to each other, a contradiction.
Now if the $2\delta$-nbhd of a point $r$ on $[x,p]$ contains a point on $c$, then, since $p$ is a projection of $x$ on $c,$ we get that $d(r,p)\leq 2\delta$. Similarly, if the $2\delta$-nbhd of a point $s$ on $[y,q]$ contains a point on $c$, then $d(s,q)\leq 2\delta$.
These imply, using the $\delta$-hyperbolicity, that the three segments $[x,p],\,[p,q],\,[y,q]$ are contained in the $4\delta$-nbhd of $[x,y]$. In particular, there exists two points $t_{1},t_{2}$ on $[x,y]$ that are $4\delta$-close to $p,q$ respectively. Hence,
$d(x,p)\leq d(x,t_{1})+d(t_{1},p)\leq d(x,t_{1})+4\delta\,,$
$d(p,q)\leq d(p,t_{1})+d(t_{1},t_{2})+d(t_{2},q)\leq d(t_{1},t_{2})+8\delta\,,$
$d(q,y)\leq d(q,t_{2})+d(t_{2},y)\leq d(t_{2},y)+4\delta\,.$
Thus,
$d(x,p)+d(p,q)+d(q,y)\leq d(x,y)+16\delta\,.$
Lemma 8.2.
(See Proposition 2.24 in [Co]) Let $d<\frac{1}{2}$, and let $\Gamma$ be the random group of level $l$ and density $d$. Let $X$ be the Cayley graph of $\Gamma$ (w.r.t. the fixed generating set $a$).
Then, with probability that tends to $1$, the following property is satisfied in $\Gamma$.
Let $G\in\Gamma$ be a hyperbolic element, and let $A_{G}$ be an axis for $G$. Let $x\in\Gamma$, and assume that $d(Gx,x)\leq A+[G]$, for some real number $A\geq 0$.
Then,
  1. 1.
    If $[G]\geq 200\alpha_{0}l$, then
    $d(x,A_{G})\leq\frac{A}{2}+50\alpha_{0}l\,.$
  2. 2.
    If $[G]<200\alpha_{0}l$, then
    $d(x,A_{G})\leq A+(7500\alpha_{0}+3\zeta_{0})l\,,$
    where $\zeta_{0}$ is given in 7.1.
Proof.
Since $d(Gx,x)=d(1,x^{-1}Gx)$, and $d(x,A_{G})=d(1,A_{x^{-1}Gx})$, and $[x^{-1}Gx]=[G]$, we can assume that $x=1$. For simplicity we write $\delta=\alpha_{0}l$.
Let $p,q\in A_{G}$ be projections for $x=1$ and $Gx=G$ on $A_{G}$, and consider a geodesic segments $[x,p],\,[Gx,q]$ between $x$ and $p$, and $Gx$ and $q$. Since $p\in A_{G}$, we have that $Gp$ belongs to the $5\delta$-nbhd of $A_{G}$. Hence,
$[G]\leq d(p,Gp)\leq[G]+15\delta\,.$
Moreover, since $p$ is a projection for $x$ on $A_{G}$, we have that $Gp$ is a projection for $Gx$ on $GA_{G}$. But $A_{G}$ and $GA_{G}$ are both axes for $G$, and in particular the Hausdorff distance between them is at most $5\delta$. Hence, the projection $q$ of $Gx$ on $A_{G}$ and the projection $Gp$ of $Gx$ on $GA_{G}$, are $30\delta$-close to each other
$d(q,Gp)\leq 30\delta\,.$
Hence,
$d(p,q)\leq d(p,Gp)+d(Gp,q)\leq[G]+45\delta\,,$
$d(p,q)\geq d(p,Gp)-d(Gp,q)\geq[G]-15\delta\,.$
In particular, if $[G]\geq 200\delta$, then $d(p,q)\geq 50\delta$, and thus by 8.1, we get that
$d(x,p)+d(Gx,q)\leq d(x,Gx)-d(p,q)+50\delta\leq A+65\delta\,.$
But $d(Gx,q)\geq d(Gx,Gp)-30\delta=d(x,p)-30\delta$, so we conclude that
$d(x,p)\leq\frac{A}{2}+50\delta\,,$
in case $[G]\geq 200\delta$.
Now assume that $[G]<200\delta$.
Write $G=zgz^{-1}$, for geodesic words $z,g\in F_{k}$, with $|g|=[G]$. Since
$|G|=d(x,Gx)\leq[G]+A\leq 200\delta+A\,,$
and $|g|=[G]\leq 200\delta$, we can assume by 7.1 that
$|z|\leq A+(408\alpha_{0}+\zeta_{0})l\,,$
where $\zeta_{0}=\frac{100\alpha_{0}}{(\frac{1}{2}-d)}$. In particular, $x=1$ is of distance at most $A+(408\alpha_{0}+\zeta_{0})l$ from the nerve $\gamma_{G}$ of $G$ induced by the elements $zg^{n}$, $n\in\mathbb{Z}$.
Hence, according 7.8, we deduce that $x=1$ is of distance at most $A+(7500\alpha_{0}+3\zeta_{0})l$ from $A_{G}$. ∎
Theorem 8.3.
(See Proposition 2.41 in [Co]) Let $d<\frac{1}{2}$, and let $\Gamma$ be the random group of level $l$ and density $d$. Let $X$ be the Cayley graph of $\Gamma$ (w.r.t. the fixed generating set $a$).
Then, with probability that tends to $1$, the following property is satisfied in $\Gamma$.
Let $G,H\in\Gamma$ be two non-commuting elements, and let $A_{G},A_{H}$ be two axes for them respectively. Then
$\text{diam}(A_{G}^{+60\alpha_{0}l}\cap A_{H}^{+60\alpha_{0}l})\leq 4\cdot([G]+[H])+D(\Gamma,X)+10\alpha_{0}\cdot l\ln l\,.$
Proof.
For simplicity we write $\delta=\alpha_{0}l$.
According to the definition of $D(\Gamma,X)$, we can assume that $[G]>1000\delta$. By raising $H$ to a suitable power if necessary, we can assume that $[H]\geq 200\delta$, and prove then that
$\text{diam}(A_{G}^{+60\delta}\cap A_{H}^{+60\delta})\leq 4\cdot([G]+[H])+D(\Gamma,X)+9\delta\ln l\,.$
Indeed, assume by contradiction that
$\text{diam}(A_{G}^{+60\delta}\cap A_{H}^{+60\delta})>4\cdot([G]+[H])+D(\Gamma,X)+9\delta\ln l\,.$
By the $\alpha_{0}l$-hyperbolicity, we get that there exists two sub-segments $I_{1}=[x_{1},y_{1}]\subset A_{G}$, and $I_{2}=[x_{2},y_{2}]\subset A_{H}$, whose Hausdorff distance from each other is at most $5\delta$, and their lengths satisfy $|I_{1}|,|I_{2}|\geq 6\cdot([G]+[H])+D(\Gamma,X)+8\delta\ln l$. And again by $\delta$-hyperbolicity, these two sub-segments admit two sub-segments $I_{1}^{\prime}=[x_{1}^{\prime},y_{1}^{\prime}]$, and $I_{2}^{\prime}=[x_{2}^{\prime},y_{2}^{\prime}]$, whose Hausdorff distance is at most $5\delta$ from each other, their lengths satisfy $|I_{1}^{\prime}|,|I_{2}^{\prime}|\geq 2\cdot([G]+[H])+D(\Gamma,X)+7\delta\ln l$, and whose endpoints are of distances at least $[G]+[H]+1000\delta$ from the endpoints of $I_{1}$ and $I_{2}$, respectively.
Let $p_{1},p_{2}$ be two points on $I_{1}^{\prime},I_{2}^{\prime}$ respectively, so that
$d(p_{1},p_{2})\leq 5\delta\,.$
Then, Since $p_{1}\in A_{G}$, we have
$d(Gp_{1},p_{1})\leq 10\delta+[G]\,.$
Let $q_{1}$ be a projection of $Gp_{1}$ on $A_{G}$. Since $Gp_{1}$ belongs to the $5\delta$-nbhd of $A_{G}$, we deduce that
$[G]-15\delta\leq d(p_{1},q_{1})\leq[G]+15\delta\,.$
In particular, $q_{1}$ belongs to $I_{1}$, and hence to the $5\delta$-nbhd of $I_{2}$. Let $q_{2}$ be a projection of $q_{1}$ on $I_{2}$. Note that
$d(q_{1},q_{2})\leq 5\delta,\quad d(Gp_{1},q_{2})\leq 10\delta\,.$
Hence
$d(p_{2},q_{2})\leq d(p_{2},p_{1})+d(p_{1},q_{1})+d(q_{1},q_{2})\leq[G]+25\delta\,,$
$d(p_{2},q_{2})\geq d(p_{1},q_{1})-d(p_{2},p_{1})-d(q_{1},q_{2})\geq[G]-25\delta\,.$
Now since $Gp_{1}$ belongs to the $10\delta$-nbhd of $A_{H}$, we get that
$[H]-30\delta\leq d(HGp_{1},Gp_{1})\leq[H]+30\delta\,,$
and that $HGp_{1}$ belongs to the $15\delta$-nbhd of $A_{H}$. Let $r_{2}$ be a projection of $HGp_{1}$ on $A_{H}$. Of course
$d(HGp_{1},r_{2})\leq 15\delta\,,$
and we deduce that
$d(q_{2},r_{2})\leq d(q_{2},Gp_{1})+d(Gp_{1},HGp_{1})+d(HGp_{1},r_{2})\leq[H]+55\delta\,,$
$d(q_{2},r_{2})\geq d(Gp_{1},HGp_{1})-d(q_{2},Gp_{1})-d(HGp_{1},r_{2})\geq[H]-55\delta\,.$
By replacing $H$ by $H^{-1}$ if necessary, we may assume that $q_{2}$ lies between $p_{2}$ and $r_{2}$ (see 7.7). Hence,
$[G]+[H]-80\delta\leq d(p_{2},r_{2})=d(p_{2},q_{2})+d(q_{2},r_{2})\leq[G]+[H]+80\delta\,.$
In particular $r_{2}$ belongs to $I_{2}$, and hence to the $5\delta$-nbhd of $I_{1}$. Since $d(HGp_{1},r_{2})\leq 15\delta$, we have that $HGp_{1}$ belongs to the $20\delta$-nbhd of $I_{1}$. Let $r_{1}$ be a projection of $HGp_{1}$ on $I_{1}$. Then,
$d(HGp_{1},r_{1})\leq 20\delta\,,\quad d(r_{1},r_{2})\leq 35\delta\,.$
Then,
$d(p_{1},r_{1})\leq d(p_{1},p_{2})+d(p_{2},r_{2})+d(r_{2},r_{1})\leq[G]+[H]+120\delta\,,$
$d(p_{1},r_{1})\geq d(p_{2},r_{2})-d(p_{1},p_{2})-d(r_{2},r_{1})\geq[G]+[H]-120\delta\,.$
Now by applying the same argument again, we deduce that $GHp_{2}$ lies in the $15\delta$-nbhd of $A_{G}$, and if $r_{1}^{\prime}$ is a projection of $GHp_{2}$ on $A_{G}$, then
$[G]+[H]-80\delta\leq d(p_{1},r_{1}^{\prime})\leq[G]+[H]+80\delta\,.$
Hence,
$d(GHp_{2},HGp_{1})\leq d(GHp_{2},r_{1}^{\prime})+d(r_{1}^{\prime},r_{1})+d(r_{1},HGp_{1})\leq 190\delta\,.$
Hence,
$d(GHp_{1},HGp_{1})\leq d(GHp_{1},HGp_{2})+d(HGp_{2},HGp_{1})\leq 195\delta\,.$
Thus, the commutator $u=G^{-1}H^{-1}GH$ satisfies that
$d(up,p)\leq 195\delta$
for every $p\in I_{1}^{\prime}$, which implies also that
$[u]\leq 195\delta\,.$
According to the assumption, $u\neq 1$. Now since $d(up,p)\leq 195\delta$ for all $p\in I_{1}^{\prime}$, we get by 8.2 that $I_{1}^{\prime}$ lies in the $(10405\alpha_{0}+3\zeta_{0})l=\lambda_{0}\delta$-nbhd of an axis $A_{u}$ of $u$, where $\lambda_{0}=\frac{(10405\alpha_{0}+3\zeta_{0})}{\alpha_{0}}$.
Since
$|I_{1}^{\prime}|\geq 2[G]+D(\Gamma,X)+7\delta\ln l\,,$
then, by the $\delta$-hyperbolicity of $\Gamma$, there exist sub-segments $J_{1}$ and $J_{2}$ of $I_{1}^{\prime}\subset A_{G}$ and $A_{u}$ respectively, so that the Hausdorff distance between $J_{1}$ and $J_{2}$ is at most $5\delta$, and $|J_{1}|,|J_{2}|\geq 2[G]+D(\Gamma,X)+6\delta\ln l$. Say $G$ translates “right”, and denote by $e$ the “left” endpoint of $J_{1}$ (see 7.7). Let $f$ be a point on $J_{1}$ so that
$D(\Gamma,X)+200\delta\leq d(e,f)\leq[G]+D(\Gamma,X)+200\delta\,.$
Let $p_{e},p_{f}$ be projections of $e,f$ on $J_{2}$ respectively. Then,
$d(p_{e},p_{f})\geq d(e,f)-10\alpha_{0}l\geq D(\Gamma,X)+190\delta\,.$
Now, since $|J_{1}|\geq 2[G]+D(\Gamma,X)+6\delta\ln l$ and $G$ translates “right”, each of $Ge$ and $Gf$ is of $10\delta$-distance from $J_{1}$. Hence each of $Gp_{e}$ and $Gp_{f}$ is of $15\delta$-distance from $J_{1}$, and hence, each of $Gp_{e}$ and $Gp_{f}$ is of $20\delta$-distance from $J_{2}$. In particular,
$Gp_{e},Gp_{f}\in A_{u}^{+60\delta}\,.$
But $G[p_{e},p_{f}]$, where $[p_{e},p_{f}]$ is the corresponding sub-segment of $A_{u}$, is a sub-segment of an axis $A_{GuG^{-1}}=GA_{u}$ for the element $GuG^{-1}$, whose translation length satisfies $[GuG^{-1}]=[u]\leq 1000\delta$. Thus, by the definition of $D(\Gamma,X)$, either $|G[p_{e},p_{f}]|\leq D(\Gamma,X)$, or otherwise $GuG^{-1}$ commutes with $u$. We already know that $|G[p_{e},p_{f}]|=d(p_{e},p_{f})\geq D(\Gamma,X)+190\delta$, so we conclude that the elements $GuG^{-1}$ and $u$ commute. Hence, Since $G,u$ are hyperbolic elements in the hyperbolic group $\Gamma$, we conclude that $G$ commutes with $u$. But $u=G^{-1}H^{-1}GH$, and thus, the elements $G$ and $H^{-1}GH$ commute. And again, this implies that $G$ and $H$ commute. ∎
Definition 8.4.
Let $d<\frac{1}{2}$. An ascending sequence of groups in the model of density $d$, is a sequence of groups $\{\Gamma_{l_{n}}\}_{n}$, so that:
  1. 1.
    The sequence $\{l_{n}\}_{n}$ is an unbounded non-decreasing sequence of integers.
  2. 2.
    For all $n$, the group $\Gamma_{l_{n}}$ belongs to the level $l_{n}$ of the model (of density $d$).
Theorem 8.5.
Let $f,t$ be two real functions with $f(x),t(x)=o(x\ln^{2}x)$, and let $d<\frac{1}{2}$. Let $\{\Gamma_{l_{n}}\}_{n}$ be an ascending sequence of groups in the model.
For all $n$, let $g_{n},h_{n}$ be two elements in the group $\Gamma_{l_{n}}$, and let $A_{g_{n}},A_{h_{n}}$ be two axes for them respectively. Denote
$I_{n}=\text{diam}(A_{g_{n}}^{+f(l_{n})}\cap A_{h_{n}}^{+f(l_{n})})\,.$
Assume that the sequences $[g_{n}],[h_{n}]$ of translation lengths of $g_{n},h_{n}$ resp. in $\Gamma_{l_{n}}$, is bounded by
$[g_{n}],[h_{n}]\leq t(l_{n})\,.$
If the sequence
$\frac{I_{n}}{l_{n}\ln^{2}l_{n}},$
is bounded from below by a positive number, then $g_{n},h_{n}$ commute eventually.
Proof.
By raising $g_{n}$ and $h_{n}$ to large enough powers, and modifying the function $t$ if necessary, we can assume that the translation lengths of $g_{n}$ and $h_{n}$ in $\Gamma_{l_{n}}$ satisfy $[g_{n}],[h_{n}]\geq 200\alpha_{0}l_{n}$.
Since $f(x)=o(x\ln^{2}x)$, and the sequence
$\frac{I_{n}}{l_{n}\ln^{2}l_{n}},$
is bounded from below by a positive number, we deduce by the $\alpha_{0}l_{n}$-hyperbolicity of $\Gamma_{l_{n}}$ that the sequence
$\frac{\text{diam}(A_{g_{n}}^{+60\alpha_{0}l_{n}}\cap A_{h_{n}}^{+60\alpha_{0}l_{n}})}{l_{n}\ln^{2}l_{n}}$
is bounded from below by a positive number.
Now if, by contradiction, the elements $g_{n},h_{n}$ do not commute eventually, then, according to 8.3, (for a subsequence of $l_{n}$) we get that
$\frac{\text{diam}(A_{g_{n}}^{+60\alpha_{0}l_{n}}\cap A_{h_{n}}^{+60\alpha_{0}l_{n}})}{l_{n}\ln^{2}l_{n}}\leq\frac{4\cdot([g_{n}]+[h_{n}])+D(\Gamma_{l_{n}},X)+10\alpha_{0}l_{n}\ln l_{n}}{l_{n}\ln^{2}l_{n}}\,.$
And hence, according to 7.9,
$\frac{\text{diam}(A_{g_{n}}^{+60\alpha_{0}l_{n}}\cap A_{h_{n}}^{+60\alpha_{0}l_{n}})}{l_{n}\ln^{2}l_{n}}\leq\frac{20l_{n}\cdot\ln l_{n}+8t(l_{n})}{l_{n}\ln^{2}l_{n}}\,,$
which approaches 00 as $n$ goes to $\infty$. A contradiction. ∎
9 Limit Groups over Ascending Sequences of Random Groups
[…]
9.1 Limit Groups over Ascending Sequences of Random Groups
Theorem 9.1.
Let $F_{q}$ be a free group, with a basis $x=(x_{1},...,x_{q})$. Let $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$ be a sequence of homomorphisms over an ascending sequence of groups in the model.
Assume that the sequence of stretching factors
$\mu_{n}=\underset{f_{n}\in\Gamma_{l_{n}}}{\min}\underset{1\leq u\leq q}{\max}d_{\Gamma_{l_{n}}}(1,\tau_{f_{n}}\circ h_{l_{n}}(x_{u}))\,,$
satisfies that
$\mu_{n}\geq l_{n}\ln^{2}l_{n}$
for all $n$, where $\tau_{f_{n}}$ is the inner automorphism $g\mapsto f_{n}gf_{n}^{-1}$ of $\Gamma_{l_{n}}$.
For all $n$, we pick an element $f_{n}\in\Gamma_{l_{n}}$, so that
$\mu_{n}=\underset{1\leq u\leq q}{\max}d_{\Gamma_{l_{n}}}(1,\tau_{f_{n}}\circ h_{l_{n}}(x_{u}))\,.$
Then, the sequence $\tau_{f_{n}}\circ h_{l_{n}}$, subconverges (in the Gromov topology) to an isometric non-trivial action of $F_{q}$ on a pointed real tree $(Y,y_{0})$.
Proof.
For all $n$, let $\delta_{n}$ be the hyperbolicity constant of the group $\Gamma_{l_{n}}$.
According to 4.4, we have that $\delta_{n}=\alpha_{0}l_{n}$. Then,
$\frac{\delta_{n}}{\mu_{n}}\overset{n\rightarrow\infty}{\longrightarrow}0\,.$
Hence, according to Proposition 1.1 in [Se1], the limit space of the sequence of the rescaled Cayley graphs $\{(X_{l_{n}}/\mu_{n},1)\}_{n}$ of the groups $\{\Gamma_{l_{n}}\}_{n}$, is 00-hyperbolic, i.e., a real tree. ∎
Theorem 9.2.
Keep the notation of 9.1.
We denote
$K_{h_{l_{n}},\infty}=\{g\in F_{q}:\forall y\in Y,\quad gy=y\}\,.$
Let $L_{h_{l_{n}},\infty}$ be the quotient group $F_{q}/K_{h_{l_{n}},\infty}$. Then,
  1. 1.
    $L_{h_{l_{n}},\infty}$ is a f.g. group.
  2. 2.
    $L_{h_{l_{n}},\infty}$ is torsion-free.
  3. 3.
    If $Y$ is a real line, then $L_{h_{l_{n}},\infty}$ is free abelian.
  4. 4.
    If an element $g\in F_{q}$ stabilizes (pointwise) a non-trivial tripod in $Y$, then $h_{l_{n}}(g)=1$ eventually.
  5. 5.
    Let $g\in F_{q}$ be an element that does not belong to $K_{h_{l_{n}},\infty}$. Then, $h_{l_{n}}(g)\neq 1$ eventually.
  6. 6.
    Let $[y_{1},y_{2}]\subset[y_{3},y_{4}]$ be a pair of non-degenerate segments in $Y$, and assume that $stab([y_{3},y_{4}])$, the (pointwise) stabilizer of $[y_{3},y_{4}]$ is non-trivial. Then, $stab([y_{3},y_{4}])$ is abelian, and
    $stab([y_{1},y_{2}])=stab([y_{3},y_{4}])\,.$
Proof.
Point 1 is clear.
For the other statements, let $g_{1},g_{2}\in F_{q}$.
If $g_{1}$ either stabilizes a non-trivial tripod in $Y$, or $Y$ is a real line and $g_{1}$ does not translate it, then, by the definition of convergence, according to 8.2, and since random groups are torsion-free, for large enough $n$, we must have that $h_{l_{n}}(g_{1})=1$ in $\Gamma_{l_{n}}$.
Assume now that $g_{1},g_{2}$ stabilizes a non-degenerate segment in $Y$. Then, according to 8.5, the elements $h_{l_{n}}(g_{1})$ and $h_{l_{n}}(g_{2})$ of $\Gamma_{l_{n}}$, must commute eventually.
Finally, if $g_{1}\in stab([y_{1},y_{2}])$ and $g_{2}\in stab([y_{3},y_{4}])$, then $g_{2}$ stabilizes the tripod $[y_{3},g_{1}y_{3},y_{2}]$. ∎
Definition 9.3.
Let $F_{q}$ be a free group, with a basis $x=(x_{1},...,x_{q})$. Let $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$ be a sequence of homomorphisms over an ascending sequence of groups in the model.
Assume that there exists a positive number $\epsilon>0$, so that the sequence of stretching factors
$\mu_{n}=\underset{f_{n}\in\Gamma_{l_{n}}}{\min}\underset{1\leq u\leq q}{\max}d_{\Gamma_{l_{n}}}(1,\tau_{f_{n}}\circ h_{l_{n}}(x_{u}))\,,$
satisfies that
$\mu_{n}\geq\epsilon l_{n}\ln^{2}l_{n}$
for all $n$. For all $n$, let $f_{n}\in\Gamma_{l_{n}}$ be an element so that
$\mu_{n}=\underset{1\leq u\leq q}{\max}d_{\Gamma_{l_{n}}}(1,\tau_{f_{n}}\circ h_{l_{n}}(x_{u}))\,.$
The sequence $g_{n}=\tau_{f_{n}}\circ h_{l_{n}}$, is called an $lln^{2}$-sequence over the ascending sequence $\Gamma_{l_{n}}$ of groups in the model, or briefly, an $lln^{2}$-sequence.
In what follows in this section, we will usually write $h_{l_{n}}$ in place of $\tau_{f_{n}}\circ h_{l_{n}}$.
In case the sequence $\tau_{f_{n}}\circ h_{l_{n}}$ converges (in the Gromov topology) to an isometric non-trivial action of $F_{q}$ on a real tree, then, in accordance with the notation of 9.2, the quotient group
$L_{h_{l_{n}},\infty}=F_{q}/K_{h_{l_{n}},\infty}$
is called an $lln^{2}$-limit group over the ascending sequence $\Gamma_{l_{n}}$ of groups in the model, or briefly, an $lln^{2}$-limit group.
We say also that the $lln^{2}$-sequence $h_{l_{n}}$ converges into the $lln^{2}$-limit group $L_{h_{l_{n}},\infty}$.
Definition 9.4.
Let $F_{q}$ be a free group, with a fixed basis $x=(x_{1},...,x_{q})$, and let $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$ be a sequence of homomorphisms over an ascending sequence of groups in the model. Let $\pi:F_{k}\rightarrow\Gamma$ be a given presentation of a group $\Gamma$, and let $h:F_{q}\rightarrow\Gamma$ be a homomorphism.
  1. 1.
    The length of the homomorphism $h:F_{q}\rightarrow\Gamma$ is defined to be
    $||h||=\underset{1\leq u\leq 1}{\max}d_{\Gamma}(1,h(x_{u}))\,.$
  2. 2.
    Let $\Sigma$ be a collection of words in the free group $F_{q}$. Let $\pi:F_{k}\rightarrow\Gamma$ be a given presentation of a group $\Gamma$. Let $h:F_{q}\rightarrow\Gamma$ be a homomorphism with $h(\Sigma)=1$. We say that the homomorphism $h$ can be $\Sigma$-lifted, if there exists some homomorphism $\tilde{h}:F_{q}\rightarrow F_{k}$ with $\tilde{h}(\Sigma)=1$ and $h=\pi\circ\tilde{h}$.
  3. 3.
    We say that the sequence $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$ is convergent if for all $w\in F_{q}$, either $h_{l_{n}}(w)=1$ eventually, or $h_{l_{n}}(w)\neq 1$ eventually. The collection
    $K_{\infty}(h_{l_{n}})=\{w\in F_{q}:h_{l_{n}}(w)=1\text{ eventually}\}$
    is called the kernal of the sequence $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$. In this case, the group
    $F_{q}/K_{\infty}(h_{l_{n}})$
    is called the algebraic limit of the sequence $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$.
  4. 4.
    We say that the sequence $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$ is a non-lift sequence, if there exists a finite set of words $\Sigma\in F_{q}$, so that $h_{l_{n}}(\Sigma)=1$ eventually, but for all $n$, the homomorphism $h_{l_{n}}$ cannot be $\Sigma$-lifted.
  5. 5.
    If the sequence $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$ is a non-lift sequence that converges to an $lln^{2}$-limit group $L_{\infty}$, then the group $L_{\infty}$ is called a non-lift $lln^{2}$-limit group.
Remark 9.5.
  1. 1.
    Since we have dropped the collection $\mathcal{N}$, we can assume that any non-lift sequence is in necessity an $lln^{2}$-sequence.
  2. 2.
    Our interest in the remainder of this chapter will be in non-lift $lln^{2}$-limit groups. Paradoxically, our aim in this chapter is showing that there exist no non-lift $lln^{2}$-limit groups.
Definition 9.6.
Let $F_{q}$ be a free group, with a fixed basis $x=(x_{1},...,x_{q})$. For $j=1,2$, , let $\{h_{l_{n}^{j}}^{j}:F_{q}\rightarrow\Gamma_{l_{n}^{j}}^{j}\}_{n}$ be a non-lift sequence of homomorphisms over an ascending sequence of groups in the model, that converges to the non-lift $lln^{2}$-limit group $\eta_{j}:F_{q}\rightarrow L_{j}$.
If the function $\eta_{1}(x_{i})\mapsto\eta_{2}(x_{i})$, $i=1,...,q$, defines a homomorphism $L_{1}\rightarrow L_{2}$, then, we say that $L_{2}$ is an $lln^{2}$-limit quotient of $L_{1}$, and write
$L_{1}\geq L_{2}\,.$
If further the map $\eta_{1}(x_{i})\mapsto\eta_{2}(x_{i})$, $i=1,...,q$, is not injective, we say that $L_{2}$ is a proper $lln^{2}$-limit quotient of $L_{1}$, and write
$L_{1}>L_{2}\,.$
Lemma 9.7.
Let $F_{q}$ be a free group, with a fixed basis $x=(x_{1},...,x_{q})$. Let $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$ be a non-lift sequence that converges to a non-lift $lln^{2}$-limit group $\eta:F_{q}\rightarrow L_{\infty}$. Then,
  1. 1.
    $L_{\infty}$ is not free, nor free abelian.
  2. 2.
    Let $L_{\infty}=L_{1}\ast...\ast L_{g}\ast F$ be a Grushko decomposition for $L_{\infty}$, so that $L_{j}$ is non-cyclic, $j=1,...,g$, and $F$ is a free group. Then, there exists $j_{0}=1,...,g$, so that the subgroup $L_{j_{0}}$ is a non-lift $lln^{2}$-limit group.
    Moreover, let $b=(b_{1},...,b_{p})$ be a generating set for $L_{j_{0}}$. Let $F_{p}$ be a free group with a fixed basis $y=(y_{1},...,y_{p})$. Let $\beta:F_{p}\rightarrow F_{q}$ be defined by mapping $y_{i}$ to some $\eta$-preimage of $b_{i}$, $i=1,...,p$. Then, the sequence
    $\{h_{l_{n}}\circ\beta:F_{p}\rightarrow\Gamma_{l_{n}}\}$
    contains a non-lift subsequence that converges to $L_{j_{0}}$.
Proof.
For Point 1, if $L_{\infty}$ was free or free abelian, then, according to 9.2, and by moving to a subsequence if necessary, the algebraic limit $A_{\infty}$ of the sequence $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$ must be a finitely generated free group, or a finitely generated free abelian group. Hence, the homomorphisms in the sequence $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$ factor through $A_{\infty}$ eventually. But $A_{\infty}$ is free or free abelian, and the groups $\Gamma_{l_{n}}$ are torsion-free hyperbolic. This contradicts the assumption that the sequence $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$ is a non-lift sequence.
For point 2, let $b^{1},...,b^{g},f$ be generating sets of sizes $p_{1},...,p_{g},r$ for the components $L_{1},...,L_{g},F$ respectively. Let
$\beta_{i}:F_{p_{i}}\rightarrow F_{q},\quad\beta_{f}:F_{r}\rightarrow F_{q}$
be maps that map the $j$-th element in a fixed basis for the respective free group $F_{p_{i}}$ or $F_{r}$, to an $\eta$-preimage of the $j$-th element in the respective set $b^{i}$ or $f$, $i=1,...,g$.
We consider the sequences
$\{h_{l_{n}}^{i}:F_{p_{i}}\rightarrow\Gamma_{l_{n}}\},\,i=1,...,g,\,\text{ and }\{h_{l_{n}}^{f}:F_{r}\rightarrow\Gamma_{l_{n}}\}\,,$
given by
$h_{l_{n}}^{i}=h_{l_{n}}\circ\beta_{i},\quad h_{l_{n}}^{f}=h_{l_{n}}\circ\beta_{f}\,.$
We also denote $\eta_{i}:F_{p_{i}}\rightarrow L_{i}$, and $\eta_{f}:F_{r}\rightarrow F$ the maps $\eta_{i}=\eta\circ\beta_{i}$, and $\eta_{f}=\eta\circ\beta_{f}$ respectively.
According to the proof of Point 1, the sequence $\{h_{l_{n}}^{f}:F_{r}\rightarrow\Gamma_{l_{n}}\}$ contains no non-lift subsequence. Moreover, for large enough $n$, the homomorphism $h_{l_{n}}^{f}:F_{r}\rightarrow\Gamma_{l_{n}}$ can be lifted by a homomorphism $\tilde{h}_{l_{n}}^{f}:F_{r}\rightarrow F_{k}$ that factors through $\eta_{f}:F_{r}\rightarrow F$, that is, $h_{l_{n}}^{f}=\pi_{\Gamma_{l_{n}}}\circ\tilde{h}_{l_{n}}^{f}$ and there exists some homomorphism $\lambda_{n}^{f}:L_{i}\rightarrow F_{k}$ so that $\tilde{h}_{l_{n}}^{f}=\lambda_{n}^{f}\circ\eta_{f}$.
Now assume by contradiction that for all $i=1,...,g$, the sequence $\{h_{l_{n}}^{i}:F_{p_{i}}\rightarrow\Gamma_{l_{n}}\}$ contains no non-lift subsequence. Then, for all $i=1,...,g$, according to the definition of non-lift sequences, for every finite collection $\Sigma$ in $F_{p_{i}}$, if $h_{l_{n}}^{i}(\Sigma)=1$ eventually, then, the homomorphisms $h_{l_{n}}^{i}$ can be $\Sigma$-lifted eventually.
Let $i=1,...,g$. By Guba’s theorem, every system over the free group $F_{k}$, is equivalent to a finite subsystem. In particular, the system describing the defining relations of the subgroup $L_{i}$ (w.r.t. the set of generators $b^{i}$), is equivalent to a finite subsystem over the free group $F_{k}$. Recall that for every $w$ in $F_{p_{i}}$, we have that $\eta_{i}(w)=1$ if and only if $w$ is mapped by $h_{l_{n}}^{i}$ to the trivial element for all large enough $n$.
Hence, we conclude that for all large enough $n$, the homomorphism $h_{l_{n}}^{i}:F_{p_{i}}\rightarrow\Gamma_{l_{n}}$ can be lifted by a homomorphism $\tilde{h}_{l_{n}}^{i}:F_{p_{i}}\rightarrow F_{k}$ that factors through $\eta_{i}:F_{p_{i}}\rightarrow L_{i}$, that is, $h_{l_{n}}^{i}=\pi_{\Gamma_{l_{n}}}\circ\tilde{h}_{l_{n}}^{i}$ and there exists some homomorphism $\lambda_{n}^{i}:L_{i}\rightarrow F_{k}$ so that $\tilde{h}_{l_{n}}^{i}=\lambda_{n}^{i}\circ\eta_{i}$.
Summing together, we obtain that for all large enough $n$, the homomorphisms $\lambda_{n}^{1},...,\lambda_{n}^{g},\lambda_{n}^{f}$ implies a homomorphism
$\lambda_{n}:L_{\infty}\rightarrow F_{k}\,,$
so that
$\pi_{\Gamma_{l_{n}}}\circ\lambda_{n}\circ\eta=h_{l_{n}}\,.$
In particular, the homomorphism $\lambda_{n}\circ\eta:F_{q}\rightarrow F_{k}$ lifts the homomorphism $h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}$.
In order to get a contradiction, recall that the sequence $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$ is non-lift. I.e., there exists a finite collection $\Sigma_{0}$ of elements in $F_{q}$, so that $h_{l_{n}}(\Sigma_{0})=1$ eventually, but for all $n$, the homomorphism $h_{l_{n}}$ cannot be $\Sigma_{0}$-lifted. However, as the sequence $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$ converges to $L_{\infty}$, the canonical map $\eta:F_{q}\rightarrow L_{\infty}$ must map the elements of $\Sigma_{0}$ to $1$. Hence, $\lambda_{n}\circ\eta$ is a $\Sigma_{0}$-lift for $h_{l_{n}}$ for all large enough $n$, a contradiction. ∎
9.2 Shortening Quotients of Limit Groups over Ascending Sequences of Random Groups
Lemma 9.8.
(See Lemma 1.4 in [Se1]) Let $F_{q}$ be a free group, with a fixed basis $x=(x_{1},...,x_{q})$. Let $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$ be a non-lift sequence, that converges to a non-lift $lln^{2}$-limit group
$\eta:F_{q}\rightarrow L_{\infty}\,.$
Then, the following properties are satisfied in $L_{\infty}$.
  1. 1.
    Let $u_{1},u_{2},u_{3}$ be non-trivial elements of $L_{\infty}$, and suppose that $[u_{1},u_{2}]=1$ and $[u_{1},u_{3}]=1$. Then, $[u_{2},u_{3}]=1$. It follows that every abelian subgroup in $L_{\infty}$ is contained in a unique maximal abelian subgroup.
  2. 2.
    Every maximal abelian subgroup of $L_{\infty}$ is malnormal.
  3. 3.
    Every solvable subgroup of $L_{\infty}$ is abelian.
Proof.
In light of 9.2, exactly the same proof of Lemma 1.4 in [Se1] proves the statement in the lemma, with the only exception that instead of using malnormality of maximal cyclic subgroups of $F_{k}$, we use malnormality of maximal cyclic subgroups of torsion-free hyperbolic groups (in proving Point 2). ∎
Let $L_{\infty}$ be a non-lift $lln^{2}$-limit group, and assume that $L_{\infty}$ is freely indecomposable. In light of 9.8, 9.1, and 9.2, Chapter 2 of [Se1] can be applied (as it is) in our context, in order to construct the abelian JSJ decomposition of the freely indecomposable non-lift $lln^{2}$-limit group $L_{\infty}$. For the exact definition of the JSJ decomposition, see Chapter 2 of [Se1]. For the exact statement on how the (abelian) JSJ decomposition encodes all the abelian splittings of $L_{\infty}$, see Theorem 2.7 in [Se1].
Definition 9.9.
Let $F_{q}$ be a free group, with a fixed basis $x=(x_{1},...,x_{q})$. Let $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$ be a non-lift sequence, that converges to a non-lift $lln^{2}$-limit group
$\eta:F_{q}\rightarrow L_{\infty}\,.$
According to 9.7, the group $L_{\infty}$ is non-free.
We are going to define a “shortening quotient” of $L_{\infty}$, inspired by the manner explained in [[Gr3], sections 3 and 4], and that in [Se-Ja].
  • Assume first that $L_{\infty}$ is freely-indecomposable.
Let $\Lambda_{L_{\infty}}$ be the abelian JSJ decomposition of $L_{\infty}$. Let $V^{1},...,V^{m}$, $E^{1},...,E^{s}$, and $t^{1},...,t^{b}$ be the corresponding vertex groups, edge groups, and Bass-Serre generators (w.r.t some maximal subtree of $\Lambda_{L_{\infty}}$) respectively, of $\Lambda_{L_{\infty}}$. Since $L_{\infty}$ is finitely generated, for all $j=1,...,m$, the vertex group $V^{j}$ is generated by a finite subset $v_{1}^{j},...,v_{r_{j}}^{j}$, together with the adjacent edge groups in $\Lambda_{L_{\infty}}$. For all $j=1,...,s$, we choose an ordered basis for the edge group $E^{j}$, and write $E_{p}^{j}$ for the first $p$ elements in that ordered basis.
For all $n$, let $U_{n}$ be the group given by the presentation
$U_{n}=\langle v_{u}^{j},t^{j},E_{n}^{j}:\mathcal{R}_{n}\rangle\,,$
where $\mathcal{R}_{n}$ denotes the collection of all the relations of length $n$ between the generators $v_{u}^{j},t^{j},E_{n}^{j}$ as an elements of $L_{\infty}$. Let
$\varphi_{n}:U_{n}\rightarrow L_{\infty}$
be the obvious mapping.
Since $L_{\infty}$ is generated by
$V^{1},...,V^{m},E^{1},...,E^{s},t^{1},...,t^{b}$
we get that, for some integer $p$, there exist words $w_{i}=w_{i}(v_{u}^{j},t^{j},E_{p}^{j})$, $i=1,...,q$, in the formal elements $v_{u}^{j},t^{j},E_{n}^{j}$, so that $w_{i}$ represents the element $g_{i}$ in $L_{\infty}$, where
$g_{1}=\eta(x_{1}),...,g_{q}=\eta(x_{q})$
($x=(x_{1},...,x_{q})$ is the fixed basis of $F_{q}$). According to that, for all $n$, we define the map
$\psi_{n}:F_{q}\rightarrow U_{n}\,,$
by $\psi_{n}(x_{i})=w_{i}$. Note that $\varphi_{n}\circ\psi_{n}=\eta$ for all $n$ (recall $\eta:F_{q}\rightarrow L_{\infty}$ is the canonical map).
For every integer $n$, the relations of the group $U_{n}$ are taken from those in $L_{\infty}$. Hence, in accordance with 9.2, there exists an ascending sequence $m(n)$ of integers, so that the homomorphism $h_{l_{m(n)}}$ factors through $\psi_{n}:F_{q}\rightarrow U_{n}$, i.e., for all $n$, there exists a homomorphism
$\lambda_{n}:U_{n}\rightarrow\Gamma_{l_{m(n)}}\,,$
so that $\lambda_{n}\circ\psi_{n}=h_{l_{m(n)}}$.
The sequence of groups $U_{n}$ is thought of as an “approximation sequence” for the decomposition $\Lambda_{L_{\infty}}$. Indeed, since $L_{\infty}$ is freely indecomposable non-cyclic, and since the sequence $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$ converges to a non-trivial action of $L_{\infty}$ on a pointed real tree so that the properties listed in 9.2 are satisfied, we get according to the shortening argument presented in [[Gr3], sections 3 and 4], that for all large enough $n$, the homomorphism $h_{l_{m(n)}}$ can be shortened (strictly) by “intermediate” composing with an automorphism of $Mod(U_{n})$, or post-composing with an inner automorphism of $\Gamma_{l_{m(n)}}$.
That is, for all large enough $n$, there exists a modular automorphism $\phi_{n}\in Mod(U_{n})$, together with some element $f_{n}\in\Gamma_{l_{m(n)}}$, so that
$||h_{l_{m(n)}}||>||\tau_{f_{n}}\circ\lambda_{n}\circ\phi_{n}\circ\psi_{n}||\,.$
For all large enough $n$, we choose $\phi_{n}\in Mod(U_{n})$ and $f_{n}\in\Gamma_{l_{m(n)}}$ so that the right hand side of the above inequality is minimal possible. Then, we consider the sequence of shortened homomorphisms
$v_{l_{m(n)}}=\tau_{f_{n}}\circ\lambda_{n}\circ\phi_{n}\circ\psi_{n}:F_{q}\rightarrow\Gamma_{l_{m(n)}}\,.$
Such a sequence $\{v_{l_{m(n)}}:F_{q}\rightarrow\Gamma_{l_{m(n)}}\}$ is called a shortened sequence for the sequence $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$.
Lemma 9.10.
Let $\{v_{l_{m(n)}}:F_{q}\rightarrow\Gamma_{l_{m(n)}}\}$ be a shortened sequence for the non-lift sequence $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$. Then, the sequence $\{v_{l_{m(n)}}:F_{q}\rightarrow\Gamma_{l_{m(n)}}\}$ contains a non-lift subsequence.
Proof.
Assume by contradiction the claim is false. By moving to a subsequence if necessary, we may assume that the sequence $\{v_{l_{m(n)}}:F_{q}\rightarrow\Gamma_{l_{m(n)}}\}$ is convergent. Then, according to the definition of non-lift sequences, for every finite collection $\Sigma$ in $F_{q}$, if $v_{l_{m(n)}}(\Sigma)=1$ eventually, then, the homomorphisms $v_{l_{m(n)}}$ can be $\Sigma$-lifted for all large enough $n$.
By Guba’s theorem, every system over the free group $F_{k}$, is equivalent to a finite subsystem. In particular, the system describing the defining relations of the group $L_{\infty}$ (w.r.t. the set of generators $\eta(x)$, where $x$ is the fixed basis of $F_{q}$), is equivalent to a finite subsystem over the free group $F_{k}$. Recall that for every $w$ in $F_{q}$, we have that $\eta(w)=1$ if and only if $w$ is mapped by $\psi_{n}:F_{q}\rightarrow U_{n}$ to the trivial element for all large enough $n$.
Hence, we conclude that for all large enough $n$, the homomorphism $v_{l_{m(n)}}:F_{q}\rightarrow\Gamma_{l_{m(n)}}$ can be lifted by a homomorphism $\tilde{v}_{l_{m(n)}}:F_{q}\rightarrow F_{k}$ that factors through $\eta:F_{q}\rightarrow L_{\infty}$, that is, $v_{l_{m(n)}}=\pi_{\Gamma_{l_{m(n)}}}\circ\tilde{v}_{l_{m(n)}}$ and there exists some homomorphism
$\tilde{u}_{n}:L_{\infty}\rightarrow F_{k}$
so that $\tilde{v}_{l_{m(n)}}=\tilde{u}_{n}\circ\eta$.
We denote
$u_{n}=\pi_{\Gamma_{l_{m(n)}}}\circ\tilde{u}_{n}:L_{\infty}\rightarrow\Gamma_{l_{m(n)}}\,,$
and we note that, tautologically, the homomorphism $\tilde{u}_{n}$ lifts the homomorphism $u_{n}$. Hence, the homomorphism $\tilde{u}_{n}\circ\varphi_{n}:U_{n}\rightarrow F_{k}$ lifts the homomorphism
$\tau_{f_{n}}\circ\lambda_{n}\circ\phi_{n}:U_{n}\rightarrow\Gamma_{l_{m(n)}}\,.$
Let $\tilde{f}_{n}\in F_{k}$ be a word representing the element $f_{n}$ in $\Gamma_{l_{m(n)}}$. Then, since
$\pi_{\Gamma_{l_{m(n)}}}\circ\tau_{\tilde{f}_{n}^{-1}}\circ\tilde{u}_{n}\circ\varphi_{n}\circ\phi_{n}^{-1}$ $=$
$=\tau_{f_{n}^{-1}}\circ\pi_{\Gamma_{l_{m(n)}}}\circ\tilde{u}_{n}\circ\varphi_{n}\circ\phi_{n}^{-1}$
$=\tau_{f_{n}^{-1}}\circ\tau_{f_{n}}\circ\lambda_{n}\circ\phi_{n}\circ\phi_{n}^{-1}$
$=\lambda_{n}\,,$
we have that the homomorphism
$\tilde{\lambda}_{n}=\tau_{\tilde{f}_{n}^{-1}}\circ\tilde{u}_{n}\circ\varphi_{n}\circ\phi_{n}^{-1}:U_{n}\rightarrow F_{k}$
(denoted $\tilde{\lambda}_{n}$) lifts the homomorphism
$\lambda_{n}:U_{n}\rightarrow\Gamma_{l_{m(n)}}\,.$
We recall the homomorphism
$h_{l_{m(n)}}=\lambda_{n}\circ\psi_{n}:F_{q}\rightarrow\Gamma_{l_{m(n)}}\,,$
and we note that the homomorphism
$\tilde{h}_{l_{m(n)}}=\tilde{\lambda}_{n}\circ\psi_{n}:F_{q}\rightarrow F_{k}$
lifts it.
As we have mentioned above, by Guba’s theorem, the system describing the defining relations of the group $L_{\infty}$, is equivalent to a finite subsystem over the free group $F_{k}$. Thus, since every element $w$ of $F_{q}$ with $\eta(w)=1$ is mapped by $\psi_{n}$ to the trivial element in $U_{n}$ for all large enough $n$, we conclude that for all large enough $n$, the homomorphism $\tilde{h}_{l_{m(n)}}$ factors through $\eta:F_{q}\rightarrow L_{\infty}$. Hence, the original sequence $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$ is not non-lift, a contradiction. ∎
We continue the definition with our shortened sequence $\{v_{l_{m(n)}}:F_{q}\rightarrow\Gamma_{l_{m(n)}}\}$ (for the original non-lift sequence $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$).
According to 9.10, by moving to a subsequence, we may assume the shortened sequence $\{v_{l_{m(n)}}:F_{q}\rightarrow\Gamma_{l_{m(n)}}\}$ is non-lift again. Hence, according to 9.1, we may assume further that the shortened sequence $\{v_{l_{m(n)}}:F_{q}\rightarrow\Gamma_{l_{m(n)}}\}$ converges to a (non-lift) $lln^{2}$-limit group $\eta_{Q}:F_{q}\rightarrow Q_{\infty}$.
Such an $lln^{2}$-limit group $Q_{\infty}$ is called an $lln^{2}$-shortening quotient of the non-lift $lln^{2}$-limit group $L_{\infty}$.
Note that, since every element $w$ of $F_{q}$ with $\eta(w)=1$ is mapped by $\psi_{n}$ to the trivial element in $U_{n}$ for all large enough $n$, we have that $Q_{\infty}$ is indeed a quotient of $L_{\infty}$; the map $\eta(x)\mapsto\eta_{Q}(x)$ is a quotient map:
$L_{\infty}\geq Q_{\infty}\,.$
Moreover, as we have explained above, it cannot happen that $L_{\infty}\cong Q_{\infty}$, for otherwise, one can shorten the homomorphism $v_{l_{m(n)}}$ by composing with an automorphism in $Mod(U_{n})$ (the same $U_{n}$ defined above) and inner automorphism of $\Gamma_{l_{m(n)}}$, for infinitely many integers $n$, which contradicts the construction of $v_{l_{m(n)}}$. Hence, actually $Q_{\infty}$ is a proper quotient of $L_{\infty}$:
$L_{\infty}>Q_{\infty}\,.$
  • Now assume that $L_{\infty}$ admits a non-trivial Grushko decomposition $L_{\infty}=L_{1}\ast...\ast L_{g}\ast F$, so that $L_{j}$ is non-cyclic, $j=1,...,g$, and $F$ is a free group.
According to 9.7, we may assume that $L_{1}$ is a non-lift $lln^{2}$-limit group. Moreover, choosing a generating set $b^{1}=(b_{1}^{1},...,b_{p_{1}}^{1})$ for $L_{1}$, choosing a free group $F_{p_{1}}$ with basis $y^{1}=(y_{1}^{1},...,y_{p_{1}}^{1})$, and passing to a subsequence, we may assume that the sequence
$\{h_{l_{n}}^{1}=h_{l_{n}}\circ\beta_{1}:F_{p_{1}}\rightarrow\Gamma_{l_{n}}\}$
is a non-lift sequence, where $\beta_{1}:F_{p_{1}}\rightarrow F_{q}$ is the map defined by mapping $y_{i}^{1}$ to some $\eta$-preimage of $b_{i}^{1}$, $i=1,...,p_{1}$. We denote also
$\eta_{1}:F_{p_{1}}\rightarrow L_{1}$
the map $\eta_{1}=\eta\circ\beta_{1}$.
We further, abuse the notation and denote the free component $L_{2}\ast...\ast L_{g}\ast F$ by $L_{2}$.
We fix a generating set $b^{2}=(b_{1}^{2},...,b_{p_{2}}^{2})$ for $L_{2}$, a free group $F_{p_{2}}$ with basis $y^{2}=(y_{1}^{2},...,y_{p_{2}}^{2})$, and a map $\beta_{2}:F_{p_{2}}\rightarrow F_{q}$ defined by mapping $y_{i}^{2}$ to some $\eta$-preimage of $b_{i}^{2}$, $i=1,...,p_{2}$. We consider the sequence
$\{h_{l_{n}}^{2}=h_{l_{n}}\circ\beta_{2}:F_{p_{2}}\rightarrow\Gamma_{l_{n}}\}\,,$
and we denote by
$\eta_{2}:F_{p_{2}}\rightarrow L_{2}$
the map $\eta_{2}=\eta\circ\beta_{2}$.
Now the sequence
$\{h_{l_{n}}^{1}:F_{p_{1}}\rightarrow\Gamma_{l_{n}}\}$
is a non-lift sequence that converges to the freely-indecomposable $lln^{2}$-limit group $L_{1}$. Hence, we may bring the notation of the first part of the definition (when $L_{\infty}$ was freely-indecomposable) for the sequence $\{h_{l_{n}}^{1}:F_{p_{1}}\rightarrow\Gamma_{l_{n}}\}$ and the non-lift $lln^{2}$-limit group $L_{1}$, with the difference that we add an superscript $1$, e.g., in place of $U_{n}$, we write $U_{n}^{1}$.
We further define new groups $U_{n}^{2}$ to be “approximations” for the group $L_{2}$, i.e., for all $n$ we define
$U_{n}^{2}=\langle b^{2}:\mathcal{R}_{n}^{2}\rangle\,,$
where $\mathcal{R}_{n}^{2}$ denotes the collection of all the relations of length $n$ between the generators $b^{2}$ as an elements of $L_{2}$.
We consider the groups
$U_{n}=U_{n}^{1}\ast U_{n}^{2}\,.$
Since $L_{\infty}$ is generated by the corresponding elements in the JSJ of $L_{1}$
$V^{1},...,V^{m},E^{1},...,E^{s},t^{1},...,t^{b}$
together with $b^{2}$, we get that, for some integer $r$, there exist words $w_{i}=w_{i}(v_{u}^{j},t^{j},E_{r}^{j},b^{2})$, $i=1,...,q$, in the formal elements $v_{u}^{j},t^{j},E_{n}^{j},b^{2}$, so that $w_{i}$ represents the element $g_{i}$ in $L_{\infty}$, where
$g_{1}=\eta(x_{1}),...,g_{q}=\eta(x_{q})$
($x=(x_{1},...,x_{q})$ is the fixed basis of $F_{q}$). According to that, for all $n$, we define the map
$\psi_{n}:F_{q}\rightarrow U_{n}\,,$
by $\psi_{n}(x_{i})=w_{i}$. Note that $\varphi_{n}\circ\psi_{n}=\eta$ for all $n$ (recall $\eta:F_{q}\rightarrow L_{\infty}$ is the canonical map).
We consider the maps
$\psi_{n}^{j}:F_{p_{j}}\rightarrow U_{n}$
given by $\psi_{n}^{j}=\psi_{n}\circ\beta_{j}$, and we note that for all large enough $n$, the image of $\psi_{n}^{j}$ is contained entirely in the component $U_{n}^{j}$, $j=1,2$.
As in the previous part of the definition, there exists an ascending sequence $m(n)$ of integers, so that the homomorphism $h_{l_{m(n)}}$ factors through $\psi_{n}:F_{q}\rightarrow U_{n}$, i.e., for all large enough $n$, there exists a homomorphism
$\lambda_{n}:U_{n}\rightarrow\Gamma_{l_{m(n)}}\,,$
so that $\lambda_{n}\circ\psi_{n}=h_{l_{m(n)}}$. For $j=1,2$, we denote by
$\lambda_{n}^{j}:U_{n}^{j}\rightarrow\Gamma_{l_{m(n)}}\,,$
the restriction of $\lambda_{n}$ on $U_{n}^{j}$.
Now, for all large enough $n$, there exists a modular automorphism $\phi_{n}^{1}\in Mod(U_{n}^{1})$, together with some element $f_{n}\in\Gamma_{l_{m(n)}}$, so that
$||h_{l_{m(n)}}^{1}||>||\tau_{f_{n}}\circ\lambda_{n}\circ\phi_{n}^{1}\circ\psi_{n}^{1}||\,.$
For all large enough $n$, we choose $\phi_{n}^{1}\in Mod(U_{n})$ and $f_{n}\in\Gamma_{l_{m(n)}}$ so that the right hand side of the above inequality is minimal possible. Then, we consider the sequence of homomorphisms
$v_{l_{m(n)}}=\tau_{f_{n}}\circ\lambda_{n}\circ\phi_{n}^{1}\circ\psi_{n}:F_{q}\rightarrow\Gamma_{l_{m(n)}}\,.$
For $j=1,2$, we let $v_{l_{m(n)}}^{j}:F_{p_{j}}\rightarrow\Gamma_{l_{m(n)}}$ be the homomorphism
$v_{l_{m(n)}}^{j}=v_{l_{m(n)}}\circ\beta_{j}\,.$
We note that for all large enough $n$, we have that
$v_{l_{m(n)}}^{2}=\tau_{f_{n}}\circ\lambda_{n}\circ\psi_{n}^{2}=\tau_{f_{n}}\circ h_{l_{m(n)}}^{2}\,.$
And since the algebraic limit of the sequence $\{h_{l_{n}}^{2}:F_{p_{2}}\rightarrow\Gamma_{l_{n}}\}$ is $L_{2}$, we have that the algebraic limit of (every subsequence of) the sequence $\{v_{l_{m(n)}}^{2}:F_{p_{2}}\rightarrow\Gamma_{l_{n}}\}$ is $L_{2}$ too.
On the other hand, the sequence $\{v_{l_{m(n)}}^{1}:F_{p_{1}}\rightarrow\Gamma_{l_{n}}\}$ is a shortened sequence for the sequence $\{h_{l_{n}}^{1}:F_{p_{1}}\rightarrow\Gamma_{l_{n}}\}$, and hence, according to 9.10, we may assume it is a non-lift sequence. Moreover, we may assume that the sequence $\{v_{l_{m(n)}}^{1}:F_{p_{1}}\rightarrow\Gamma_{l_{n}}\}$ converges to an $lln^{2}$-shortening quotient $Q_{1}$ of $L_{1}$.
In accordance with that, we call the sequence $\{v_{l_{m(n)}}:F_{q}\rightarrow\Gamma_{l_{m(n)}}\}$ a shortened sequence for the sequence $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$.
Lemma 9.11.
(In the case when $L_{\infty}$ is freely-decomposable too). Let $\{v_{l_{m(n)}}:F_{q}\rightarrow\Gamma_{l_{m(n)}}\}$ be a shortened sequence for the non-lift sequence $\{h_{l_{n}}:F_{q}\rightarrow\Gamma_{l_{n}}\}$. Then, the sequence $\{v_{l_{m(n)}}:F_{q}\rightarrow\Gamma_{l_{m(n)}}\}$ contains a non-lift subsequence.
Proof.
Assume by contradiction the claim is false. We already know that the sequence $\{v_{l_{m(n)}}^{1}:F_{p_{1}}\rightarrow\Gamma_{l_{n}}\}$ is non-lift. Let $\Sigma$ be a finite subset of $F_{p_{1}}$ so that $v_{l_{m(n)}}^{1}(\Sigma)=1$ eventually, and for all $n$, the homomorphism $v_{l_{m(n)}}^{1}$ cannot be $\Sigma$-lifted.
Let $\Sigma^{\prime}$ be the finite subset $\beta_{1}(\Sigma)$ of $F_{q}$. Since $v_{l_{m(n)}}^{1}=v_{l_{m(n)}}\circ\beta_{1}$, we conclude that $v_{l_{m(n)}}(\Sigma^{\prime})=1$ eventually. Since we are assuming the claim is false, for all large enough $n$, the homomorphism $v_{l_{m(n)}}$ can be $\Sigma^{\prime}$-lifted. But this implies that for all large enough $n$, the homomorphism $v_{l_{m(n)}}^{1}$ can be $\Sigma$-lifted, a contradiction. ∎
Hence, in total, by moving to a subsequence, the shortened sequence $\{v_{l_{m(n)}}:F_{q}\rightarrow\Gamma_{l_{m(n)}}\}$ is a non-lift sequence, that converges to the non-lift $lln^{2}$-limit group $\eta_{Q}:F_{q}\rightarrow Q_{\infty}=Q_{1}\ast L_{2}$, and since every element $w$ of $F_{q}$ with $\eta(w)=1$ is mapped by $\psi_{n}$ to the trivial element in $U_{n}$ for all large enough $n$, we have that the map $\eta(x)\mapsto\eta_{Q}(x)$ is a quotient map
$L_{\infty}\geq Q_{\infty}\,.$
Moreover, since $Q_{1}$ is a proper quotient of $L_{1}$, we have that $Q_{\infty}$ is a proper quotient of $L_{\infty}$
$L_{\infty}>Q_{\infty}\,.$
Such an $lln^{2}$-limit group $Q_{\infty}$ is called an $lln^{2}$-shortening quotient of the non-lift $lln^{2}$-limit group $L_{\infty}$.
Corollary 9.12.
Let $L_{\infty}$ be a non-lift $lln^{2}$-limit group. Then, $L_{\infty}$ admits an infinite (strictly) decreasing sequence of non-lift $lln^{2}$-limit quotients
$L_{\infty}>Q_{\infty}^{1}>Q_{\infty}^{2}>...\,.$
Proof.
As we have explained in 9.9, every non-lift $lln^{2}$-limit group $L_{\infty}$ admits an $lln^{2}$-shortening quotient $Q_{\infty}$, so that $Q_{\infty}$ itself is a non-lift $lln^{2}$-limit group which is a proper quotient of $L_{\infty}$. Hence, by induction, one can create an infinite decreasing sequence as the required one. ∎
Theorem 9.13.
Every strictly decreasing sequence
$L_{1}>L_{2}>L_{3}>...$
of non-lift $lln^{2}$-limit groups, terminates after finitely many steps.
Proof.
We follow the pf of theorem 13 in [Se-Ja]. Assume by contradiction the claim is false.
We construct then the following strictly decreasing sequence of non-lift $lln^{2}$-limit groups.
For brevity, given a non-lift $lln^{2}$-limit group $R$, we denote by $\mathcal{A}_{R}$ the collection of all the non-lift $lln^{2}$-limit groups $R_{1}^{\prime}$, for which $R>R_{1}^{\prime}$ and there exists a strictly decreasing infinite sequence of non-lift $lln^{2}$-limit groups that begins with $R_{1}^{\prime}$
$R_{1}^{\prime}>R_{2}^{\prime}>R_{3}^{\prime}>...\,.$
For convenience, given a free group $F_{q}$ with a free basis $x=(x_{1},...,x_{q})$, we write $\mathcal{A}_{F_{q}}$ to denote the collection of all the non-lift $lln^{2}$-limit groups $R$ whose canonical generating set consists of $q$ elements, and for which $\mathcal{A}_{R}$ is non-empty. For all $R$ in $\mathcal{A}_{F_{q}}$, we denote $\eta_{R}:F_{q}\rightarrow R$ the quotient map that maps $x_{i}$ to the $i$-th element in the canonical generating set of $R$, for all $i=1,...,q$.
From the collection of free groups (bases) $F_{q}$ for which $\mathcal{A}_{F_{q}}$ is non-empty, we choose an $F_{q}$ with $q$ minimal.
We define the strictly decreasing sequence
$R_{1}>R_{2}>R_{3}>...$
by induction. First, we denote $R_{0}=F_{q}$. Then, assuming that $R_{n-1}$ is defined, and that $\mathcal{A}_{R_{n-1}}$ is non-empty, we let $R_{n}$ to be an element of the collection $\mathcal{A}_{R_{n-1}}$, so that the map $\eta_{R_{n}}=\eta_{n}:F_{q}\rightarrow R_{n}$, maps to the identity the maximal possible number of elements in the ball of radius $n$ in the Cayley graph of $F_{q}$ (w.r.t. the fixed basis $x$).
For brevity, given a sequence of homomorphisms $\{v_{m}:F_{q}\rightarrow\Gamma_{m}\}$, an integer $m_{0}$, and a collection of elements $A\subset F_{q}$, we say that the sequence $\{v_{m}:F_{q}\rightarrow\Gamma_{m}\}$ stabilizes $A$ $m_{0}$-eventually, if for all $w\in A$, we have that [$v_{m}(w)=1$ for all $m\geq m_{0}$, or $v_{m}(w)\neq 1$ for all $m\geq m_{0}$].
Now for all $n\geq 1$, the group $R_{n}$ is in particular a non-lift $lln^{2}$-limit group. For all $n\geq 1$, let
$\{v_{m}^{n}:F_{q}\rightarrow\Gamma_{l_{m}^{n}}^{n}\}_{m}$
be a non-lift sequence that converges to $R_{n}$. For all $n\geq 1$, using the definition of non-lift sequences, let $\Sigma_{n}$ be a finite subset of $F_{q}$, so that $v_{m}^{n}(\Sigma_{n})=1$ for all large enough $m$, and assume that the homomorphism $v_{m}^{n}$ cannot be $\Sigma_{n}$-lifted for all $m$. Given two integers $n\geq i\geq 1$, we recall that $R_{n}\geq R_{i}$. Hence, $v_{m}^{n}(\Sigma_{i})=1$ for all large enough $m$. Thus, we may assume that the sequence $\{\Sigma_{n}\}_{n}$ is an ascending sequence
$\Sigma_{1}\subset\Sigma_{2}\subset\Sigma_{3}\subset...\,.$
In accordance with 9.7 and 9.2, for all $n\geq 1$, let $m(n)$ be a large enough integer, so that
  1. 1.
    $v_{m(n)}^{n}(\Sigma_{n})=1$,
  2. 2.
    the sequence $\{v_{m}^{n}:F_{q}\rightarrow\Gamma_{l_{m}^{n}}^{n}\}_{m}$ stabilizes the ball of radius $n$ of the Cayley graph of $F_{q}$ $m(n)$-eventually,
  3. 3.
    and so that after all, the sequence $\{\Gamma_{l_{m(n)}^{n}}^{n}\}_{n}$ is an ascending sequence of groups in the model.
Of course, the sequence of integers $\{m(n)\}_{n}$ can further be chosen to be strictly increasing, and thus, to simplify the notation and write
$\Gamma_{l_{m(n)}}=\Gamma_{l_{m(n)}^{n}}^{n}$
without ambiguity.
We define the sequence of homomorphisms $\{h_{n}:F_{q}\rightarrow\Gamma_{l_{m(n)}}\}_{n}$ by setting $h_{n}=v_{m(n)}^{n}$ for all $n$.
Now we prove that the constructed sequence $\{h_{n}:F_{q}\rightarrow\Gamma_{l_{m(n)}}\}_{n}$ over the ascending sequence $\{\Gamma_{l_{m(n)}}\}_{n}$ of groups in the model, contains a non-lift subsequence. Indeed, assume the converse.
We denote by $\Sigma_{\infty}$ the union of all the subsets $\Sigma_{n}$
[…]
By Guba’s theorem, every system over the free group $F_{k}$, is equivalent to a finite subsystem. In particular, the system $\Sigma_{\infty}$ (in the variables $x$, where $x$ is the fixed basis of $F_{q}$), is equivalent to a finite subsystem $\Sigma_{r_{0}}$ over $F_{k}$. Note that, by the construction of $\{\Sigma_{i}\}_{i}$ and $\{h_{n}=v_{m(n)}^{n}\}_{n}$, for every $i$, we have that $h_{n}(\Sigma_{i})=1$ eventually. In particular, we have that $h_{n}(\Sigma_{r_{0}})=1$ for all large enough $n$.
However, we are assuming that the sequence $\{h_{n}:F_{q}\rightarrow\Gamma_{l_{m(n)}}\}_{n}$ contains no non-lift subsequence, and hence, for all large enough $n$, the homomorphism $h_{n}=v_{m(n)}^{n}$ can be $\Sigma_{r_{0}}$-lifted. Since the system $\Sigma_{r_{0}}$ is equivalent to $\Sigma_{\infty}$ over $F_{k}$, we deduce that for all large enough $n$, the homomorphism $h_{n}=v_{m(n)}^{n}$ can be $\Sigma_{\infty}$-lifted, and in particular $\Sigma_{n}$-lifted. But this contradicts the defining property of the homomorphism $v_{m(n)}^{n}$ as one that cannot be $\Sigma_{n}$-lifted.
We deduce that the sequence $\{h_{n}:F_{q}\rightarrow\Gamma_{l_{m(n)}}\}_{n}$ contains a non-lift subsequence, and in accordance with 9.1, by passing to a subsequence, we can continue the proof under the assumption that the sequence $\{h_{n}:F_{q}\rightarrow\Gamma_{l_{m(n)}}\}_{n}$ is a non-lift sequence that converges into a non-lift $lln^{2}$-limit group that we denote by
$\eta_{\infty}:F_{q}\rightarrow R_{\infty}\,.$
By construction, the group $R_{\infty}$ is a proper quotient of $R_{n}$, for all $n$.
According to 9.12, the non-lift $lln^{2}$-limit group $R_{\infty}$, admits an infinite decreasing sequence of non-lift $lln^{2}$-limit quotients
$R_{\infty}>L_{1}>L_{2}>...\,.$
In particular, the group $L_{1}$ belongs to the collection $\mathcal{A}_{R_{n}}$ (defined in the beginning of the proof), for all $n$. Let $w\in F_{q}$ be a word representing a non-trivial element in $R_{\infty}$, but $w$ represents the trivial element in $L_{1}$. Denote by $n_{0}$ the length of $w$ in $F_{q}$. Then, the canonical map $\eta_{L_{1}}:F_{q}\rightarrow L_{1}$ maps to the identity strictly more elements of the ball of radius $n_{0}$ in the Cayley graph of $F_{q}$, than the map $\eta_{R_{n_{0}}}:F_{q}\rightarrow R_{n_{0}}$. This contradicts the defining property of $R_{n_{0}}$. ∎
Corollary 9.14.
There exist no non-lift $lln^{2}$-limit groups.
Proof.
According to 9.12 every non-lift $lln^{2}$-limit group admits an infinite strictly decreasing sequence of non-lift $lln^{2}$-limit quotients. And according to 9.13, every such sequence must terminate after finitely many steps. Hence there cannot exist a non-lift $lln^{2}$-limit group. ∎
9.3 Restricted Limit Groups over Ascending Sequences of Random Groups
Since we want to be able to treat sentences that contain constants, we should be able to lift solutions for systems of equations with constants. Recall that the constants are the words of the fixed free group $F_{k}$ with the fixed basis $a=(a_{1},...,a_{k})$.
Definition 9.15.
Let $F_{q}$ be a free group, with a basis $x=(x_{1},...,x_{q})$. Let $\{h_{l_{n}}:F_{q}\ast F_{k}\rightarrow\Gamma_{l_{n}}\}$ be a sequence of homomorphisms over an ascending sequence of groups in the model. Let $\pi:F_{k}\rightarrow\Gamma$ be a given presentation of a group $\Gamma$, and let $h:F_{q}\ast F_{k}\rightarrow\Gamma$ be a homomorphism.
  1. 1.
    We say that the homomorphism $h:F_{q}\ast F_{k}\rightarrow\Gamma$ is a restricted homomorphism, if $h(a_{i})=a_{i}$ (the right hand side means precisely $\pi(a_{i})$) for all $i=1,...,k$.
  2. 2.
    We say that the sequence $\{h_{l_{n}}:F_{q}\ast F_{k}\rightarrow\Gamma_{l_{n}}\}$ is a restricted sequence, If for all $n$, the homomorphism $h_{l_{n}}$ is restricted.
Theorem 9.16.
Let $F_{q}$ be a free group, with a basis $x=(x_{1},...,x_{q})$. Let $\{h_{l_{n}}:F_{q}\ast F_{k}\rightarrow\Gamma_{l_{n}}\}$ be a restricted sequence of homomorphisms over an ascending sequence of groups in the model.
Assume that the sequence of stretching factors
$\mu_{n}=\underset{1\leq u\leq q}{\max}d_{\Gamma_{l_{n}}}(1,h_{l_{n}}(x_{u}))\,,$
(note that here, due to the existence of constants, we do not post-compose with inner automorphisms) satisfies that
$\mu_{n}\geq l_{n}\ln^{2}l_{n}$
for all $n$.
Then, the sequence $h_{l_{n}}$, subconverges (in the Gromov topology) to an isometric non-trivial action of $F_{q}\ast F_{k}$ on a pointed real tree $(Y,y_{0})$, in which the subgroup $F_{k}$ is elliptic. Moreover, the real tree is not a real line.
Proof.
The same proof of 9.1. If $Y$ was a real line, then the generators of $F_{k}$ represent commuting elements in the groups $\Gamma_{l_{n}}$ for infinitely many integers $n$, which cannot happen according to 4.3. ∎
Definition 9.17.
Let $F_{q}$ be a free group, with a basis $x=(x_{1},...,x_{q})$. Let $\{h_{l_{n}}:F_{q}\ast F_{k}\rightarrow\Gamma_{l_{n}}\}$ be a restricted sequence of homomorphisms over an ascending sequence of groups in the model.
Assume that there exists a positive number $\epsilon>0$, so that the sequence of stretching factors
$\mu_{n}=\underset{1\leq u\leq q}{\max}d_{\Gamma_{l_{n}}}(1,h_{l_{n}}(x_{u}))\,,$
satisfies that
$\mu_{n}\geq\epsilon l_{n}\ln^{2}l_{n}$
for all $n$.
The sequence $h_{l_{n}}$, is called a restricted $lln^{2}$-sequence over the ascending sequence $\Gamma_{l_{n}}$ of groups in the model, or briefly, a restricted $lln^{2}$-sequence.
In case the sequence $h_{l_{n}}$ is a restricted $lln^{2}$-sequence that converges (in the Gromov topology) to an isometric non-trivial action of $F_{q}\ast F_{k}$ on a real tree, then, in accordance with the notation of 9.2, the quotient group
$L_{h_{l_{n}},\infty}=F_{q}\ast F_{k}/K_{h_{l_{n}},\infty}$
is called a restricted $lln^{2}$-limit group over the ascending sequence $\Gamma_{l_{n}}$ of groups in the model, or briefly, a restricted $lln^{2}$-limit group.
We say also that the restricted $lln^{2}$-sequence $h_{l_{n}}$ converges into the restricted $lln^{2}$-limit group $L_{h_{l_{n}},\infty}$.
Remark 9.18.
Let $\eta:F_{q}\ast F_{k}\rightarrow L_{\infty}$ be a restricted $lln^{2}$-limit group (with the canonical generating set associated to it). Then, the canonical map $\eta$ maps $F_{k}$ isomorphically into $L_{\infty}$. This follows by 4.3 and the appropriate version of 9.2 for the restricted case (see the discussion below).
Definition 9.19.
Let $F_{q}$ be a free group, with a fixed basis $x=(x_{1},...,x_{q})$, and let $\{h_{l_{n}}:F_{q}\ast F_{k}\rightarrow\Gamma_{l_{n}}\}$ be a restricted sequence of homomorphisms over an ascending sequence of groups in the model. Let $\pi:F_{k}\rightarrow\Gamma$ be a given presentation of a group $\Gamma$, and let $h:F_{q}\ast F_{k}\rightarrow\Gamma$ be a restricted homomorphism.
  1. 1.
    Let $\Sigma$ be a collection of words in the free group $F_{q}\ast F_{k}$. We say that the homomorphism $h$ can be $\Sigma$-lifted restrictly , if there exists some restricted homomorphism $\tilde{h}:F_{q}\ast F_{k}\rightarrow F_{k}$ with $\tilde{h}(\Sigma)=1$ and $h=\pi\circ\tilde{h}$.
  2. 2.
    We say that the restricted sequence $\{h_{l_{n}}:F_{q}\ast F_{k}\rightarrow\Gamma_{l_{n}}\}$ is a restricted non-lift sequence, if there exists a finite set of words $\Sigma\subset F_{q}\ast F_{k}$, so that $h_{l_{n}}(\Sigma)=1$ eventually, but for all $n$, the homomorphism $h_{l_{n}}$ cannot be $\Sigma$-lifted restrictly.
  3. 3.
    If the sequence $\{h_{l_{n}}:F_{q}\ast F_{k}\rightarrow\Gamma_{l_{n}}\}$ is a restricted non-lift sequence that converges to a restricted $lln^{2}$-limit group $L_{\infty}$, then the group $L_{\infty}$ is called a restricted non-lift $lln^{2}$-limit group.
Remark 9.20.
Since we have dropped the collection $\mathcal{N}$, we can assume that any restricted non-lift sequence is in necessity a restricted $lln^{2}$-sequence.
Now the formulation of 9.2 for the restricted case (w.r.t. to the notation of 9.16 instead of that of 9.1) can be proved by the same argument in 9.2. LABEL:Def:53 stays unchanged for the restricted case. Point 1 of 9.7 is replaced by the fact that the limit tree is not a real line in the restricted case. The abelian decompositions that we are interested in, for a restricted non-lift $lln^{2}$-limit group $L_{\infty}$, are only those for which the subgroup $F_{k}\leq L_{\infty}$ is elliptic in them (i.e., can be conjugated into a vertex group). Hence, Point 2 of 9.7 naturally obtains a formulation that is consistent with this. LABEL:Lem:55 stays unchanged for the restricted case.
Let $L_{\infty}$ be a restricted non-lift $lln^{2}$-limit group, and assume that $L_{\infty}$ is freely indecomposable w.r.t. the subgroup $F_{k}$.
For defining the restricted abelian JSJ decomposition associated to the restricted non-lift $lln^{2}$-limit group $L_{\infty}$, we recall again that the abelian decompositions of $L_{\infty}$ that we are interested in, are only those for which the subgroup $F_{k}\leq L_{\infty}$ is elliptic in them.
In light of the versions of 9.8, 9.1, and 9.2 for the restricted case, Chapter 2 of [Se1] naturally generalizes (in the same way as it generalizes for restricted limit groups in the standard context of [Se1]), in order to construct the restricted abelian JSJ decomposition of the restricted freely indecomposable (w.r.t. $F_{k}$) non-lift $lln^{2}$-limit group $L_{\infty}$.
In particular, the subgroup $F_{k}\leq L_{\infty}$ is elliptic in the restricted JSJ of $L_{\infty}$. For the exact statement on how the restricted (abelian) JSJ decomposition encodes all the abelian splittings of $L_{\infty}$ in which $F_{k}$ is elliptic, see Theorem 8.1 (the restricted case) of [Se1].
LABEL:Def:56 together with 9.10 and 9.11, obtain a natural form for the restricted case, without changing the arguments. Thus, 9.12 obtains also a natural form with the same argument for the restricted case.
LABEL:Thm:60 stays unchanged for the restricted case.
And hence 9.14 obtains the following form:
Theorem 9.21.
There exist no restricted non-lift $lln^{2}$-limit groups.
10 Consequences
10.1 Lifting All the Solutions of a Given System over Random Groups
Let $k\geq 2$ be an integer, let $a=(a_{1},...,a_{k})$ be a given tuple of formal letters, and let $F_{k}$ be the free group with basis $a$. Let $d<\frac{1}{2}$ be a real number, and consider the Gromov model of density $d$.
Theorem 10.1.
Let $\Sigma_{0}=\Sigma_{0}(y,a)$ be a system of equations in the variables $y$ and constants (maybe empty) $a$. Then, the random group $\Gamma$ of density $d$ satisfies the following property with overwhelming probability.
Every solution of the system $\Sigma_{0}$ in $\Gamma$, can be lifted to a solution of $\Sigma_{0}$ in $F_{k}$.
Proof.
[…]
We denote the collection of all the groups in level $l$ of the model of density $d$ by $\mathcal{M}_{l}$.
According to 9.21, there exists an integer $l_{0}$, so that for all integers $l\geq l_{0}$, every group $\Gamma_{l}$ in $\mathcal{M}_{l}\backslash\mathcal{N}_{l}$ satisfies the $\Sigma_{0}$-l.p., i.e., every solution of $\Sigma_{0}$ in $\Gamma_{l}$ can be lifted to a solution of $\Sigma_{0}$ in $F_{k}$.
[…]
$\frac{|\mathcal{N}_{l}|}{|\mathcal{M}_{l}|}\overset{l\rightarrow\infty}{\longrightarrow}0\,,$
and the claim follows. ∎
Theorem 10.2.
Let $\Sigma_{0}=\Sigma_{0}(y,a)$ be a system of equations in the variables $y$ and constants (maybe empty) $a$. Let $MR$ be the Makanin-Razborov diagram of the system $\Sigma_{0}$ constructed over the free group $F_{k}$. Then, the random group $\Gamma$ of density $d$ satisfies the following property with overwhelming probability.
Every solution of the system $\Sigma_{0}$ in $\Gamma$, factors through $MR$.
Proof.
Every solution of $\Sigma_{0}$ in $\Gamma$ can be lifted to a solution of $\Sigma_{0}$ in $F_{k}$. By the construction of $MR$ (see [[Se1], sections 5 and 8]), every solution of $\Sigma_{0}$ in $F_{k}$ factors through $MR$. Hence, every solution of $\Sigma_{0}$ in $\Gamma$ factors through $MR$. ∎
10.2 Universal Sentence over Random Groups
Let $k\geq 2$ be an integer, let $a=(a_{1},...,a_{k})$ be a given tuple of formal letters, and let $F_{k}$ be the free group with basis $a$. Let $d<\frac{1}{2}$ be a real number, and consider the Gromov model of density $d$.
Theorem 10.3.
Let $\psi=\psi(x,a)$ be a sentence in the Boolean algebra of one-quantifier sentences (which may or may not contain constants). Let $\Gamma$ be the random group of density $d$.
Then, the sentence $\psi$ is a truth sentence over the free group $F_{k}$, if and only if, with overwhelming probability $\psi$ is a truth sentence over $\Gamma$.
Proof.
Assume that $\psi$ is an existential sentence. Then $\psi$ is equivalent (tautologically - over any group) to a sentence of the form
$\stackrel{{\scriptstyle[}}{{i}}=1]{r}{\vee}\left(\exists x\quad\Sigma_{i}(x,a)=1\wedge v_{1}^{i}(x,a)\neq 1\wedge...\wedge v_{n_{i}}^{i}(x,a)\neq 1\right)\,,$
where $r,n_{i}$ are integers, $\Sigma_{i}(x,a)=1$ is a system of equations, and $v_{j}^{i}(x,a)$ is a word in the free group $F(x,a)$, $i=1,...,r$, $j=1,...,n_{i}$.
We denote by $p_{l}$ the probability that the random group of level $l$ and density $d$ satisfies the sentence $\psi$.
Assume that $\psi$ is true over $F_{k}$. Then, according to 4.3, it follows directly that with overwhelming probability, the random group $\Gamma$ of density $d$ satisfies $\psi$
$p_{l}\overset{l\rightarrow\infty}{\longrightarrow}1\,.$
[…]
Since $\psi$ is true over $\Gamma_{l}$, there exists $i_{0}=1,...,r$, and some solution $x_{0}$ for the system $\Sigma_{i_{0}}(x,a)$ over $\Gamma_{l}$, so that
$v_{1}^{i_{0}}(x_{0},a)\neq 1\wedge...\wedge v_{n_{i_{0}}}^{i_{0}}(x_{0},a)\neq 1\,$
in $\Gamma_{l}$.
Let $\tilde{x}_{0}$ be a lift for the solution $x_{0}$ to a solution for the system $\Sigma_{i_{0}}(x,a)$ over the free group $F_{k}$. Then
$\Sigma_{i_{0}}(\tilde{x}_{0},a)=1\wedge v_{1}^{i_{0}}(\tilde{x}_{0},a)\neq 1\wedge...\wedge v_{n_{i_{0}}}^{i_{0}}(\tilde{x}_{0},a)\neq 1\,$
in $F_{k}$, which implies that $\psi$ is a truth sentence over $F_{k}$.
This argument, obviously, implies the statement in the theorem. ∎